### v1

# Networked Communication for Mean-Field Games with Function Approximation and Empirical Mean-Field Estimation

Abstract
Recent works have provided algorithms by which decentralised agents, which may be connected via a communication network, can learn equilibria in Mean-Field Games from a single, non-episodic run of the empirical system. However, these algorithms are given for tabular settings: this computationally limits the size of players’ observation space, meaning that the algorithms are not able to handle anything but small state spaces, nor to generalise beyond policies depending on the ego player’s state to so-called ‘population-dependent’ policies. We address this limitation by introducing function approximation to the existing setting, drawing on the Munchausen Online Mirror Descent method that has previously been employed only in finite-horizon, episodic, centralised settings. While this permits us to include the population’s mean-field distribution in the observation for each player’s policy, it is arguably unrealistic to assume that decentralised agents would have access to this global information: we therefore additionally provide new algorithms that allow agents to estimate the global empirical distribution based on a local neighbourhood, and to improve this estimate via communication over a given network. Our experiments showcase how the communication network allows decentralised agents to estimate the mean-field distribution for population-dependent policies, and that exchanging policy information helps networked agents to outperform both independent and even centralised agents in function-approximation settings, by an even greater margin than in tabular settings.
1 Introduction
The mean-field game (MFG) framework (Lasry and Lions 2007; Huang, Malhamé, and Caines 2006) can be used to address the difficulty faced by multi-agent reinforcement learning (MARL) regarding computational scalability as the number of agents increases. It models a representative agent as interacting not with other individual agents in the population on a per-agent basis, but instead with a distribution over the other agents, known as the mean field. The MFG framework analyses the limiting case when the population consists of an infinite number of symmetric and anonymous agents, that is, they have identical reward and transition functions which depend on the mean-field distribution rather than on the actions of specific other players. The solution to this game is the mean-field Nash equilibrium (MFNE), which can be used as an approximation for the Nash equilibrium (NE) in a finite-agent game, with the error in the solution reducing as the number of agents N tends to infinity (Saldi, Başar, and Raginsky 2018; Anahtarci, Kariksiz, and Saldi 2023; Yardim, Goldman, and He 2024; Toumi, Malhame, and Le Ny 2024; Hu and Zhang 2024). MFGs have thus been applied to a wide variety of real-world problems: see Laurière et al. (2022) for applications and references.
Recent works argue that classical algorithms for solving MFGs rely on assumptions and methods that are likely to be undesirable in real-world applications, emphasising that desirable qualities for practical MFG algorithms include: learning from the empirical distribution of $N$ agents (without generation or manipulation of this distribution by the algorithm itself or an external oracle/simulator); learning from a single, continued system run that is not arbitrarily reset as in episodic learning; model-free learning; decentralisation; and fast practical convergence (Yardim et al. 2023; Benjamin and Abate 2024). While these works address these desiderata, they do so only in settings in which the state and action spaces are small enough that the Q-function can be represented by a table, limiting their approaches’ scalability.
Moreover, in those works, as in many others on MFGs, agents only observe their local state as input to their Q-function (which defines their policy). This is sufficient when the solved MFG is expected to have a stationary distribution (‘stationary MFGs’) (Laurière et al. 2022; Xie et al. 2021; Anahtarci, Kariksiz, and Saldi 2023; uz Zaman et al. 2023; Yardim et al. 2023; Benjamin and Abate 2024). However, in reality there are numerous reasons for which agents may benefit from being able to respond to the current distribution. Recent work has thus increasingly focused on these more general settings where it is necessary for agents to have so-called ‘master policies’ (a.k.a. population-dependent policies) which depend on both the mean-field distribution and their local state (Cardaliaguet et al. 2015; Carmona, Delarue, and Lacker 2016; Perrin et al. 2022; Wu et al. 2024; Laurière et al. 2022; Lauriere et al. 2022; Laurière et al. 2022).
The distribution is a high-dimensional (and thus large) observation object, and also takes a continuum of values. Therefore a population-dependent Q-function cannot be represented exactly in a table and must be approximated. To address these limitations while maintaining the desiderata for real-world applications given in recent works, we introduce function approximation to the MFG setting of decentralised agents learning from a single, non-episodic run of the empirical system, allowing this setting to handle larger state spaces and to accept the mean-field distribution as an observation input. To overcome the difficulties of training non-linear approximators in this context, we use the so-called ‘Munchausen’ trick, introduced by Vieillard, Pietquin, and Geist (2020) for single-agent RL, and extended to MFGs by Lauriere et al. (2022), and to MFGs with population-dependent policies by Wu et al. (2024).
We particularly explore this in the context of networked communication between decentralised agents (Benjamin and Abate 2024). We demonstrate that this communication brings two specific benefits over the purely independent setting (while also removing the undesirable assumption of a centralised learner, which in the real world may be unrealistic, a computational bottleneck and a vulnerable single point of failure). Firstly, when the Q-function is approximated rather than exact, some updates lead to better performing policies than others. Allowing networked agents to propagate better performing policies through the population leads to faster practical convergence than in the purely independent-learning case, and often even than in the centralised case. Secondly, we argue that in the real world it is unrealistic to assume that decentralised agents, endowed with local state observations and limited (if any) communication radius, would be able to observe the global mean-field distribution and use it as input to their Q-functions / policy. We therefore further contribute two setting-dependent algorithms by which decentralised agents can estimate the global distribution from local observations, and further improve their estimates by communication with neighbours.
We focus on ‘coordination games’, where agents can increase their individual rewards by following the same strategy as others and therefore have an incentive to communicate policies, even if the MFG setting itself is technically non-cooperative. Thus our work can be applied to real-world problems in e.g. traffic signal control, formation control in swarm robotics, and consensus and synchronisation e.g. for sensor networks. In summary, our contributions are:
  • We introduce, for the first time, function approximation to MFG settings with decentralised agents. To do this:
    • We use Munchausen RL for the first time in an infinite-horizon MFG context, unlike the finite-horizon settings in Lauriere et al. (2022) and Wu et al. (2024).
    • In doing so, we also present the first use of function approximation for solving MFGs from continued, non-episodic runs of the empirical system; this is unlike Yardim et al. (2023) and Benjamin and Abate (2024), which look at tabular settings.
  • Function approximation allows us to explore larger state spaces, and also settings where agents’ policies depend on the mean-field distribution as well as their local state.
  • Rather than assuming that agents have access to this global knowledge, as in prior works, we present two novel algorithms that allow decentralised agents to locally estimate the empirical distribution and to improve these estimates by communication with others.
  • We present numerical results showcasing these algorithms, especially the two benefits of the decentralised communication scheme, which significantly outperforms both the independent and centralised settings.
The paper is structured as follows. Preliminaries are given in Sec. 2 and our core learning and policy improvement algorithm is presented in Sec. 3. Our mean-field estimation and communication algorithms are given in Sec. 4, and experiments are presented in Sec. 5. See Appx. B for a more detailed ‘Related work’ section. Ways to extend our algorithms are discussed in the ‘Future Work’ section (Appx. C).
2 Preliminaries
2.1 Mean-field games
We use the following notation. $N$ is the number of agents in a population, with $\mathcal{S}$ and $\mathcal{A}$ representing the finite state and common action spaces, respectively. The set of probability measures on a finite set $\mathcal{X}$ is denoted $\Delta_{\mathcal{X}}$, and $\mathbf{e}_{x}\in\Delta_{\mathcal{X}}$ for $x\in\mathcal{X}$ is a one-hot vector with only the entry corresponding to $x$ set to 1, and all others set to 0. For time $t\geq 0$, $\hat{\mu}_{t}$ = $\frac{1}{N}\sum^{N}_{i=1}\sum_{s\in\mathcal{S}}$ $\mathbb{1}_{s^{i}_{t}=s}\mathbf{e}_{s}$in $\Delta_{\mathcal{S}}$ is a vector of length $|\mathcal{S}|$ denoting the empirical categorical state distribution of the $N$ agents at time $t$. For agent $i\in 1\dots N$, $i$’s policy at time $t$ depends on its observation $o^{i}_{t}$. We explore three different forms that this observation object can take:
  • In the conventional setting, the observation is simply $i$’s current local state $s^{i}_{t}$, such that $\pi^{i}(a|o^{i}_{t})=\pi^{i}(a|s^{i}_{t})$.
  • When the policy is population-dependent, if we assume perfect observability of the global mean-field distribution then we have $o^{i}_{t}=(s^{i}_{t},\hat{\mu}_{t})$.
  • It is unrealistic to assume that decentralised agents with a possibly limited communication radius can observe the global mean-field distribution, so we allow agents to form a local estimate $\tilde{\hat{\mu}}^{i}_{t}$ which can be improved by communication with neighbours. Here we have $o^{i}_{t}=(s^{i}_{t},\tilde{\hat{\mu}}^{i}_{t})$.
In the following definitions we focus on the population-dependent case when $o^{i}_{t}=(s^{i}_{t},\hat{\mu}_{t})$, and clarify afterwards the connection to the other observation cases. Thus the set of policies is $\Pi$ = $\pi$ : $\mathcal{S}\times\Delta_{\mathcal{S}}\rightarrow\Delta_{\mathcal{A}}$}, and the set of Q-functions is denoted $\mathcal{Q}=\{q:\mathcal{S}\times\Delta_{\mathcal{S}}\times\mathcal{A}\rightarrow\mathbb{R}\}$.
Definition 1 (N-player symmetric anonymous games)
[…]
$s^{i}_{t+1}\sim P(\cdot|s^{i}_{t},a^{i}_{t},\hat{\mu}_{t}),\;\;\;r^{i}_{t}=R(s^{i}_{t},a^{i}_{t},\hat{\mu}_{t}).$
At the limit as $N\rightarrow\infty$, the infinite population of agents can be characterised as a limit distribution $\mu\in\Delta_{\mathcal{S}}$; the infinite-agent game is termed an MFG. The so-called ‘mean-field flow’ $\boldsymbol{\mu}$ is given by the infinite sequence of mean-field distributions s.t. $\boldsymbol{\mu}=$$(\mu_{t})_{t\geq 0}$.
Definition 2 (Induced mean-field flow)
We denote by $I(\pi)$ the mean-field flow $\boldsymbol{\mu}$ induced when all the agents follow $\pi$, where this is generated from $\pi$ as follows:
$\mu_{t+1}(s^{\prime})=\sum_{s,a}\mu_{t}(s)\pi(a|s,\mu_{t})P(s^{\prime}|s,a,\mu_{t}).$
When the mean-field flow is stationary such that the distribution is the same for all $t$, i.e. $\mu_{t}=\mu_{t+1}$ $\forall t\geq 0$, the policy $\pi^{i}(a|s^{i}_{t},{\mu}_{t})$ need not depend on the distribution, such that $\pi^{i}(a|s^{i}_{t},{\mu}_{t})$ = $\pi^{i}(a|s^{i}_{t})$, i.e. we recover the classical population-independent policy. However, for such a population-independent policy the initial distribution $\mu_{0}$ must be known and fixed in advance, whereas otherwise it need not be. We also give the following definitions.
Definition 3 (Mean-field discounted return)
In a MFG where all agents follow policy $\pi$ giving a mean-field flow $\boldsymbol{\mu}=$$(\mu_{t})_{t\geq 0}$, the expected discounted return of the representative agent is given by
$V(\pi,\boldsymbol{\mu})=\mathbb{E}\left[\sum^{\infty}_{t=0}\gamma^{t}(R(s_{t},a_{t},\mu_{t}))\bigg{|}\begin{subarray}{c}s_{0}\sim\mu_{0}\\ a_{t}\sim\pi(\cdot|s_{t},\mu_{t})\\ s_{t+1}\sim P(\cdot|s_{t},a_{t},{\mu_{t}})\end{subarray}\right].$
Definition 4 (Best-response (BR) policy)
A policy $\pi^{*}$ is a best response (BR) against the mean-field flow $\boldsymbol{\mu}$ if it maximises the discounted return $V(\cdot,\boldsymbol{\mu})$; the set of these policies is denoted $BR(\boldsymbol{\mu})$, i.e.
$\pi^{*}\in BR(\boldsymbol{\mu}):=\mathop{\arg\max}\limits_{\pi}V(\pi,\boldsymbol{\mu}).$
Definition 5 (Mean-field Nash equilibrium (MFNE))
A pair ($\pi^{*},\boldsymbol{\mu}^{*})$ is a mean-field Nash equilibrium (MFNE) if the following two conditions hold:
  • $\pi^{*}$ is a best response to $\boldsymbol{\mu}^{*}$, i.e. $\pi^{*}\in BR(\boldsymbol{\mu})$;
  • $\boldsymbol{\mu}^{*}$ is induced by $\pi^{*}$, i.e. $\boldsymbol{\mu}^{*}$ = $I(\pi^{*})$.
$\pi^{*}$ is thus a fixed point of the map $BR\circ I$, i.e. $\pi^{*}\in BR(I(\pi^{*}))$. If a population-dependent policy is a MFNE policy for any initial distribution $\mu_{0}$, it is a ‘master policy’.
Previous works have shown that, in tabular settings, it is possible for a finite population of decentralised agents (each of which is permitted to have a distinct population-independent policy $\pi^{i}$) to learn the MFNE using only the empirical distribution $\hat{\mu}_{t}$, rather than the exactly calculated mean-field flow (Yardim et al. 2023; Benjamin and Abate 2024). This MFNE is in turn an approximate NE for the original finite-agent game. In this work we provide algorithms to perform this in non-tabular and population-dependent settings, and demonstrate them empirically.
2.2 (Munchausen) Online Mirror Descent
Instead of performing the computationally expensive process of finding a $BR$ at each iteration, we can use a form of policy iteration for MFGs called Online Mirror Descent (OMD). This involves beginning with an initial policy $\pi_{0}$, and then at each iteration $k$, evaluating the current policy $\pi_{k}$ with respect to its induced mean-field flow $\boldsymbol{\mu}=I(\pi_{k})$ to compute its Q-function $Q_{k+1}$. To stabilise the learning process, we then use a weighted sum over this and past Q-functions, and set $\pi_{k+1}$ to be the softmax over this weighted sum, i.e. $\pi_{k+1}(\cdot|s,\mu)=softmax\left(\frac{1}{\tau_{q}}\sum^{k}_{\kappa=0}{Q_{\kappa}}(s,\mu,\cdot)\right)$. $\tau_{q}$ is a temperature parameter that scales the entropy in Munchausen RL (Vieillard, Pietquin, and Geist 2020); note that this is a different temperature to the one agents use when selecting which communicated parameters to adopt, denoted $\tau^{comm}_{k}$ and discussed in Sec. 3.2.
If the Q-function is approximated non-linearly using neural networks, it is difficult to compute this weighted sum. The so-called ‘Munchausen trick’ addresses this by computing a single Q-function that mimics the weighted sum using implicit regularisation based on the Kullback-Leibler (KL) divergence between $\pi_{k}$ and $\pi_{k+1}$ (Vieillard, Pietquin, and Geist 2020). Using this reparametrisation gives Munchausen OMD (MOMD), detailed further in Sec. 3.1 (Lauriere et al. 2022; Wu et al. 2024). MOMD does not bias the MFNE, and has the same convergence guarantees as OMD (Hadikhanloo 2017; Pérolat et al. 2022a; Wu et al. 2024).
2.3 Networks
We conceive of the finite population as exhibiting two time-varying networks. The basic definition of such a network is:
Definition 6 (Time-varying network)
[…]
One of these networks $\mathcal{G}^{comm}_{t}$ defines which agents are able to communicate information to each other at time $t$. The second network $\mathcal{G}^{obs}_{t}$ is a graph defining which agents are able to observe each other’s states, which we use in general settings for estimating the mean-field distribution from local information. The structure of the two networks may be identical (e.g. if embodied agents can both observe the position (state) of, and exchange information with, other agents within a certain physical distance from themselves), or different (e.g. if agents can observe the positions of nearby agents, but only exchange information with agents by which they are connected via a telephone line, which may connect agents over long distances).
We also define an alternative version of the observation graph that is useful in a specific subclass of environments, which can most intuitively be thought of as those where agents’ states are positions in physical space. When this is the case, we usually think of agents’ ability to observe each other as depending more abstractly on whether states are visible to each other. We define this visibility graph as follows:
Definition 7 (Time-varying state-visibility graph)
[…]
We view an agent in $s$ as able to obtain a count of the number of agents in $s^{\prime}$ if $s^{\prime}$ is visible to $s$. The benefit of this graph $\mathcal{G}^{vis}_{t}$ over $\mathcal{G}^{obs}_{t}$ is that there is mutual exclusivity: either an agent in state $s$ is able to obtain a total count of all of the agents in state $s^{\prime}$ (if $s^{\prime}$ is visible to $s$), or it cannot obtain information about any agent in state $s^{\prime}$ (if those states are not visible to each other). Additionally, this graph permits an agent in state $s$ to observe that there are no agents in state $s^{\prime}$ as long as $s^{\prime}$ is visible to $s$. These benefits are not available if the observability graph is defined strictly between agents as in $\mathcal{G}^{obs}_{t}$, such that using $\mathcal{G}^{vis}_{t}$ facilitates more efficient estimation of the global mean-field distribution from local information in settings where $\mathcal{G}^{vis}_{t}$ is applicable (see Sec. 4).
3 Learning and policy improvement
3.1 Q-network and update
Algorithm 1 Networked learning with non-linear function approximation
0:  loop parameters $K,M,L,E,C_{p}$, learning parameters $\gamma,\tau_{q},|B|,cl,\nu$, $\{\tau^{comm}_{k}\}_{k\in\{0,\dots,K-1\}}$
0:  initial states {$s^{i}_{0}$}i, $i=1,\ldots,N$; $t\leftarrow 0$
1:  $\forall i$ : Randomly initialise parameters $\theta^{i}_{0}$ of Q-networks $\check{Q}_{\theta^{i}_{0}}(o^{i}_{t},\cdot)$, and set $\pi^{i}_{0}(a|o)=$ softmax$\left(\frac{1}{\tau_{q}}\check{Q}_{\theta^{i}_{0}}(o,\cdot)\right)(a)$
2:  for $k=0,\dots,K-1$ do
3:     $\forall i$: Empty $i$’s buffer
4:     for $m=0,\dots,M-1$ do
5:        Take step $\forall i:a^{i}_{t}\sim\pi^{i}_{k}(\cdot|o^{i}_{t}),r^{i}_{t}=R(s^{i}_{t},a^{i}_{t},\hat{\mu}_{t}),s^{i}_{t+1}\sim P(\cdot|s^{i}_{t},a^{i}_{t},\hat{\mu}_{t})$; $t\leftarrow t+1$
6:        $\forall i$: Add $\zeta^{i}_{t}$ to $i$’s buffer
7:     end for
8:     for $l=0,\dots,L-1$ do
9:        $\forall i:$ Sample batch $B^{i}_{k,l}$ from $i$’s buffer
10:        Update $\theta$ to minimise ^$\hat{\mathcal{L}}(\theta,\theta^{\prime})$ as in Def. 8
11:        If $l\mod$ $\nu=0$, set $\theta^{\prime}\leftarrow\theta$
12:     end for
13:     $\check{Q}_{\theta^{i}_{k+1}}(o,\cdot)\leftarrow\check{Q}_{\theta^{i}_{k,L}}(o,\cdot)$
14:     $\forall i$ : Set $\pi^{i}_{k+1}(a|o)=$ softmax$\left(\frac{1}{\tau_{q}}\check{Q}_{\theta^{i}_{k+1}}(o,\cdot)\right)(a)$
15:     $\forall i:$ $\sigma^{i}_{k+1}=0$
16:     for $e=0,\dots,E-1$ evaluation steps do
17:        Take step $\forall i:a^{i}_{t}\sim\pi^{i}_{k}(\cdot|o^{i}_{t}),r^{i}_{t}=R(s^{i}_{t},a^{i}_{t},\hat{\mu}_{t}),s^{i}_{t+1}\sim P(\cdot|s^{i}_{t},a^{i}_{t},\hat{\mu}_{t})$; $t\leftarrow t+1$
18:        $\forall i:$ $\sigma^{i}_{k+1}=\sigma^{i}_{k+1}+\gamma^{e}(r^{i}_{t}+h(\pi^{i}_{k+1}(s^{i}_{t})))$
19:        $t\leftarrow t+1$
20:     end for
21:     for $C_{p}$ rounds do
22:        $\forall i:$ Broadcast $\sigma^{i}_{k+1},\pi^{i}_{k+1}$
23:        $\forall i:J^{i}_{t}\leftarrow\{j\in\mathcal{N}:(i,j)\in\mathcal{E}_{t}$}
24:        $\forall i:$ Select $\mathrm{adopted}^{i}\sim$ Pr$\left(\mathrm{adopted}^{i}=j\right)$ = $\frac{\exp{(\sigma^{j}_{k+1}}/\tau^{comm}_{k})}{\sum_{x\in J^{i}_{t}}\exp{(\sigma^{x}_{k+1}}/\tau^{comm}_{k})}$ $\forall j\in J^{i}_{t}$
25:        $\forall i:\sigma^{i}_{k+1}\leftarrow\sigma^{\mathrm{adopted}^{i}}_{k+1},\pi^{i}_{k+1}\leftarrow\pi^{\mathrm{adopted}^{i}}_{k+1}$
26:        Take step $\forall i:a^{i}_{t}\sim\pi^{i}_{k}(\cdot|o^{i}_{t}),r^{i}_{t}=R(s^{i}_{t},a^{i}_{t},\hat{\mu}_{t}),s^{i}_{t+1}\sim P(\cdot|s^{i}_{t},a^{i}_{t},\hat{\mu}_{t})$; $t\leftarrow t+1$
27:     end for
28:  end for
29:  return  policies $\pi^{i}_{K}$}i, $i=1,\ldots,N$
Lines 1-14 of our novel Alg. 1 contain the core Q-function/policy update method. Agent $i$ maintains a neural network parametrised by $\theta^{i}_{k}$ to approximate its Q-function: $\check{Q}_{\theta^{i}_{k}}(o,\cdot)$. The agent’s policy is given by $\pi_{\theta^{i}_{k}}(a|o)=$ softmax$\left(\frac{1}{\tau_{q}}\check{Q}_{\theta^{i}_{k}}(o,\cdot)\right)(a)$. When appropriate, we denote the policy $\pi^{i}_{k}(a|o)$ for simplicity. Each agent maintains a buffer (with size $M$) of collected transitions of the form $\left(o^{i}_{t},a_{t},r_{t},o^{i}_{t+1}\right)$. At each iteration $k$, they empty their buffer (Line 3) before collecting $M$ new transitions (Lines 4-7); each decentralised agent $i$ then trains its Q-network $\check{Q}_{\theta^{i}_{k}}$ via $L$ training updates as follows (Lines 8-12). For training purposes, $i$ also maintains a target network $\check{Q}_{\theta^{i,^{\prime}}_{k,l}}$ with the same architecture but parameters $\theta^{i,^{\prime}}_{k,l}$ copied from $\theta^{i}_{k,l}$ less regularly than $\theta^{i}_{k,l}$ themselves are updated, i.e. only every $\nu$ learning iterations (Line 11). At each iteration $l$, the agent samples a random batch $B^{i}_{k,l}$ of $|B|$ transitions from its buffer (Line 9). It then trains its neural network using stochastic gradient descent to minimise the following empirical loss (Line 10):
Definition 8 (Empirical loss for Q-network)
The empirical loss is given by
^$\hat{\mathcal{L}}(\theta,\theta^{\prime})=\frac{1}{|B|}\sum_{transition\in B^{i}_{k,l}}\left|\check{Q}_{\theta^{i}_{k,l}}(o_{t},a_{t})-T\right|^{2},$
where the target $T$ is
$T=r_{t}\;+\;[\tau_{q}\ln\pi_{\theta^{i,^{\prime}}_{k,l}}(a_{t}|o_{t})]^{0}_{cl}\;+$
$\gamma\sum_{a\in\mathcal{A}}\pi_{\theta^{i,^{\prime}}_{k,l}}(a|o_{t+1})\left(\check{Q}_{\theta^{i,^{\prime}}_{k,l}}(o_{t+1},a)-\tau_{q}\ln\pi_{\theta^{i,^{\prime}}_{k,l}}(a|o_{t+1})\right).$
For $cl<0$, $[\cdot]^{0}_{cl}$ is a clipping function used in Munchausen RL to prevent numerical issues if the policy is too close to deterministic, as the log-policy term is otherwise unbounded (Vieillard, Pietquin, and Geist 2020; Wu et al. 2024).
3.2 Communication and adoption of parameters
We use the communication network $\mathcal{G}^{comm}_{t}$ to share two types of information at different points in Alg 1. One is used to improve local estimates of the mean-field distribution (see Sec. 4). The other, described here, is used to privilege the spread of better performing policy updates through the population, allowing faster convergence in this networked case than in the independent and even centralised cases.
We extend the work in Benjamin and Abate (2024) to the function-approximation case, where in our work agents broadcast the parameters of the Q-network that defines their policy, rather than the Q-function table. At each iteration $k$, after independently updating their Q-network and policy (Lines 3-14), agents approximately evaluate their new policies by collecting rewards for $E$ steps, and assign the estimated value to $\sigma^{i}_{k+1}$ (Lines 15-20). They then broadcast their Q-network parameters along with $\sigma^{i}_{k+1}$ (Line 22). Receiving these from their neighbours on the communication network, agents select which set of parameters to adopt by taking a softmax over their own and the received estimate values $\sigma^{j}_{k+1}$ (Lines 23-25). They repeat this broadcast and adoption process for $C_{p}$ rounds, which is distinct from the $C_{e}$ communication rounds discussed in Sec. 4. By this process, decentralised agents can adopt policy parameters that are estimated to be better performing than their own, bringing the benefits to convergence speed discussed in Sec. 5.1.
4 Mean-field estimation and communication
We first describe the most general version of our algorithm for decentralised estimation of the empirical categorical mean-field distribution, assuming the more general setting where $\mathcal{G}^{obs}_{t}$ applies (see discussion in Sec. 2.3). We subsequently detail how the algorithm can be made more efficient in environments where the more abstract visibility graph $\mathcal{G}^{vis}_{t}$ applies, as in our experimental settings. In both cases, the algorithm runs to generate the observation object when a step is taken in the main Alg. 1, i.e. to produce $o^{i}_{t}=(s^{i}_{t},\tilde{\hat{\mu}}^{i}_{t})$ for the steps $a^{i}_{t}\sim\pi^{i}_{k}(\cdot|o^{i}_{t})$ in Lines 5, 17 and 26. Both versions of the algorithm are subject to implicit assumptions, which we highlight and discuss methods for addressing in our ‘Future Work’ section in Appx. C.
Algorithm for the general setting
Algorithm 2 Mean-field estimation and communication algorithm in general settings
0:  Time-dependent observation graph $\mathcal{G}^{obs}_{t}$, time-dependent communication graph $\mathcal{G}^{comm}_{t}$, states $\{s^{i}_{t}\}_{i=1}^{N}$, number of communication rounds $C_{e}$
1:  $\forall i,s:$ Initialise count vector $\hat{\upsilon}^{i}_{t}[s]$ with $\emptyset$
2:  $\forall i:$ $\hat{\upsilon}^{i}_{t}[s^{j}_{t}]$leftarrow $\{ID^{j}\}_{j\in\mathcal{N}:(i,j)\in\mathcal{E}^{obs}_{t}}$
3:  for $c_{e}$ in $1,\dots,C_{e}$ do
4:     $\forall i:$ Broadcast $\hat{\upsilon}^{i}_{t,c_{e}}$
5:     $\forall i:J^{i}_{t}\leftarrow\{j\in\mathcal{N}:(i,j)\in\mathcal{E}^{comm}_{t}\}$
6:     $\forall i,s$ : $\hat{\upsilon}^{i}_{t,(c_{e}+1)}[s]\leftarrow\hat{\upsilon}^{i}_{t,c_{e}}[s]\cup\{\hat{\upsilon}^{j}_{t,c_{e}}[s]\}_{j\in J^{i}_{t}}$
7:  end for
8:  $\forall i:$ $counted\_agents^{i}_{t}\leftarrow\sum_{s\in\mathcal{S}:\hat{\upsilon}^{i}_{t}[s]\neq\emptyset}|\hat{\upsilon}^{i}_{t}[s]|$
9:  $\forall i:$ $uncounted\_agents^{i}_{t}\leftarrow N-counted\_agents^{i}_{t}$
10:  $\forall i,s$ : $\tilde{\hat{\mu}}^{i}_{t}[s]\leftarrow\frac{uncounted\_agents^{i}_{t}}{N\times|\mathcal{S}|}$
11:  $\forall i,s$ where $\hat{\upsilon}^{i}_{t}[s]$ is not $\emptyset$ : $\tilde{\hat{\mu}}^{i}_{t}[s]\leftarrow\tilde{\hat{\mu}}^{i}_{t}[s]+\frac{|\hat{\upsilon}^{i}_{t}[s]|}{N}$
12:  return  mean-field estimates $\forall i$ $\tilde{\hat{\mu}}^{i}_{t}$
The method for the general setting functions is given in Alg 2. In this setting, we assume that each agent is associated with a unique ID to avoid the same agents being counted multiple times. Each agent maintains a ‘count’ vector ${\hat{\upsilon}}^{i}_{t}$ of length $|\mathcal{S}|$ i.e. it is the same shape as the vector denoting the true empirical categorical distribution of agents. Each state position in the vector can hold a list of IDs. Before any actions are taken at each time step $t$, each agent’s count vector ${\hat{\upsilon}}^{i}_{t}$ is initialised as full of $\emptyset$ (‘no count’) markers for each state. Then, for each agent $j$ with which agent $i$ is connected via the observation graph, $i$ places $j$’s unique ID in its count vector in the correct state position. Next, for $C_{e}\geq 0$ communication rounds, agents exchange their local counts with neighbours on the communication network, and merge these counts with their own count vector, filtering out the unique IDs of those that have already been counted. If $C_{e}=0$ then the local count will remain purely independent. By exchanging these partially filled vectors, agents are able to improve their local counts by adding the states of agents that they have not been able to observe directly themselves.
Algorithm 3 Mean-field estimation and communication algorithm for environments with $\mathcal{G}^{vis}_{t}$
0:  Time-dependent visibility graph $\mathcal{G}^{vis}_{t}$, time-dependent communication graph $\mathcal{G}^{comm}_{t}$, states $\{s^{i}_{t}\}_{i=1}^{N}$, number of communication rounds $C_{e}$
1:  $\forall i,s:$ Initialise count vector $\hat{\upsilon}^{i}_{t}[s]$ with $\emptyset$
2:  $\forall i$ and $\forall s^{\prime}\in\mathcal{S^{\prime}}:(s^{i}_{t},s^{\prime})\in\mathcal{E}^{vis}_{t}$ : $\hat{\upsilon}^{i}_{t}[s^{\prime}]$leftarrow $\sum_{j\in 1,\dots,N:s^{j}_{t}=s^{\prime}}1$
3:  for $c_{e}$ in $1,\dots,C_{e}$ do
4:     $\forall i:$ Broadcast $\hat{\upsilon}^{i}_{t,c_{e}}$
5:     $\forall i:J^{i}_{t}=i\cup\{j\in\mathcal{N}:(i,j)\in\mathcal{E}^{comm}_{t}\}$
6:     $\forall i,s$ and $\forall j\in J^{i}_{t}:\hat{\upsilon}^{i}_{t,(c_{e}+1)}[s]\leftarrow\hat{\upsilon}^{j}_{t,c_{e}}[s]$ if $\hat{\upsilon}^{j}_{t,c_{e}}[s]\neq\emptyset$
7:  end for
8:  $\forall i:$ $counted\_agents^{i}_{t}\leftarrow\sum_{s\in\mathcal{S}:\hat{\upsilon}^{i}_{t}[s]\neq\emptyset}\hat{\upsilon}^{i}_{t}[s]$
9:  $\forall i:$ $uncounted\_agents^{i}_{t}\leftarrow N-counted\_agents^{i}_{t}$
10:  $\forall i:$ $unseen\_states^{i}_{t}\leftarrow\sum_{s\in\mathcal{S}:\hat{\upsilon}^{i}_{t}[s]=\emptyset}1$
11:  $\forall i,s$ where $\hat{\upsilon}^{i}_{t}[s]$ is not $\emptyset$ : $\tilde{\hat{\mu}}^{i}_{t}[s]\leftarrow\frac{\hat{\upsilon}^{i}_{t}[s]}{N}$
12:  $\forall i,s$ where $\hat{\upsilon}^{i}_{t}[s]$ is $\emptyset$ : $\tilde{\hat{\mu}}^{i}_{t}[s]\leftarrow\frac{uncounted\_agents^{i}_{t}}{N\times unobserved\_states^{i}_{t}}$
13:  return  mean-field estimates $\forall i$ $\tilde{\hat{\mu}}^{i}_{t}$
After the $C_{e}$ communication rounds, each state position ${\hat{\upsilon}}^{i}_{t}[s]$ either still maintains the $\emptyset$ marker if no agents have been counted in this state, or it contains $x_{s}>0$ unique IDs. The local mean-field estimate $\tilde{\hat{\mu}}^{i}_{t}$ is then obtained from ${\hat{\upsilon}}^{i}_{t}$ as follows. All states that have a count $x_{s}$ have this count converted into the categorical probability $x_{s}/N$ (we assume that agents know the total number of agents in the finite population, even if they cannot observe them all at each $t$). The total number of agents counted in ${\hat{\upsilon}}^{i}_{t}$ is given by $counted\_agents$ = $\sum_{s\in\mathcal{S}}x_{s}$, and the agents that have not been observed are $uncounted\_agents$ = $N$ $counted\_agents$. In this general setting, the unobserved agents are assumed to be uniformly distributed across all the states, so $uncounted\_agents/(N\times|\mathcal{S}|)$ is added to all the values in $\tilde{\hat{\mu}}^{i}_{t}$, replacing the $\emptyset$ marker for states for which no agents have been observed.
Algorithm for visibility-based environments
We now describe the differences in our estimation algorithm (Alg. 3) for the subclass of environments where $\mathcal{G}^{vis}_{t}$ applies in place of $\mathcal{G}^{obs}_{t}$, i.e. the mutual observability of agents depends in turn on the mutual visibility of states. Recall the benefit of $\mathcal{G}^{vis}_{t}$ over $\mathcal{G}^{obs}_{t}$ is that the former allows an agent in state $s$ to obtain a correct, complete count $x_{s^{\prime}}\geq 0$ of all the agents in state $s^{\prime}$, for any state $s^{\prime}$ that is visible to $s$ (note the count may be zero). Unique IDs are thus not required as there is no risk of counting the same agent twice when receiving communicated counts: either all agents in $s^{\prime}$ have been counted, or no count has yet been obtained for $s^{\prime}$. This simplifies the algorithm and helps preserve agent anonymity and privacy.
Secondly, uncounted agents cannot be in states for which a count has already been obtained, since the count is complete and correct, even if the count is $x_{s^{\prime}}=0$. Therefore after the $C_{e}$ communication rounds, the uncounted agents $uncounted/N$ need be uniformly distributed only across the positions in the vector that still have the $\emptyset$ marker, and not across all states as in the general setting. This makes the estimation more accurate in this special setting.
5 Experiments
Figure 1: ‘Target agreement’ task, population-independent policies, 50x50 grid.
Figure 2: ‘Cluster’ task, population-independent policies, 100x100 grid.
Figure 3: ‘Push object’ task, population-dependent policies with estimated mean-field distribution, 10x10 grid.
Figure 4: ‘Evade’ task, population-dependent policies with estimated mean-field distribution, 10x10 grid.
We provide two sets of experiments. The first set showcases that our function-approximation algorithm (Alg. 1) can scale to large state spaces for population-independent policies, and that in such settings networked, communicating agents can outperform purely-independent and even centralised agents, and do so by an even greater margin than in the tabular settings from Benjamin and Abate (2024). The second set demonstrates that Alg. 1 can handle population-dependent policies, as well as the ability of Alg. 3 to practically estimate the mean-field distribution locally.
For the types of game used in our demonstrations we follow the gold standard in prior works on MFGs, namely grid-world environments where agents can move in the four cardinal directions or remain in place (uz Zaman et al. 2023; Lauriere et al. 2022; Algumaei et al. 2023; Laurière 2021; Cui, Fabian, and Koeppl 2023; Lauriere et al. 2022; Benjamin and Abate 2024; Wu et al. 2024). We present results from four coordination tasks defined by the reward/transition functions of the agents - see Appx. A.1 for a full technical description. The first two tasks are those used with population-independent policies in Benjamin and Abate (2024), though while they show results for an 8x8 and a ‘larger’ 16x16 grid, our results for these tasks are for 100x100 and 50x50 grids:
  • Cluster. Agents are rewarded for gathering together. The agents are given no indication where they should cluster, agreeing this themselves over time.
  • Target agreement. Agents are rewarded for visiting any of a given number of targets, but their reward is proportional to the number of other agents co-located at the target. Agents must thus coordinate on which single target they will all meet at to maximise their individual rewards.
We also showcase the ability of our algorithm to handle two more complex tasks, using population-dependent policies and estimated mean-field observations:
  • Evade shark in shoal. At each $t$, a ‘shark’ in the environment takes a step towards the grid point containing the most agents according to the empirical mean-field distribution. The shark’s position forms part of agents’ local states in addition to their own position. Agents are rewarded more for being further from the shark, and also for clustering with other agents. As well as featuring a non-stationary distribution, we add ‘common noise’ to the environment, with the shark taking a random step with probability 0.01. Such noise that affects the local states of all agents in the same way, making the evolution of the distribution stochastic, makes population-independent policies sub-optimal (Laurière et al. 2022).
  • Push object to edge. Agents are rewarded for how close they are to an ‘object’ in the environment, and for how close this object is to the edge of the grid. The object’s position forms part of agents’ local states in addition to their own position. The object moves in a direction with a probability proportional to the number of agents on its opposite side, i.e. agents must coordinate on which side of the object from which to ‘push’ it, to ensure it moves toward the edge of the grid.
In these spatial environments, both the communication network $\mathcal{G}^{comm}_{t}$ and the visibility graph $\mathcal{G}^{vis}_{t}$ are determined by the physical distance from agent $i$; we show plots for various radii, expressed as fractions of the maximum possible distance (the diagonal length of the grid). To demonstrate the benefits of the networked architecture by comparison, we also plot the results of modified versions of our algorithm for centralised and independent learners. In the centralised setting, the Q-network updates of agent $i=1$ are automatically pushed to all other agents, and the true global mean-field distribution is always used in place of the local estimate i.e. $\tilde{\hat{\mu}}^{i}_{t}$ = ${\hat{\mu}}_{t}$. In the independent case, there are no links in $\mathcal{G}^{comm}_{t}$ or $\mathcal{G}^{vis}_{t}$, i.e. $\mathcal{E}^{comm}_{t}=\mathcal{E}^{vis}_{t}=\emptyset$. Hyperparameters are discussed in Appx. A.3.
5.1 Results
We evaluate our experiments via two metrics. Exploitability is the most common metric in works on MFGs, and is a measure of proximity to the MFNE. It quantifies how much a best-responding agent can benefit by deviating from the set of policies that generate the current mean-field distribution, with a decreasing exploitability meaning the population is closer to the MFNE. However, there are several issues with this metric in our setting, such that it may give limited or noisy information (see Appx. A.2 for a full discussion). We thus also give a second metric, as in Benjamin and Abate (2024): the population’s average discounted return. This allows us to compare how much agents are learning policies that increase their returns, even when the exploitability metric gives us limited ability to distinguish between the desirability of the MFNEs to which populations are converging.
Population-independent policies in large state-spaces
Figs. 1 and 2 illustrate that introducing function approximation to algorithms in this setting allows them to converge within a practical number of iterations ($k\ll 100$), even for large state spaces (100x100 grids). By contrast, the tabular algorithms in Benjamin and Abate (2024) appear only just to converge by $k=200$ for the same tasks for the larger of their two grids, which is only 16x16.
In Fig. 1 the networked agents generally have lower exploitability than both centralised and independent agents, while in Fig. 2 the networked agents’ exploitability is similar to that of the other cases. However, the networked agents significantly outperform the other architectures in terms of average return on both tasks, indicating that the communication scheme helps agents to find substantially ‘preferable’ equilibria. Moreover, the margin by which the networked agents can outperform the centralised agents is much greater than in Benjamin and Abate (2024), showing that the benefits of the communication scheme are even greater in non-tabular settings. See Appx. A for further experiments with these large state spaces.
Population-dependent policies in complex environments
Figs. 3 and 4, where agents estimate the mean-field distribution via Alg. 3, differ minimally from Figs. 5 and 6 in Appx. A, where agents directly receive the global mean-field distribution. This shows that our estimation algorithm allows agents to appropriately estimate the distribution, even with only one round of communication for agents to help each other improve their local counts. Only in the ‘push object’ task in Fig. 3, and there only with the smaller broadcast radii, do agents slightly underperform the returns of agents in the global observability case in Fig. 5, as is reasonable.
For the reasons given in Appx. A.2, the exploitability metric gives limited information in the ‘push object’ task in Fig. 3. In the ‘evade’ task in Fig. 4, exploitability suggests that centralised learners outperform the other cases. However, all of the networked cases significantly outperform the independent learners in terms of the average return to which they converge in both tasks. In the ‘push object’ task networked learners also significantly outperform centralised learners in all but the case with the smallest broad communication radius, while in the ‘evade’ task all networked cases perform similarly to the centralised case. Recall though that in the real world a centralised architecture is a strong assumption, a computational bottleneck and single point of failure.
Appendix A Experiments
For reproducibility, the code to run our experiments is provided with our Supplementary Materials, and will be made publicly available upon publication.
Experiments were conducted on a Linux-based machine with 2 x Intel Xeon Gold 6248 CPUs (40 physical cores, 80 threads total, 55 MiB L3 cache). We use the JAX framework to accelerate and vectorise our code. Random seeds are set in our code in a fixed way dependent on the trial number to allow easy replication of experiments - please see the code files in the Supplementary Materials for details.
A.1 Games
We conduct numerical tests with four games (defined by the agents’ reward/transition functions), chosen for being particularly amenable to intuitive and visualisable understanding of whether the agents are learning behaviours that are appropriate and explainable for the respective objective functions. In all cases, rewards are normalised in [0,1] after they are computed.
Cluster.
This is the inverse of the ‘exploration’ game in (Lauriere et al. 2022), where in our case agents are encouraged to gather together by the reward function $R(s^{i}_{t},a^{i}_{t},\hat{\mu}_{t})=$ log$(\hat{\mu}_{t}(s^{i}_{t}))$. That is, agent $i$ receives a reward that is logarithmically proportional to the fraction of the population that is co-located with it at time $t$. We give the population no indication where they should cluster, agreeing this themselves over time.
Agree on a single target.
Unlike in the above ‘cluster’ game, the agents are given options of locations at which to gather, and they must reach consensus among themselves. If the agents are co-located with one of a number of specified targets $\phi\in\Phi$ (in our experiments we place one target in each of the four corners of the grid), and other agents are also at that target, they get a reward proportional to the fraction of the population found there; otherwise they receive a penalty of -1. In other words, the agents must coordinate on which of a number of mutually beneficial points will be their single gathering place. Define the magnitude of the distances between $x,y$ at $t$ as $dist_{t}(x,y)$. The reward function is given by $R(s^{i}_{t},a^{i}_{t},\hat{\mu}_{t})=r_{targ}(r_{collab}(\hat{\mu}_{t}(s^{i}_{t})))$, where
$r_{targ}(x)=\begin{cases}x\quad&\text{if}\,\exists\phi\in\Phi\text{ s.t. dist}_{t}(s^{i}_{t},\phi)=0\\ -1\quad&\text{otherwise,}\\ \end{cases}$
$r_{collab}(x)=\begin{cases}x\quad&\text{if}\,\hat{\mu}_{t}(s^{i}_{t})>1/N\\ -1\quad&\text{otherwise.}\\ \end{cases}$
Evade shark in shoal.
Define the magnitude of the horizontal and vertical distances between $x,y$ at $t$ as $dist^{h}_{t}(x,y)$ and $dist^{v}_{t}(x,y)$ respectively. The state $s^{i}_{t}$ now consists of agent $i$’s position $x^{i}_{t}$ and the ‘shark’s’ position $\phi_{t}$. At each time step, the shark steps towards the most populated grid point $x^{*}_{t}=\arg\max_{x\in\mathcal{S}}\hat{\mu}_{t}(x)$. A horizontal step is taken if $dist^{h}_{t}(\phi_{t},x^{*}_{t})$geq $dist^{v}_{t}(\phi_{t},x^{*}_{t})$, otherwise a vertical step is taken. Additionally, the shark moves in a random direction with probability 0.01. Agents are rewarded more for being further from the shark, and also for clustering with other agents. The reward function is given by
$R(s^{i}_{t},a^{i}_{t},\hat{\mu}_{t})=\;$ $dist^{h}_{t}(\phi_{t},x^{i}_{t})+$
$dist^{v}_{t}(\phi_{t},x^{i}_{t})+\text{norm}_{dist}(\log(\hat{\mu}_{t}(x^{i}_{t}))),$
where $\text{norm}_{dist}(\cdot)$ indicates that the final term is normalised to have the same maximum and minimum values as the total combined vertical and horizontal distance.
Push object to edge.
[…]
$R(s^{i}_{t},a^{i}_{t},\hat{\mu}_{t})=\;$ $dist^{h}_{t}(\phi_{t},x^{i}_{t})+dist^{v}_{t}(\phi_{t},x^{i}_{t})+$
$dist^{h}_{t}(\phi_{t},\text{edge}^{*}_{t})+dist^{v}_{t}(\phi_{t},\text{edge}^{*}_{t}).$
Diffuse.
We conduct an additional experiment not mentioned in the main text, which is similar to the ‘exploration’ tasks in Lauriere et al. (2022), Wu et al. (2024) and other MFG works. In our version agents are rewarded for being located in more sparsely populated areas but only if they are stationary. The reward function is given by $R(s^{i}_{t},a^{i}_{t},\hat{\mu}_{t})=r_{stationary}(-\hat{\mu}_{t}(s^{i}_{t}))$, where
$r_{stationary}(x)=\begin{cases}x\quad&\text{if}\,a^{i}_{t}\text{ is `remain stationary'}\\ -1\quad&\text{otherwise.}\\ \end{cases}$
A.2 Experimental metrics
To give as informative results as possible about both performance and proximity to the MFNE, we provide two metrics for each experiment. Both metrics are plotted with mean and standard deviation, computed over the ten trials (each with a random seed) of the system evolution in each setting.
Exploitability
Works on MFGs most commonly use the exploitability metric to evaluate how close a given policy $\pi$ is to a NE policy $\pi^{*}$ (Lauriere et al. 2022; Perrin et al. 2020; Laurière et al. 2022; Algumaei et al. 2023; Pérolat et al. 2022b; Benjamin and Abate 2024; Wu et al. 2024). The metric usually assumes that all agents are following the same policy $\pi$, and quantifies how much an agent can benefit by deviating from $\pi$ by measuring the difference between the return given by $\pi$ and that of a $BR$ policy with respect to the distribution indicated by $\pi$:
Definition 9 (Exploitability of $\pi$)
The exploitability ${E}x$ of policy $\pi$ is given by:
${E}x(\pi)=V(BR(I(\pi)),I(\pi))-V(\pi,I(\pi)).$
If $\pi$ has a large exploitability then an agent can significantly improve its return by deviating from $\pi$, meaning that $\pi$ is far from $\pi^{*}$, whereas an exploitability of 0 implies that $\pi=\pi^{*}$. Prior works conducting empirical testing have generally focused on the centralised setting, so this classical definition, as well as most evaluations, only consider exploitability when all agents are following a single policy $\pi_{k}$. However, Benjamin and Abate (2024) notes that purely independent agents, as well as networked agents, may have divergent policies $\pi^{i}_{k}\neq\pi^{j}_{k}$ $\forall i,k\in 1,\dots,N$, as in our own setting. We therefore are interested in the ‘exploitability’ of the population’s joint policy $\boldsymbol{\pi}$ := ($\pi^{1},\dots,\pi^{N}$) $\in\Pi^{N}$.
Since we do not have access to the exact $BR$ policy as in some related works (Lauriere et al. 2022; Wu et al. 2024), we must instead approximate the exploitability, similarly to (Perrin et al. 2021; Benjamin and Abate 2024). We freeze the policy of all agents apart from a deviating agent, for which we store its current policy and then conduct 50 $k$ loops of policy improvement. To approximate the expectations in Def. 9, we take the best return of the deviating agent across 10 additional $k$ loops, as well as the mean of all the other agents’ returns across these same 10 loops. (While the policies of all non-deviating agents is $\pi_{k}$ in the centralised case, if the non-deviating agents do not share a single policy, then this method is in fact approximating the exploitability of their joint policy $\boldsymbol{\pi}^{-d}_{k}$, where $d$ is the deviating agent.) We then revert the agent back to its stored policy, before learning continues for all agents as per the main algorithm. Due to the expensive computations required for this metric, we evaluate it every second $k$ iteration of the main algorithm for Figs. 1, 2, 7, 8 and 9, and every fourth iteration for the population-dependent experiments.
The exploitability metric has a number of limitations in our setting. Our approximation takes place via MOMD policy improvement steps (as in the main algorithm) for a independent, deviating agent while the policies of the rest of the population are frozen. As such, the quality of our approximation is limited by the number of policy improvement/expectation rounds, which must be restricted for the sake of running speed of the experiments. Moreover, since one of the findings of our paper is that networked agents can improve their policies faster than independent or centralised agents, especially when non-linear function approximation is used, it is arguably unsurprising that approximating the $BR$ by an independently deviating agent sometimes gives an unclear and noisy metric. This includes the exploitability going below zero, which should not be possible if the policies and distributions are computed exactly.
Moreover, in coordination games (the setting on which we focus), agents benefit by following the same behaviour as others, and so a deviating agent generally stands to gain less from a $BR$ policy than it might in the non-coordination games on which many other works focus. For example, the return of a best-responding agent in the ‘push object’ game still depends on the extent to which other agents coordinate on which direction in which to push the box, meaning it cannot significantly increase its return by deviating. This means that the downward trajectory of the exploitability metric is less clear in our plots than in other works. This is likely why the approximated exploitability gets lower in the ‘disperse’ task in Fig. 8 than in the other tasks. Given the limitations presented by approximating exploitability, we also provide the second metric to indicate the progress of learning.
Average discounted return
We record the average discounted return of the agents’ policies $\pi_{k}^{i}$ during the $M$ iterations - this allows us to observe that settings that converge to similar exploitability values may not have similar average agent returns, suggesting that some algorithms are better than others at finding not just NE, but preferable NE. See for example Figs. 1 and 7, where the networked agents converge to similar exploitability as the independent and centralised agents, but receive higher average returns.
A.3 Hyperparameters
See Table 1 for our hyperparameter choices. We can group our hyperparameters into those controlling the size of the experiment, those controlling the size of the Q-network, those controlling the number of iterations of each loop in the algorithms and those affecting the learning/policy updates or policy adoption.
In our experiments we generally want to demonstrate that our communication-based algorithms outperform the centralised and independent architectures by allowing policies that are estimated to be better performing to proliferate through the population, such that convergence occurs within fewer iterations and computationally faster, even when the Q-function is poorly approximated and/or the mean-field is poorly estimated, as is likely to be the case in real-world scenarios. Moreover we want to show that there is a benefit even to a small amount of communication, so that communication rounds themselves do not excessively add to time complexity. As such, we generally select hyperparameters at the lowest end of those we tested during development, to show that our algorithms are particularly successful given what might otherwise be considered ‘undesirable’ hyperparameter choices.
Hyperparam. Value Comment
Trials 10 We run 10 trials with different random seeds for each experiment. We plot the mean and standard deviation for each metric across the trials.
Gridsize 10x10 / 50x50 / 100x100 Experiments with population-dependent policies are run on the 10x10 grid (Figs. 3, 4, 5 and 6), while experiments on large state spaces are run on 50x50 and 100x100 grids (Figs. 1, 2, 7, 8 and 9).
Population 500 We chose 500 for our demonstrations to show that our algorithm can handle large populations, indeed often larger than those demonstrated in other mean-field works, especially for grid-world environments, while also being feasible to simulate wrt. time and computation constraints (Yongacoglu, Arslan, and Yüksel 2022; Cui and Koeppl 2021; Cui, Fabian, and Koeppl 2023; Guo et al. 2020; Subramanian and Mahajan 2019; Yang et al. 2018; Subramanian et al. 2020, 2022, 2021; Wu et al. 2024; Benjamin and Abate 2024).
Number of neurons in input layer cf. comment The agent’s position is represented by two concatenated one-hot vectors indicating the agent’s row and column. An additional two such vectors are added for the shark’s/object’s position in the ‘evade’ and ’push object’ tasks. For population-dependent policies, the mean-field distribution is a flattened vector of the same size as the grid. As such, the input size in the ‘evade’ and ’push object’ tasks is $[(4\times\mathrm{dimension})+(\mathrm{dimension}^{2})]$; in the other settings it is $[2\times\mathrm{dimension}]$.
Neurons per hidden layer cf. comment We draw inspiration from common rules of thumb when selecting the number of neurons in hidden layers, e.g. it should be between the number of input neurons and output neurons / it should be 2/3 the size of the input layer plus the size of the output layer / it should be a power of 2 for computational efficiency. Using these rules of thumb as rough heuristics, we select the number of neurons per hidden layer by rounding the size of the input layer down to the nearest power of 2. The layers are all fully connected.
Hidden layers 2 We experimented with 2 and 3 hidden layers in the Q-networks. While 3 hidden layers gave similar or slighly better performance, we selected 2 for increased computational speed for conducting our experiments.
Activation function ReLU This is a common choice in deep RL.
$K$ 100 $K$ is chosen to be large enough to see at least one of the metrics converging.
$M$ 50 We tested $M$ in {50,100} and found that the lower value was sufficient to achieve convergence while minimising training time. It may be possible to converge with even smaller choices of $M$.
$L$ 50 We tested $L$ in {50,100} and found that the lower value was sufficient to achieve convergence while minimising training time. It may be possible to converge with even smaller choices of $L$.
$E$ 20 We tested $E$ in {20,50,100}, and choose the lowest value to show the benefit to convergence even from very few evaluation steps. It may be possible to reduce this value further and still achieve similar results.
$C_{p}$ 1 As in Benjamin and Abate (2024), we choose this value to show the convergence benefits brought by even a single communication round, even in networks that may have limited connectivity; higher choices are likely to have even better performance.
$C_{e}$ 1 Similar to $C_{p}$, we choose this value to show the ability of our algorithm to appropriately estimate the mean field even with only a single communication round, even in networks that may have limited connectivity.
$\gamma$ 0.9 Standard choice across RL literature.
$\tau_{q}$ 0.03 We tested $\tau_{q}$ in {0.01,0.02,0.03,0.04,0.05}, as well as linearly decreasing $\tau_{q}$ from 0.05 $\rightarrow$ 0 as $k$ increases. However, only 0.03 gave stable increase in return. Note that this is the value also chosen in Vieillard, Pietquin, and Geist (2020).
$|B|$ 32 This is a common choice of batch size that trades off noisy updates and computational efficiency.
$cl$1 We use the same value as in Vieillard, Pietquin, and Geist (2020).
$\nu$ $L-1$ We tested $\nu$ in $\{1,4,20,L-1\}$. We found that in our setting, updating $\theta^{\prime}\leftarrow\theta$ once per $k$ iteration s.t. $\theta^{\prime}_{k+1,l}=\theta_{k,l}$ $\forall l$ gave sufficient learning that was similar to the other potential choices of $\nu$, so we do this for simplicity, rather than arbitrarily choosing a frequency to update $\theta^{\prime}$ during each $k$ loop. Setting the target to be the policy from the previous iteration is similar to the method in Lauriere et al. (2022). Whilst Wu et al. (2024) updates the target within the $L$ loops for stability, we do not find this to be a problem in our experiments.
Optimiser Adam As in Vieillard, Pietquin, and Geist (2020), we use the Adam optimiser with initial learning rate 0.01.
$\tau_{k}$ cf. comment $\tau_{k}$ increases linearly from 0.001 to 1 across the $K$ iterations. This is a simplification of the annealing scheme used in Benjamin and Abate (2024). Further optimising the annealing process may lead to better results.
Table 1: Hyperparameters
A.4 Additional experiments
Figure 5: ‘Push object’ task, population-dependent policies with global observability of mean-field distribution, 10x10 grid.
Figure 6: ‘Evade’ task, population-dependent policies with global observability of mean-field distribution, 10x10 grid.
Figure 7: ‘Target agreement’ task, population-independent policies, 100x100 grid.
Figure 8: ‘Disperse’ task, population-independent policies, 100x100 grid.
Figure 9: ‘Cluster’ task, population-independent policies, 50x50 grid.
We provide additional experiments on large grids in Figs. 7, 8 and 9. In the ‘target agreement’ task in Fig. 7, the networked agents all significantly outperform both centralised and independent agents in term of average return, despite the independent agents appearing to have slightly lower exploitability. This is because independent agents are hardly improving their policies at all, such that there is little a deviating agent can do to increase its return in this coordination game, such that exploitability appears low, despite this being an undesirable equilibrium (see Appx. A.2 for further discussion on the limited information provided by the exploitation metric). In the ‘cluster’ task in Fig. 9, the networked agents obtain significantly higher return than the independent agents. While centralised agents have the lowest exploitability, networked agents of almost all communication radii outperform them in terms of average return, indicating that the communication scheme helps populations reach better performing equilibria.
In the additional ‘disperse’ task in Fig. 8, networked agents significantly outperform independent and centralised agents in terms of average return. They also outperform centralised agents in terms of exploitability, and significantly outperform independent agents in terms of exploitability.
Explanation of better performance by networked agents than centralised agents
By adopting policy parameters $\pi^{j}_{k+1}$ from neighbours with probability related to their estimated value $\sigma^{j}_{k+1}$, networked agents can adopt parameters that are estimated to be better performing than their own, which intuitively means networked agents improve their return quicker than agents learning entirely independently. Moreover, since the centralised case involves a single central agent training its Q-network and then pushing its update automatically to all the others, this single set of updated parameters will be taken on by all agents whether or not they are actually particularly well performing. Since the networked agents are more likely to select better performing parameters out of a variety of options according to our algorithm, we find networked agents usually in fact learn quicker than centralised agents, and indeed the benefit is even greater in the function approximation setting than in the tabular case given in Benjamin and Abate (2024). It is significant that our communication scheme not only allows us to avoid the undesirable assumption of a centralised learner, but even to outperform it.
Appendix B Related work
MFGs are a quickly growing research area, so we only discuss the works most closely related to this present work, and instead refer the reader to Benjamin and Abate (2024) for detailed discussion around the setting of networked communication for MFGs, and to Laurière et al. (2022) for a broader survey of MFGs. Our work is most closely related to Benjamin and Abate (2024), which introduced networked communication to the infinite-horizon MFG setting. However, this work focuses only on tabular settings rather than using function approximation as in ours, and only addresses population-independent policies.
Lauriere et al. (2022) uses Munchausen Online Mirror Descent (MOMD), similar to our method for learning with neural networks, but there are numerous differences to our setting: most relevantly, they study a finite-horizon episodic setting, where the mean-field distribution is updated in an exact way and an oracle supplies a centralised learner with rewards and transitions for it to learn a population-independent policy. Wu et al. (2024) uses MOMD to learn population-dependent policies, albeit also with a centralised method that exactly updates the mean-field distribution in a finite-horizon episodic setting. Perrin et al. (2022) learns population-dependent policies with function approximation in infinite-horizon settings like our own, but does so in a centralised, two-timescale manner without using the empirical mean-field distribution.
Yongacoglu, Arslan, and Yüksel (2022) addresses decentralised learning from a continuous, non-episodic run of the empirical system using either full or compressed information about the mean-field distribution, but agents are assumed to receive this information directly, rather than estimating it locally as in the algorithm we now present. They also do not consider function approximation or inter-agent communication in their algorithms. In the closely related but distinct area of Mean-Field RL, Subramanian et al. (2020) does estimate the empirical mean-field distribution from the local neighbourhood, however agents are seeking to estimate the mean action rather than the mean-field distribution over states as in our MFG setting. Their agents also do not have access to a communication network by which they can improve their estimates.
Appendix C Extensions and future work
Alg. 3, described in Sec. 4 assumes that if a state $s^{\prime}$ is connected to $s$ on the visibility graph $\mathcal{G}^{vis}_{t}$, an agent in $s$ is able to accurately count all the agents in $s^{\prime}$, i.e. it either counts the exact total or cannot observe the state at all. While we assume this for simplicity, this is not inherently the case, since a real-world agent may have only noisy observations even of others located nearby, due to imperfect sensors. We suggest two ways to deal with this case. Firstly, if agents share unique IDs as in Alg. 2, then when communicating their vectors of collected IDs with each other via $\mathcal{G}^{comm}_{t}$, agents would gain the most accurate picture possible of all the agents that have been observed in a given state. However, as we note in the main body of the paper, there are various reasons why sharing IDs might be undesirable, including privacy and scalability. If instead only counts are taken, and if the noise on each agents’ count is assumed to be independent and, for example, subject to a Gaussian distribution, the algorithm can easily be updated such that communicating agents compute averages of their local and received counts to improve their accuracy, rather than simply using communication to fill in counts for previously unobserved states. (Note that we can also consider the original case without noise to involve averaging, since averaging identical values equates to using the original value). Since the algorithm is intended to aid in local estimation of the mean-field distribution, which is inherently approximate due to the uniform method for distributing the uncounted agents, we are not concerned with reaching exact consensus between agents on the communicated counts, such that we do not require repeated averaging to ensure asymptotic convergence.
We may also wish to consider more sophisticated methods for distributing the uncounted agents across states, in place of the uniform distribution we use now. Such choices may be domain-specific based on knowledge of a particular environment. For example, one might in fact use the counts to perform Bayesian updates on a specific prior, where this prior may relate to the estimated mean-field distribution at the previous time step $t-1$. If agents seek to learn to predict the evolution of the mean field based on their own policy or by learning a model, the Bayesian prior may also be based on forward prediction from the estimated mean-field distribution at $t-1$. Future work lies in conducting experiments in all of these more general settings.
Perrin et al. (2022) notes that in grid-world settings such as those in our experiments, passing the (estimated or true global) mean-field distribution as a flat vector to the Q-network ignores the geometric structure of the problem. They therefore propose to create an embedding of the distribution by first passing the vector to a convolutional neural network, essentially treating the categorical distribution as an image. This technique is also followed in (Wu et al. 2024) (for their additional experiments, but not in the main body of their paper). As future work, we can test whether such a method improves the performance of our algorithms.
As in the closely-related work by Wu et al. (2024), we conduct extensive numerical experiments to demonstrate the benefits of our algorithm over baselines, and leave the theoretical analysis, such proof of convergence in the function approximation setting, for future work.
Appendix D Additional remarks
In our Algs. 2 and 3, agents share their local counts with neighbours on the communication network $\mathcal{G}^{comm}_{t}$, and only after the $C_{e}$ communication rounds do they complete their estimated distribution by distributing the uncounted agents along their vectors. An alternative would be for each agent to immediately form a local estimate from their local count obtained via $\mathcal{G}^{obs}_{t}$ or $\mathcal{G}^{vis}_{t}$, which is only then communicated and updated via the communication network. However, we take the former approach to avoid poor local estimations spreading through the network and leading to widespread inaccuracies. Information that is certain (the count) is spread as widely as possible, before being locally converted into an estimate of the total mean field. The same would be the case in our proposed extension for averaging noisy counts i.e. only the counts would be averaged, with the estimates completed by distributing the remaining agents after the $C_{e}$ communication rounds.

### v2

@@
 Abstract
-Recent works have provided algorithms by which decentralised agents, which may be connected via a communication network, can learn equilibria in Mean-Field Games from a single, non-episodic run of the empirical system.
-However, these algorithms are given for tabular settings: this computationally limits the size of players’ observation space, meaning that the algorithms are not able to handle anything but small state spaces, nor to generalise beyond policies depending on the ego player’s state to so-called ‘population-dependent’ policies.
+Recent algorithms allow decentralised agents, possibly connected via a communication network, to learn equilibria in Mean-Field Games from a non-episodic run of the empirical system.
+However, these algorithms are for tabular settings: this computationally limits the size of agents’ observation space, meaning the algorithms cannot handle anything but small state spaces, nor generalise beyond policies depending only on the agent’s local state to so-called ‘population-dependent’ policies.
 We address this limitation by introducing function approximation to the existing setting, drawing on the Munchausen Online Mirror Descent method that has previously been employed only in finite-horizon, episodic, centralised settings.
-While this permits us to include the population’s mean-field distribution in the observation for each player’s policy, it is arguably unrealistic to assume that decentralised agents would have access to this global information: we therefore additionally provide new algorithms that allow agents to estimate the global empirical distribution based on a local neighbourhood, and to improve this estimate via communication over a given network.
-Our experiments showcase how the communication network allows decentralised agents to estimate the mean-field distribution for population-dependent policies, and that exchanging policy information helps networked agents to outperform both independent and even centralised agents in function-approximation settings, by an even greater margin than in tabular settings.
+While this permits us to include the mean field in the observation for players’ policies, it is unrealistic to assume decentralised agents have access to this global information: we therefore also provide new algorithms allowing agents to locally estimate the global empirical distribution, and to improve this estimate via inter-agent communication.
+We show theoretically that exchanging policy information helps networked agents outperform both independent and even centralised agents in function-approximation settings.
+Our experiments demonstrate this happening empirically, by an even greater margin than in tabular settings, and show that the communication network allows decentralised agents to estimate the mean field for population-dependent policies.
 1 Introduction
-The mean-field game (MFG) framework (Lasry and Lions 2007; Huang, Malhamé, and Caines 2006) can be used to address the difficulty faced by multi-agent reinforcement learning (MARL) regarding computational scalability as the number of agents increases.
-It models a representative agent as interacting not with other individual agents in the population on a per-agent basis, but instead with a distribution over the other agents, known as the mean field.
+The mean-field game (MFG) framework Lasry and Lions (2007); Huang et al.
+(2006) can be used to address the difficulty faced by multi-agent reinforcement learning (MARL) regarding computational scalability as the number of agents grows Yardim and He (2024); Zeng et al.
+(2024).
+It models a representative agent as interacting not with other individual agents in the population on a per-agent basis, but instead with a distribution over the other agents, called the mean field.
 The MFG framework analyses the limiting case when the population consists of an infinite number of symmetric and anonymous agents, that is, they have identical reward and transition functions which depend on the mean-field distribution rather than on the actions of specific other players.
-The solution to this game is the mean-field Nash equilibrium (MFNE), which can be used as an approximation for the Nash equilibrium (NE) in a finite-agent game, with the error in the solution reducing as the number of agents N tends to infinity (Saldi, Başar, and Raginsky 2018; Anahtarci, Kariksiz, and Saldi 2023; Yardim, Goldman, and He 2024; Toumi, Malhame, and Le Ny 2024; Hu and Zhang 2024).
+The solution to this game is the mean-field Nash equilibrium (MFNE), which can be used as an approximation for the Nash equilibrium (NE) in a finite-agent game (which is harder to solve in itself), with the error in the solution reducing as the number of agents N tends to infinity Saldi et al.
+(2018); Anahtarci et al.
+(2023); Yardim et al.
+(2024); Toumi et al.
+(2024); Hu and Zhang (2024); Chen et al.
+(2024).
 MFGs have thus been applied to a wide variety of real-world problems: see Laurière et al.
-(2022) for applications and references.
-Recent works argue that classical algorithms for solving MFGs rely on assumptions and methods that are likely to be undesirable in real-world applications, emphasising that desirable qualities for practical MFG algorithms include: learning from the empirical distribution of $N$ agents (without generation or manipulation of this distribution by the algorithm itself or an external oracle/simulator); learning from a single, continued system run that is not arbitrarily reset as in episodic learning; model-free learning; decentralisation; and fast practical convergence (Yardim et al.
-2023; Benjamin and Abate 2024).
+(2022a) for examples.
+Recent works argue that classical algorithms for solving MFGs rely on assumptions and methods that are likely to be undesirable in real-world applications (e.g.
+swarm robotics or autonomous vehicles), emphasising that desirable qualities for practical MFG algorithms include: learning from the empirical distribution of $N$ agents (i.e.
+this distribution is generated only by the policies of the agents, rather than being updated by the algorithm itself or an external oracle/simulator); learning online from a single, non-episodic system run (i.e.
+similar to above, the population cannot be arbitrarily reset by an external controller); model-free learning; decentralisation; and fast practical convergence Yardim et al.
+(2023); Benjamin and Abate (2023).
 While these works address these desiderata, they do so only in settings in which the state and action spaces are small enough that the Q-function can be represented by a table, limiting their approaches’ scalability.
 Moreover, in those works, as in many others on MFGs, agents only observe their local state as input to their Q-function (which defines their policy).
-This is sufficient when the solved MFG is expected to have a stationary distribution (‘stationary MFGs’) (Laurière et al.
-2022; Xie et al.
-2021; Anahtarci, Kariksiz, and Saldi 2023; uz Zaman et al.
-2023; Yardim et al.
-2023; Benjamin and Abate 2024).
+This is sufficient when the solved MFG is expected to have a stationary distribution (‘stationary MFGs’) Laurière et al.
+(2022a); Xie et al.
+(2021); Anahtarci et al.
+(2023); Zaman et al.
+(2023); Yardim et al.
+(2023); Benjamin and Abate (2023).
 However, in reality there are numerous reasons for which agents may benefit from being able to respond to the current distribution.
 Recent work has thus increasingly focused on these more general settings where it is necessary for agents to have so-called ‘master policies’ (a.k.a.
-population-dependent policies) which depend on both the mean-field distribution and their local state (Cardaliaguet et al.
-2015; Carmona, Delarue, and Lacker 2016; Perrin et al.
-2022; Wu et al.
-2024; Laurière et al.
-2022; Lauriere et al.
-2022; Laurière et al.
-2022).
-The distribution is a high-dimensional (and thus large) observation object, and also takes a continuum of values.
+population-dependent policies) which depend on both the mean-field distribution and their local state Cardaliaguet et al.
+(2015); Carmona et al.
+(2016); Perrin et al.
+(2022); Wu et al.
+(2024); Laurière et al.
+(2022a, b, a).
+The distribution is a large, high-dimensional observation object, taking a continuum of values.
 Therefore a population-dependent Q-function cannot be represented exactly in a table and must be approximated.
-To address these limitations while maintaining the desiderata for real-world applications given in recent works, we introduce function approximation to the MFG setting of decentralised agents learning from a single, non-episodic run of the empirical system, allowing this setting to handle larger state spaces and to accept the mean-field distribution as an observation input.
-To overcome the difficulties of training non-linear approximators in this context, we use the so-called ‘Munchausen’ trick, introduced by Vieillard, Pietquin, and Geist (2020) for single-agent RL, and extended to MFGs by Lauriere et al.
-(2022), and to MFGs with population-dependent policies by Wu et al.
+To address these limitations while maintaining the desiderata for real-world applications given in recent works, we introduce function approximation to the MFG setting of decentralised agents learning online from a single, non-episodic run of the empirical system, allowing this setting to handle larger state spaces and to accept the mean-field distribution as an observation input.To overcome the difficulties of training non-linear approximators in this context, we use the so-called ‘Munchausen’ trick, introduced by Vieillard et al.
+(2020) for single-agent RL, and extended to MFGs by Laurière et al.
+(2022b), and to MFGs with population-dependent policies by Wu et al.
 (2024).
-We particularly explore this in the context of networked communication between decentralised agents (Benjamin and Abate 2024).
-We demonstrate that this communication brings two specific benefits over the purely independent setting (while also removing the undesirable assumption of a centralised learner, which in the real world may be unrealistic, a computational bottleneck and a vulnerable single point of failure).
+We particularly explore this in the context of networked communication between decentralised agents Benjamin and Abate (2023).
+We demonstrate that communication brings two specific benefits over the purely independent setting (while also removing the undesirable assumption of a centralised learner, which in the real world may be unrealistic, a computational bottleneck and a vulnerable single point of failure).
 Firstly, when the Q-function is approximated rather than exact, some updates lead to better performing policies than others.
-Allowing networked agents to propagate better performing policies through the population leads to faster practical convergence than in the purely independent-learning case, and often even than in the centralised case.
+Allowing networked agents to propagate better performing policies through the population (reminiscent of the use of fitness functions in distributed evolutionary algorithms Eiben and Smith (2015); Hart et al.
+(2015)) leads to faster learning than in the purely independent case and often even than in the centralised case, as we show both theoretically and empirically.
 Secondly, we argue that in the real world it is unrealistic to assume that decentralised agents, endowed with local state observations and limited (if any) communication radius, would be able to observe the global mean-field distribution and use it as input to their Q-functions / policy.
 We therefore further contribute two setting-dependent algorithms by which decentralised agents can estimate the global distribution from local observations, and further improve their estimates by communication with neighbours.
-We focus on ‘coordination games’, where agents can increase their individual rewards by following the same strategy as others and therefore have an incentive to communicate policies, even if the MFG setting itself is technically non-cooperative.
+We focus on ‘coordination’ games, where agents can increase their individual rewards by following the same strategy as others and therefore have an incentive to communicate policies, even if the MFG setting itself is technically non-cooperative.
 Thus our work can be applied to real-world problems in e.g.
 traffic signal control, formation control in swarm robotics, and consensus and synchronisation e.g.
-for sensor networks.
+for sensor networks Soleimani et al.
+(2024).
+[^1] We further preempt concerns about the appropriateness of communication in competitive settings by wondering whether self-interested agents would be any more likely to want to obey a central learner as has usually been assumed.
+Moreover we show that self-interested communicating agents can obtain higher returns than independent agents even in non-coordination games (Fig.
+5), indicating that they do have incentive to communicate.
 In summary, our contributions are:
   • We introduce, for the first time, function approximation to MFG settings with decentralised agents. To do this:
-    • We use Munchausen RL for the first time in an infinite-horizon MFG context, unlike the finite-horizon settings in Lauriere et al. (2022) and Wu et al. (2024).
-    • In doing so, we also present the first use of function approximation for solving MFGs from continued, non-episodic runs of the empirical system; this is unlike Yardim et al.
-(2023) and Benjamin and Abate (2024), which look at tabular settings.
+    • We use Munchausen RL for the first time in an infinite-horizon MFG context (for finite-horizon see Laurière et al. (2022b); Wu et al. (2024)).
+    • This constitutes the first use of function approximation for solving MFGs from a single, non-episodic run of the empirical system (for tabular settings see Yardim et al.
+(2023); Benjamin and Abate (2023)).
   • Function approximation allows us to explore larger state spaces, and also settings where agents’ policies depend on the mean-field distribution as well as their local state.
-  • Rather than assuming that agents have access to this global knowledge, as in prior works, we present two novel algorithms that allow decentralised agents to locally estimate the empirical distribution and to improve these estimates by communication with others.
-  • We present numerical results showcasing these algorithms, especially the two benefits of the decentralised communication scheme, which significantly outperforms both the independent and centralised settings.
+  • Rather than assuming that agents have access to this global knowledge as in prior works, we present two additional novel algorithms allowing decentralised agents to locally estimate the empirical distribution and to improve these estimates by inter-agent communication.
+  • We show theoretically that networked agents may learn faster than both centralised and independent agents in the function-approximation setting.
+  • We support this with extensive experiments, where our results showcase the two benefits of the decentralised communication scheme, which significantly outperforms both the independent and centralised settings.
 The paper is structured as follows.
-Preliminaries are given in Sec.
-2 and our core learning and policy improvement algorithm is presented in Sec.
+We give preliminaries in Sec.
+2 and our core learning and policy improvement algorithm in Sec.
 3.
-Our mean-field estimation and communication algorithms are given in Sec.
-4, and experiments are presented in Sec.
-5.
-See Appx.
-B for a more detailed ‘Related work’ section.
-Ways to extend our algorithms are discussed in the ‘Future Work’ section (Appx.
-C).
+We present our mean-field estimation and communication algorithms in Sec.
+4, theoretical results in Sec.
+5 and experiments in Sec.
+6.
+Sup.
+Mat.
+D contains an extended ‘Related work’ section.
 2 Preliminaries
@@
   • When the policy is population-dependent, if we assume perfect observability of the global mean-field distribution then we have $o^{i}_{t}=(s^{i}_{t},\hat{\mu}_{t})$.
-  • It is unrealistic to assume that decentralised agents with a possibly limited communication radius can observe the global mean-field distribution, so we allow agents to form a local estimate $\tilde{\hat{\mu}}^{i}_{t}$ which can be improved by communication with neighbours.
+  • It is unrealistic to assume that decentralised agents with a possibly limited communication radius can observe the global mean field, so we allow agents to form a local estimate $\tilde{\hat{\mu}}^{i}_{t}$ which can be improved by communication with neighbours.
 Here we have $o^{i}_{t}=(s^{i}_{t},\tilde{\hat{\mu}}^{i}_{t})$.
@@
 Thus the set of policies is $\Pi$ = $\pi$ : $\mathcal{S}\times\Delta_{\mathcal{S}}\rightarrow\Delta_{\mathcal{A}}$}, and the set of Q-functions is denoted $\mathcal{Q}=\{q:\mathcal{S}\times\Delta_{\mathcal{S}}\times\mathcal{A}\rightarrow\mathbb{R}\}$.
-Definition 1 (N-player symmetric anonymous games)
+Definition 1 (N-player symmetric anonymous games).
 […]
@@
 $\boldsymbol{\mu}=$$(\mu_{t})_{t\geq 0}$.
-Definition 2 (Induced mean-field flow)
+Definition 2 (Induced mean-field flow).
 We denote by $I(\pi)$ the mean-field flow $\boldsymbol{\mu}$ induced when all the agents follow $\pi$, where this is generated from $\pi$ as follows:
@@
 We also give the following definitions.
-Definition 3 (Mean-field discounted return)
+Definition 3 (Mean-field discounted return).
 In a MFG where all agents follow policy $\pi$ giving a mean-field flow $\boldsymbol{\mu}=$$(\mu_{t})_{t\geq 0}$, the expected discounted return of the representative agent is given by
 $V(\pi,\boldsymbol{\mu})=\mathbb{E}\left[\sum^{\infty}_{t=0}\gamma^{t}(R(s_{t},a_{t},\mu_{t}))\bigg{|}\begin{subarray}{c}s_{0}\sim\mu_{0}\\ a_{t}\sim\pi(\cdot|s_{t},\mu_{t})\\ s_{t+1}\sim P(\cdot|s_{t},a_{t},{\mu_{t}})\end{subarray}\right].$
-Definition 4 (Best-response (BR) policy)
+Definition 4 (Best-response (BR) policy).
 A policy $\pi^{*}$ is a best response (BR) against the mean-field flow $\boldsymbol{\mu}$ if it maximises the discounted return $V(\cdot,\boldsymbol{\mu})$; the set of these policies is denoted $BR(\boldsymbol{\mu})$, i.e.
 $\pi^{*}\in BR(\boldsymbol{\mu}):=\mathop{\arg\max}\limits_{\pi}V(\pi,\boldsymbol{\mu}).$
-Definition 5 (Mean-field Nash equilibrium (MFNE))
+Definition 5 (Mean-field Nash equilibrium (MFNE)).
 A pair ($\pi^{*},\boldsymbol{\mu}^{*})$ is a mean-field Nash equilibrium (MFNE) if the following two conditions hold:
-  • $\pi^{*}$ is a best response to $\boldsymbol{\mu}^{*}$, i.e. $\pi^{*}\in BR(\boldsymbol{\mu})$;
+  • $\pi^{*}$ is a best response to $\boldsymbol{\mu}^{*}$, i.e. $\pi^{*}\in BR(\boldsymbol{\mu}^{*})$;
   • $\boldsymbol{\mu}^{*}$ is induced by $\pi^{*}$, i.e. $\boldsymbol{\mu}^{*}$ = $I(\pi^{*})$.
@@
 If a population-dependent policy is a MFNE policy for any initial distribution $\mu_{0}$, it is a ‘master policy’.
-Previous works have shown that, in tabular settings, it is possible for a finite population of decentralised agents (each of which is permitted to have a distinct population-independent policy $\pi^{i}$) to learn the MFNE using only the empirical distribution $\hat{\mu}_{t}$, rather than the exactly calculated mean-field flow (Yardim et al.
-2023; Benjamin and Abate 2024).
-This MFNE is in turn an approximate NE for the original finite-agent game.
-In this work we provide algorithms to perform this in non-tabular and population-dependent settings, and demonstrate them empirically.
+Previous works have shown that, in tabular settings, it is possible for a finite population of decentralised agents (each of which is permitted to have a distinct population-independent policy $\pi^{i}$) to learn the MFNE using only the empirical distribution $\hat{\mu}_{t}$, rather than the exactly calculated infinite flow $\boldsymbol{\mu}$ Yardim et al.
+(2023); Benjamin and Abate (2023).
+This MFNE may be the goal in itself, or it can in turn serve as an approximate NE for the harder-to-solve game involving the finite population.
+In this work we provide algorithms to perform this process in non-tabular and population-dependent settings, and demonstrate them empirically.
 2.2 (Munchausen) Online Mirror Descent
@@
 $\pi_{k+1}(\cdot|s,\mu)=softmax\left(\frac{1}{\tau_{q}}\sum^{k}_{\kappa=0}{Q_{\kappa}}(s,\mu,\cdot)\right)$.
-$\tau_{q}$ is a temperature parameter that scales the entropy in Munchausen RL (Vieillard, Pietquin, and Geist 2020); note that this is a different temperature to the one agents use when selecting which communicated parameters to adopt, denoted $\tau^{comm}_{k}$ and discussed in Sec.
-3.2.
+$\tau_{q}$ is a temperature parameter that scales the entropy in Munchausen RL Vieillard et al.
+(2020); note that this is a different temperature to the one agents use when selecting which communicated parameters to adopt, denoted $\tau^{comm}_{k}$ (Sec.
+3.2).
 If the Q-function is approximated non-linearly using neural networks, it is difficult to compute this weighted sum.
-The so-called ‘Munchausen trick’ addresses this by computing a single Q-function that mimics the weighted sum using implicit regularisation based on the Kullback-Leibler (KL) divergence between $\pi_{k}$ and $\pi_{k+1}$ (Vieillard, Pietquin, and Geist 2020).
+The so-called ‘Munchausen trick’ addresses this by computing a single Q-function that mimics the weighted sum using implicit regularisation based on the Kullback-Leibler (KL) divergence between $\pi_{k}$ and $\pi_{k+1}$ Vieillard et al.
+(2020).
 Using this reparametrisation gives Munchausen OMD (MOMD), detailed further in Sec.
-3.1 (Lauriere et al.
-2022; Wu et al.
-2024).
-MOMD does not bias the MFNE, and has the same convergence guarantees as OMD (Hadikhanloo 2017; Pérolat et al.
-2022a; Wu et al.
-2024).
+3.1 Laurière et al.
+(2022b); Wu et al.
+(2024).
+MOMD does not bias the MFNE, and has the same convergence guarantees as OMD Hadikhanloo (2017); Perolat et al.
+(2021); Wu et al.
+(2024).
 2.3 Networks
 We conceive of the finite population as exhibiting two time-varying networks. The basic definition of such a network is:
-Definition 6 (Time-varying network)
+Definition 6 (Time-varying network).
 […]
-One of these networks $\mathcal{G}^{comm}_{t}$ defines which agents are able to communicate information to each other at time $t$.
-The second network $\mathcal{G}^{obs}_{t}$ is a graph defining which agents are able to observe each other’s states, which we use in general settings for estimating the mean-field distribution from local information.
+One of these networks $\mathcal{G}^{comm}_{t}$ defines which agents can communicate information to each other at time $t$.
+The second network $\mathcal{G}^{obs}_{t}$ is a graph defining which agents can observe each other’s states, which we use in general settings for estimating the mean-field distribution from local information.
 The structure of the two networks may be identical (e.g.
 if embodied agents can both observe the position (state) of, and exchange information with, other agents within a certain physical distance from themselves), or different (e.g.
-if agents can observe the positions of nearby agents, but only exchange information with agents by which they are connected via a telephone line, which may connect agents over long distances).
+if agents can observe the positions of nearby agents, but only exchange information with agents by which they are linked via satellite, which may connect agents over long distances).
 We also define an alternative version of the observation graph that is useful in a specific subclass of environments, which can most intuitively be thought of as those where agents’ states are positions in physical space.
@@
 We define this visibility graph as follows:
-Definition 7 (Time-varying state-visibility graph)
+Definition 7 (Time-varying state-visibility graph).
 […]
@@
 4).
-3 Learning and policy improvement
-3.1 Q-network and update
 Algorithm 1 Networked learning with non-linear function approximation
 0:  loop parameters $K,M,L,E,C_{p}$, learning parameters $\gamma,\tau_{q},|B|,cl,\nu$, $\{\tau^{comm}_{k}\}_{k\in\{0,\dots,K-1\}}$
-0:  initial states {$s^{i}_{0}$}i, $i=1,\ldots,N$; $t\leftarrow 0$
-1:  $\forall i$ : Randomly initialise parameters $\theta^{i}_{0}$ of Q-networks $\check{Q}_{\theta^{i}_{0}}(o^{i}_{t},\cdot)$, and set $\pi^{i}_{0}(a|o)=$ softmax$\left(\frac{1}{\tau_{q}}\check{Q}_{\theta^{i}_{0}}(o,\cdot)\right)(a)$
+0:  initial states {$s^{i}_{0}$}${}_{i=1}^{N}$; $t\leftarrow 0$
+1:  $\forall i$ : Randomly initialise parameters $\theta^{i}_{0}$ of Q-networks $\check{Q}_{\theta^{i}_{0}}(o,\cdot)$, and set $\pi^{i}_{0}(a|o)=$ softmax$\left(\frac{1}{\tau_{q}}\check{Q}_{\theta^{i}_{0}}(o,\cdot)\right)(a)$
 2:  for $k=0,\dots,K-1$ do
@@
 13:     $\check{Q}_{\theta^{i}_{k+1}}(o,\cdot)\leftarrow\check{Q}_{\theta^{i}_{k,L}}(o,\cdot)$
-14:     $\forall i$ : Set $\pi^{i}_{k+1}(a|o)=$ softmax$\left(\frac{1}{\tau_{q}}\check{Q}_{\theta^{i}_{k+1}}(o,\cdot)\right)(a)$
-15:     $\forall i:$ $\sigma^{i}_{k+1}=0$
+14:     $\forall i$ : $\pi^{i}_{k+1}(a|o)\leftarrow$ softmax$\left(\frac{1}{\tau_{q}}\check{Q}_{\theta^{i}_{k+1}}(o,\cdot)\right)(a)$
+15:     $\forall i:$ $\sigma^{i}_{k+1}\leftarrow 0$
 16:     for $e=0,\dots,E-1$ evaluation steps do
-17:        Take step $\forall i:a^{i}_{t}\sim\pi^{i}_{k}(\cdot|o^{i}_{t}),r^{i}_{t}=R(s^{i}_{t},a^{i}_{t},\hat{\mu}_{t}),s^{i}_{t+1}\sim P(\cdot|s^{i}_{t},a^{i}_{t},\hat{\mu}_{t})$; $t\leftarrow t+1$
-18:        $\forall i:$ $\sigma^{i}_{k+1}=\sigma^{i}_{k+1}+\gamma^{e}(r^{i}_{t}+h(\pi^{i}_{k+1}(s^{i}_{t})))$
+17:        Take step $\forall i:a^{i}_{t}\sim\pi^{i}_{k}(\cdot|o^{i}_{t}),r^{i}_{t}=R(s^{i}_{t},a^{i}_{t},\hat{\mu}_{t}),s^{i}_{t+1}\sim P(\cdot|s^{i}_{t},a^{i}_{t},\hat{\mu}_{t})$
+18:        $\forall i:$ $\sigma^{i}_{k+1}=\sigma^{i}_{k+1}+\gamma^{e}\cdot r^{i}_{t}$
 19:        $t\leftarrow t+1$
@@
 22:        $\forall i:$ Broadcast $\sigma^{i}_{k+1},\pi^{i}_{k+1}$
-23:        $\forall i:J^{i}_{t}\leftarrow\{j\in\mathcal{N}:(i,j)\in\mathcal{E}_{t}$}
+23:        $\forall i:J^{i}_{t}\leftarrow\{j\in\mathcal{N}:(i,j)\in\mathcal{E}^{comm}_{t}$}
 24:        $\forall i:$ Select $\mathrm{adopted}^{i}\sim$ Pr$\left(\mathrm{adopted}^{i}=j\right)$ = $\frac{\exp{(\sigma^{j}_{k+1}}/\tau^{comm}_{k})}{\sum_{x\in J^{i}_{t}}\exp{(\sigma^{x}_{k+1}}/\tau^{comm}_{k})}$ $\forall j\in J^{i}_{t}$
@@
 28:  end for
-29:  return  policies $\pi^{i}_{K}$}i, $i=1,\ldots,N$
+29:  return  policies $\pi^{i}_{K}$}${}_{i=1}^{N}$
+3 Learning and policy improvement
+3.1 Q-network and update
 Lines 1-14 of our novel Alg.
 1 contain the core Q-function/policy update method.
-Agent $i$ maintains a neural network parametrised by $\theta^{i}_{k}$ to approximate its Q-function: $\check{Q}_{\theta^{i}_{k}}(o,\cdot)$.
+Agent $i$ has a neural network parametrised by $\theta^{i}_{k}$ to approximate its Q-function: $\check{Q}_{\theta^{i}_{k}}(o,\cdot)$.
 The agent’s policy is given by $\pi_{\theta^{i}_{k}}(a|o)=$ softmax$\left(\frac{1}{\tau_{q}}\check{Q}_{\theta^{i}_{k}}(o,\cdot)\right)(a)$.
-When appropriate, we denote the policy $\pi^{i}_{k}(a|o)$ for simplicity.
-Each agent maintains a buffer (with size $M$) of collected transitions of the form $\left(o^{i}_{t},a_{t},r_{t},o^{i}_{t+1}\right)$.
+We denote the policy $\pi^{i}_{k}(a|o)$ for simplicity when appropriate.
+Each agent maintains a buffer (of size $M$) of collected transitions of the form $\left(o^{i}_{t},a^{i}_{t},r^{i}_{t},o^{i}_{t+1}\right)$.
 At each iteration $k$, they empty their buffer (Line 3) before collecting $M$ new transitions (Lines 4-7); each decentralised agent $i$ then trains its Q-network $\check{Q}_{\theta^{i}_{k}}$ via $L$ training updates as follows (Lines 8-12).
@@
 only every $\nu$ learning iterations (Line 11).
-At each iteration $l$, the agent samples a random batch $B^{i}_{k,l}$ of $|B|$ transitions from its buffer (Line 9).
-It then trains its neural network using stochastic gradient descent to minimise the following empirical loss (Line 10):
-Definition 8 (Empirical loss for Q-network)
-The empirical loss is given by
+At each iteration $l$, the agent samples a random batch $B^{i}_{k,l}$ of $|B|$ transitions from its buffer (Line 9), and trains its neural network via stochastic gradient descent to minimise the empirical loss (Def.
+8, Line 10).
+For $cl<0$, $[\cdot]^{0}_{cl}$ is a clipping function used in Munchausen RL to prevent numerical issues if the policy is too close to deterministic, as the log-policy term is otherwise unbounded Vieillard et al.
+(2020); Wu et al.
+(2024):
+Definition 8 (Empirical loss for Q-network).
+This is given by
 ^$\hat{\mathcal{L}}(\theta,\theta^{\prime})=\frac{1}{|B|}\sum_{transition\in B^{i}_{k,l}}\left|\check{Q}_{\theta^{i}_{k,l}}(o_{t},a_{t})-T\right|^{2},$
@@
 $\gamma\sum_{a\in\mathcal{A}}\pi_{\theta^{i,^{\prime}}_{k,l}}(a|o_{t+1})\left(\check{Q}_{\theta^{i,^{\prime}}_{k,l}}(o_{t+1},a)-\tau_{q}\ln\pi_{\theta^{i,^{\prime}}_{k,l}}(a|o_{t+1})\right).$
-For $cl<0$, $[\cdot]^{0}_{cl}$ is a clipping function used in Munchausen RL to prevent numerical issues if the policy is too close to deterministic, as the log-policy term is otherwise unbounded (Vieillard, Pietquin, and Geist 2020; Wu et al.
-2024).
 3.2 Communication and adoption of parameters
@@
 The other, described here, is used to privilege the spread of better performing policy updates through the population, allowing faster convergence in this networked case than in the independent and even centralised cases.
-We extend the work in Benjamin and Abate (2024) to the function-approximation case, where in our work agents broadcast the parameters of the Q-network that defines their policy, rather than the Q-function table.
-At each iteration $k$, after independently updating their Q-network and policy (Lines 3-14), agents approximately evaluate their new policies by collecting rewards for $E$ steps, and assign the estimated value to $\sigma^{i}_{k+1}$ (Lines 15-20).
+We adapt the work in Benjamin and Abate (2023) for the function-approximation case, where in our work agents broadcast the parameters of the Q-network that defines their policy, rather than the Q-function table.
+At each iteration $k$, after independently updating their Q-network and policy (Lines 3-14), agents approximate the infinite discounted return (Def.
+3) of their new policies by collecting rewards for $E$ steps, and assign the finite-step discounted sum to $\sigma^{i}_{k+1}$ (Lines 15-20).
 They then broadcast their Q-network parameters along with $\sigma^{i}_{k+1}$ (Line 22).
-Receiving these from their neighbours on the communication network, agents select which set of parameters to adopt by taking a softmax over their own and the received estimate values $\sigma^{j}_{k+1}$ (Lines 23-25).
-They repeat this broadcast and adoption process for $C_{p}$ rounds, which is distinct from the $C_{e}$ communication rounds discussed in Sec.
-4.
-By this process, decentralised agents can adopt policy parameters that are estimated to be better performing than their own, bringing the benefits to convergence speed discussed in Sec.
-5.1.
+Receiving these from neighbours on the network, agents select which set of parameters to adopt by taking a softmax over their own and the received estimate values $\sigma^{j}_{k+1}$ (Lines 23-25).
+They repeat the process for $C_{p}$ rounds.
+This allows decentralised agents to adopt policy parameters estimated to perform better than their own, accelerating learning as shown in Sec.
+5.
 4 Mean-field estimation and communication
@@
 to produce $o^{i}_{t}=(s^{i}_{t},\tilde{\hat{\mu}}^{i}_{t})$ for the steps $a^{i}_{t}\sim\pi^{i}_{k}(\cdot|o^{i}_{t})$ in Lines 5, 17 and 26.
-Both versions of the algorithm are subject to implicit assumptions, which we highlight and discuss methods for addressing in our ‘Future Work’ section in Appx.
-C.
-Algorithm for the general setting
-Algorithm 2 Mean-field estimation and communication algorithm in general settings
+Both versions of the algorithm are subject to implicit assumptions, which we highlight and discuss methods for addressing in Sup.
+Mat.
+E.
+4.0.1 Algorithm for the general setting
+Algorithm 2 Mean-field estimation and communication in general settings
 0:  Time-dependent observation graph $\mathcal{G}^{obs}_{t}$, time-dependent communication graph $\mathcal{G}^{comm}_{t}$, states $\{s^{i}_{t}\}_{i=1}^{N}$, number of communication rounds $C_{e}$
@@
 11:  $\forall i,s$ where $\hat{\upsilon}^{i}_{t}[s]$ is not $\emptyset$ : $\tilde{\hat{\mu}}^{i}_{t}[s]\leftarrow\tilde{\hat{\mu}}^{i}_{t}[s]+\frac{|\hat{\upsilon}^{i}_{t}[s]|}{N}$
-12:  return  mean-field estimates $\forall i$ $\tilde{\hat{\mu}}^{i}_{t}$
-The method for the general setting functions is given in Alg 2.
-In this setting, we assume that each agent is associated with a unique ID to avoid the same agents being counted multiple times.
+12:  return  mean-field estimates $\{\tilde{\hat{\mu}}^{i}_{t}\}_{i=1}^{N}$
+In this setting, our method (Alg.
+2) assumes each agent is associated with a unique ID to avoid the same agents being counted multiple times.
 Each agent maintains a ‘count’ vector ${\hat{\upsilon}}^{i}_{t}$ of length $|\mathcal{S}|$ i.e.
-it is the same shape as the vector denoting the true empirical categorical distribution of agents.
+of the same shape as the vector denoting the true empirical categorical distribution of agents.
 Each state position in the vector can hold a list of IDs.
@@
 By exchanging these partially filled vectors, agents are able to improve their local counts by adding the states of agents that they have not been able to observe directly themselves.
-Algorithm 3 Mean-field estimation and communication algorithm for environments with $\mathcal{G}^{vis}_{t}$
+After the $C_{e}$ communication rounds, each state position ${\hat{\upsilon}}^{i}_{t}[s]$ either still maintains the $\emptyset$ marker if no agents have been counted in this state, or contains $x_{s}>0$ unique IDs.
+The local mean-field estimate $\tilde{\hat{\mu}}^{i}_{t}$ is then obtained from ${\hat{\upsilon}}^{i}_{t}$ as follows.
+All states that have a count $x_{s}$ have this count converted into the categorical probability $x_{s}/N$ (we assume that agents know the total number of agents in the finite population, even if they cannot observe them all at each $t$).
+The total number of agents counted in ${\hat{\upsilon}}^{i}_{t}$ is given by $counted\_agents$ = $\sum_{s\in\mathcal{S}}x_{s}$, and the agents that have not been observed are $uncounted\_agents$ = $N$ $counted\_agents$.
+In this general setting, the unobserved agents are assumed to be uniformly distributed across all the states, so $uncounted\_agents/(N\times|\mathcal{S}|)$ is added to all the values in $\tilde{\hat{\mu}}^{i}_{t}$, replacing the $\emptyset$ marker for states for which no agents have been observed.
+4.0.2 Algorithm for visibility-based environments
+We give now the differences in our estimation algorithm (Alg.
+3, Sup.
+Mat.
+A) for the subclass of environments where $\mathcal{G}^{vis}_{t}$ applies in place of $\mathcal{G}^{obs}_{t}$, i.e.
+the mutual observability of agents depends in turn on the mutual visibility of states.
+The benefit of $\mathcal{G}^{vis}_{t}$ over $\mathcal{G}^{obs}_{t}$ is that the former allows an agent in state $s$ to obtain a correct, complete count $x_{s^{\prime}}\geq 0$ of all the agents in state $s^{\prime}$, for any state $s^{\prime}$ that is visible to $s$ (note the count may be zero).
+Unique IDs are thus not required as there is no risk of counting the same agent twice when receiving communicated counts: either all agents in $s^{\prime}$ have been counted, or no count has yet been obtained for $s^{\prime}$.
+This simplifies the algorithm and helps preserve agent privacy.
+Secondly, uncounted agents cannot be in states for which a count has already been obtained, since the count is complete and correct, even if the count is $x_{s^{\prime}}=0$.
+Therefore after the $C_{e}$ communication rounds, the $uncounted\_agents$ proportion needs to be uniformly distributed only across the positions in the vector that still have the $\emptyset$ marker, and not across all states as in the general setting.
+This makes the estimation more accurate in this special setting.
+5 Theoretical results
+To demonstrate the benefits of the networked architecture by comparison, we also consider (theoretically here and experimentally in Sec.
+6) the results of modified versions of our algorithm for centralised and independent learners.
+In the centralised setting, the Q-network updates of arbitrary agent $i=1$ are automatically pushed to all other agents, and the true global mean-field distribution is always used in place of the local estimate i.e.
+$\tilde{\hat{\mu}}^{i}_{t}$ = ${\hat{\mu}}_{t}$.
+In the independent case, there are no links in $\mathcal{G}^{comm}_{t}$ or $\mathcal{G}^{vis}_{t}$, i.e.
+$\mathcal{E}^{comm}_{t}=\mathcal{E}^{vis}_{t}=\emptyset$.
+Networked agents often learn faster than centralised ones in our experiments; we justify theoretically this possibly counterintuitive result here.
+We first make two strong assumptions that give conditions under which networked agents do outperform centralised ones.
+The fact that these strong assumptions do not always hold in reality explains why networked agents may not always outperform centralised ones.
+Recall that at each iteration $k$ of Alg.
+1, after independently updating their policies in Line 14, the population has the policies $\pi^{i}_{k+1}$}${}_{i=1}^{N}$.
+There is randomness in these independent policy updates, stemming from the random sampling of each agent’s independently collected buffer.
+In Lines 15-20, agents approximate the infinite discounted returns {$V(\pi^{i}_{k+1},I(\pi^{i}_{k+1}))\}_{i=1}^{N}$ (Def.
+3) of their updated policies by computing $\sigma^{i}_{k+1}$}${}_{i=1}^{N}$: the $E$-step discounted return with respect to the empirical mean field generated when agents follow policies $\pi^{i}_{k+1}$}${}_{i=1}^{N}$ (i.e.
+they do not at this stage all follow a single identical policy).
+We can characterise the approximation as $\sigma^{i}_{k+1}$}${}_{i=1}^{N}$ = $\widehat{V}(\pi^{i}_{k+1},I(\pi^{i}_{k+1}))\}_{i=1}^{N}$.
+Assumption 1.
+Assume that $\sigma^{i}_{k+1}$}${}_{i=1}^{N}$ are sufficiently good approximations so as to respect the ordering of the true values {$V(\pi^{i}_{k+1},I(\pi^{i}_{k+1}))\}_{i=1}^{N}$, i.e.
+$\forall i,j\in\{1,\dots,N\}$: $\sigma^{i}_{k+1}>\sigma^{j}_{k+1}\iff V(\pi^{i}_{k+1},I(\pi^{i}_{k+1}))>V(\pi^{j}_{k+1},I(\pi^{j}_{k+1})).$
+Assumption 2.
+Assume that after the $C_{p}$ rounds in Lines 21-27 in which agents exchange and adopt policies from neighbours, the population is left with a single policy such that $\forall i,j\in\{1,\dots,N\}$ $\pi^{i}_{k+1}=\pi^{j}_{k+1}$.
+[^2] Most simply we can think of Ass.
+2 holding if 1) $\tau^{comm}_{k}\rightarrow 0\;\forall k$ such that the softmax essentially becomes a max function, and 2) the communication network $\mathcal{G}^{comm}_{t}$ is static and connected during the $C_{p}$ communication rounds, where $C_{p}$ is larger than the network diameter $d_{\mathcal{G}^{comm}_{t}}$ Under these conditions, previous results on max-consensus algorithms show that all agents in the network will converge on the highest $\sigma^{max}_{k+1}$ value (and hence the unique associated $\pi^{max}_{k+1}$ within a number of rounds equal to the diameter $d_{\mathcal{G}^{comm}_{t}}$ Nejad et al.
+(2009); Benjamin and Abate (2023).
+However, policy consensus as in Ass.
+2 might be achieved even outside of these conditions, including if the network is dynamic and not connected at every step, given appropriate values for $C_{p}$ and $\tau^{comm}_{k+1}\in\mathbb{R}_{>0}$.
+Call the network consensus policy $\pi^{\mathrm{net}}_{k+1}$, and its associated finitely approximated return $\sigma^{\mathrm{net}}_{k+1}$.
+Recall that the centralised case is where the Q-network update of arbitrary agent $i=1$ is automatically pushed to all the others instead of the policy evaluation and exchange in Lines 15-27; this is equivalent to a networked case where policy consensus is reached on a random one of the policies $\pi^{i}_{k+1}$}${}_{i=1}^{N}$.
+Call this policy arbitrarily given to the whole population $\pi^{\mathrm{cent}}_{k+1}$, and its associated finitely approximated return $\sigma^{\mathrm{cent}}_{k+1}$.
+Theorem 1.
+Given Ass.
+1 and 2, $\mathbb{E}[V(\pi^{\mathrm{net}}_{k+1},I(\pi^{\mathrm{net}}_{k+1}))]>\mathbb{E}[V(\pi^{\mathrm{cent}}_{k+1},I(\pi^{\mathrm{cent}}_{k+1}))]$.
+Thus in expectation networked agents will increase their returns faster than centralised ones.
+Proof.
+Recall that before the communication rounds in Line 21 (Alg.
+1), the randomly updated policies $\pi^{i}_{k+1}$}${}_{i=1}^{N}$ have associated approximated returns $\sigma^{i}_{k+1}$}${}_{i=1}^{N}$.
+Denote the mean and maximum of this set $\sigma^{\mathrm{mean}}_{k+1}$ and $\sigma^{\mathrm{max}}_{k+1}$ respectively.
+Since $\pi^{\mathrm{cent}}_{k+1}$ is chosen arbitrarily from $\pi^{i}_{k+1}$}${}_{i=1}^{N}$, it will obey $\mathbb{E}[\sigma^{\mathrm{cent}}_{k+1}]=\sigma^{\mathrm{mean}}_{k+1}$ $\forall k$, though there will be high variance.
+Conversely, the softmax adoption probability (Line 24, Alg.
+1) for the networked case means by definition that policies with higher $\sigma^{i}_{k+1}$ are more likely to be adopted at each communication round.
+Thus the $\pi^{\mathrm{net}}_{k+1}$ that gets adopted by the whole networked population will obey $\mathbb{E}[\sigma^{\mathrm{net}}_{k+1}]>\sigma^{\mathrm{mean}}_{k+1}$ (if $\tau^{comm}_{k+1}\rightarrow 0$, it will obey $\mathbb{E}[\sigma^{\mathrm{net}}_{k+1}]=\sigma^{\mathrm{max}}_{k+1}$ $\forall k$).
+As such, $\mathbb{E}[\sigma^{\mathrm{net}}_{k+1}]>\mathbb{E}[\sigma^{\mathrm{cent}}_{k+1}]$, which by Ass.
+1 implies the result.
+∎
+The adoption scheme in Line 24 biases the spread of policies towards those estimated to be better, which, given sufficiently good approximations (Ass.
+1), results in higher discounted returns in practice.
+By choosing updates in a more principled way, networked agents learn faster than the centralised case that adopts updates regardless of quality.
+This intuition applies even if we loosen Ass.
+2 that the networked population converges on a single consensus policy within the $C_{p}$ communication rounds.
+Similar logic can also be applied to understand why networked agents outperform entirely independent ones, combined with the fact that divergence between policies in the independent case worsens sample complexity over the networked and centralised cases by biasing approximations of the Q-function Yardim et al.
+(2023); Benjamin and Abate (2023).
+Significantly, the communication scheme not only allows us to avoid the undesirable assumption of a central learner, but even to outperform it.
+Moreover, the benefit of networked communication over centralised learning is even greater in our function approximation setting than in the tabular case of Benjamin and Abate (2023), perhaps due to greater variance in the quality of Q-function estimates in our case.
+6 Experiments
+We provide two sets of experiments.
+The first set showcases that our function-approximation algorithm (Alg.
+1) can scale to large state spaces for population-independent policies, and that in such settings networked, communicating agents can outperform purely-independent and even centralised agents, and do so by an even greater margin than in the tabular settings from Benjamin and Abate (2023).
+The second set (given in Sup.
+Mat.
+C.4) demonstrates that Alg.
+1 can handle population-dependent policies, as well as the ability of Alg.
+3 to practically estimate the mean-field distribution locally.
+For the types of game used in our demonstrations we follow the gold standard in prior MFG works, i.e.
+grid-world environments where agents can move in the four cardinal directions or remain in place Laurière (2021); Laurière et al.
+(2022b); Zaman et al.
+(2023); Algumaei et al.
+(2023); Cui et al.
+(2023); Benjamin and Abate (2023); Wu et al.
+(2024).
+We present results from five tasks defined by the agents’ reward/transition functions, all but one of which are coordination tasks.
+The first two tasks are those used with population-independent policies in Benjamin and Abate (2023), but while they show results for an 8x8 and a ‘larger’ 16x16 grid, our results are for 100x100 and 50x50 grids:
+  • Cluster. Agents are rewarded for gathering but given no indication where to do so, agreeing it over time.
+  • Target agreement.
+Agents are rewarded for visiting any of a given number of targets, but the reward is proportional to the number of other agents co-located at the target.
+Agents must coordinate on which single target they will all meet at to maximise their individual rewards.
+See Sup.
+Mat.
+C.1 for a full technical description of these tasks, as well as descriptions of additional more complex tasks.
+In these spatial environments, both the communication network $\mathcal{G}^{comm}_{t}$ and the visibility graph $\mathcal{G}^{vis}_{t}$ are determined by the physical distance from agent $i$; we show plots for various radii, expressed as fractions of the maximum possible distance (the grid’s diagonal length).
+We discuss hyperparameters in Sup.
+Mat.
+C.3.
+Figure 1: ‘Target agreement’, pop.-independent, 100x100 grid.
+Figure 2: ‘Cluster’, pop.-independent, 100x100 grid.
+We evaluate our experiments via two metrics.
+Exploitability is the most common metric in works on MFGs, and is a measure of proximity to the MFNE.
+It quantifies how much a best-responding agent can benefit by deviating from the set of policies that generate the current mean-field distribution, with a decreasing exploitability meaning the population is closer to the MFNE.
+However, there are several issues with this metric in our setting, particularly for our coordination tasks where competitive agents benefit from aligning behaviours, such that it may give limited or noisy information (discussed further in Sup.
+Mat.
+C.2.1).
+We thus also give a second metric, as in Benjamin and Abate (2023): the population’s average discounted return.
+This allows us to compare how quickly agents are learning to increase their returns, even when ‘exploitability’ gives us limited ability to distinguish between the desirability of the MFNEs to which populations converge.
+6.1 Results and discussion
+Figs.
+1 and 2 illustrate that introducing function approximation to algorithms in this setting allows them to converge within a practical number of iterations ($k\ll 100$), even for large state spaces (100x100 grids).
+By contrast, the tabular algorithms in Benjamin and Abate (2023) appear only just to converge by $k=200$ for the same tasks for the larger of their two grids, which is only 16x16.
+In Figs.
+1 and 2, networked agents all significantly outperform both centralised and independent agents in term of average return, despite the centralised agents appearing to have similar exploitability, and the independent agents having similar or slightly lower exploitability.
+This is because independent agents (and also the centralised ones in Fig.
+1) hardly improve their policies at all, such that there is little a deviating agent can do to increase its return in these coordination games, meaning exploitability appears low, despite this being an undesirable equilibrium (see Sup.
+Mat.
+C.2 for further discussion on the limited information provided by the exploitation metric).
+The fact that the networked agents nevertheless significantly outperform the other architectures in terms of average return indicates that communication helps agents to find substantially ‘preferable’ equilibria.
+Moreover, the margin by which the networked agents can outperform the centralised agents is much greater than in Benjamin and Abate (2023), showing that the benefits of the communication scheme are even greater in non-tabular settings.
+See Sup.
+Mat.
+C.4 for further experiments with large state spaces, as well as others demonstrating the ability of our algorithm to handle two more complex tasks, using population-dependent policies and estimated mean-field observations.
+7 Conclusion
+We novelly contributed function approximation to the online-learning setting for empirical MFGs, and also contributed two novel algorithms for locally estimating the empirical mean field for population-dependent policies.
+We have justified theoretically why our networked communication algorithm is able to learn faster than both centralised and independent agents in this function approximation setting, and demonstrated empirically the ability of our algorithms to handle large state spaces and estimate the mean field.
+Limitations and ways to extend our algorithms are in Sup.
+Mat.
+E.
+Supplementary material to ‘Networked Communication for Mean-Field Games with Function Approximation and Empirical Mean-Field Estimation’
+Appendix A Mean-field estimation algorithm for visibility-based environments
+Algorithm 3 Mean-field estimation and communication for environments with $\mathcal{G}^{vis}_{t}$
 0:  Time-dependent visibility graph $\mathcal{G}^{vis}_{t}$, time-dependent communication graph $\mathcal{G}^{comm}_{t}$, states $\{s^{i}_{t}\}_{i=1}^{N}$, number of communication rounds $C_{e}$
@@
 12:  $\forall i,s$ where $\hat{\upsilon}^{i}_{t}[s]$ is $\emptyset$ : $\tilde{\hat{\mu}}^{i}_{t}[s]\leftarrow\frac{uncounted\_agents^{i}_{t}}{N\times unobserved\_states^{i}_{t}}$
-13:  return  mean-field estimates $\forall i$ $\tilde{\hat{\mu}}^{i}_{t}$
-After the $C_{e}$ communication rounds, each state position ${\hat{\upsilon}}^{i}_{t}[s]$ either still maintains the $\emptyset$ marker if no agents have been counted in this state, or it contains $x_{s}>0$ unique IDs.
-The local mean-field estimate $\tilde{\hat{\mu}}^{i}_{t}$ is then obtained from ${\hat{\upsilon}}^{i}_{t}$ as follows.
-All states that have a count $x_{s}$ have this count converted into the categorical probability $x_{s}/N$ (we assume that agents know the total number of agents in the finite population, even if they cannot observe them all at each $t$).
-The total number of agents counted in ${\hat{\upsilon}}^{i}_{t}$ is given by $counted\_agents$ = $\sum_{s\in\mathcal{S}}x_{s}$, and the agents that have not been observed are $uncounted\_agents$ = $N$ $counted\_agents$.
-In this general setting, the unobserved agents are assumed to be uniformly distributed across all the states, so $uncounted\_agents/(N\times|\mathcal{S}|)$ is added to all the values in $\tilde{\hat{\mu}}^{i}_{t}$, replacing the $\emptyset$ marker for states for which no agents have been observed.
-Algorithm for visibility-based environments
-We now describe the differences in our estimation algorithm (Alg.
-3) for the subclass of environments where $\mathcal{G}^{vis}_{t}$ applies in place of $\mathcal{G}^{obs}_{t}$, i.e.
-the mutual observability of agents depends in turn on the mutual visibility of states.
-Recall the benefit of $\mathcal{G}^{vis}_{t}$ over $\mathcal{G}^{obs}_{t}$ is that the former allows an agent in state $s$ to obtain a correct, complete count $x_{s^{\prime}}\geq 0$ of all the agents in state $s^{\prime}$, for any state $s^{\prime}$ that is visible to $s$ (note the count may be zero).
-Unique IDs are thus not required as there is no risk of counting the same agent twice when receiving communicated counts: either all agents in $s^{\prime}$ have been counted, or no count has yet been obtained for $s^{\prime}$.
-This simplifies the algorithm and helps preserve agent anonymity and privacy.
-Secondly, uncounted agents cannot be in states for which a count has already been obtained, since the count is complete and correct, even if the count is $x_{s^{\prime}}=0$.
-Therefore after the $C_{e}$ communication rounds, the uncounted agents $uncounted/N$ need be uniformly distributed only across the positions in the vector that still have the $\emptyset$ marker, and not across all states as in the general setting.
-This makes the estimation more accurate in this special setting.
-5 Experiments
-Figure 1: ‘Target agreement’ task, population-independent policies, 50x50 grid.
-Figure 2: ‘Cluster’ task, population-independent policies, 100x100 grid.
-Figure 3: ‘Push object’ task, population-dependent policies with estimated mean-field distribution, 10x10 grid.
-Figure 4: ‘Evade’ task, population-dependent policies with estimated mean-field distribution, 10x10 grid.
-We provide two sets of experiments.
-The first set showcases that our function-approximation algorithm (Alg.
-1) can scale to large state spaces for population-independent policies, and that in such settings networked, communicating agents can outperform purely-independent and even centralised agents, and do so by an even greater margin than in the tabular settings from Benjamin and Abate (2024).
-The second set demonstrates that Alg.
-1 can handle population-dependent policies, as well as the ability of Alg.
-3 to practically estimate the mean-field distribution locally.
-For the types of game used in our demonstrations we follow the gold standard in prior works on MFGs, namely grid-world environments where agents can move in the four cardinal directions or remain in place (uz Zaman et al.
-2023; Lauriere et al.
-2022; Algumaei et al.
-2023; Laurière 2021; Cui, Fabian, and Koeppl 2023; Lauriere et al.
-2022; Benjamin and Abate 2024; Wu et al.
-2024).
-We present results from four coordination tasks defined by the reward/transition functions of the agents - see Appx.
-A.1 for a full technical description.
-The first two tasks are those used with population-independent policies in Benjamin and Abate (2024), though while they show results for an 8x8 and a ‘larger’ 16x16 grid, our results for these tasks are for 100x100 and 50x50 grids:
-  • Cluster. Agents are rewarded for gathering together. The agents are given no indication where they should cluster, agreeing this themselves over time.
-  • Target agreement.
-Agents are rewarded for visiting any of a given number of targets, but their reward is proportional to the number of other agents co-located at the target.
-Agents must thus coordinate on which single target they will all meet at to maximise their individual rewards.
-We also showcase the ability of our algorithm to handle two more complex tasks, using population-dependent policies and estimated mean-field observations:
-  • Evade shark in shoal.
-At each $t$, a ‘shark’ in the environment takes a step towards the grid point containing the most agents according to the empirical mean-field distribution.
-The shark’s position forms part of agents’ local states in addition to their own position.
-Agents are rewarded more for being further from the shark, and also for clustering with other agents.
-As well as featuring a non-stationary distribution, we add ‘common noise’ to the environment, with the shark taking a random step with probability 0.01.
-Such noise that affects the local states of all agents in the same way, making the evolution of the distribution stochastic, makes population-independent policies sub-optimal (Laurière et al.
-2022).
-  • Push object to edge.
-Agents are rewarded for how close they are to an ‘object’ in the environment, and for how close this object is to the edge of the grid.
-The object’s position forms part of agents’ local states in addition to their own position.
-The object moves in a direction with a probability proportional to the number of agents on its opposite side, i.e.
-agents must coordinate on which side of the object from which to ‘push’ it, to ensure it moves toward the edge of the grid.
-In these spatial environments, both the communication network $\mathcal{G}^{comm}_{t}$ and the visibility graph $\mathcal{G}^{vis}_{t}$ are determined by the physical distance from agent $i$; we show plots for various radii, expressed as fractions of the maximum possible distance (the diagonal length of the grid).
-To demonstrate the benefits of the networked architecture by comparison, we also plot the results of modified versions of our algorithm for centralised and independent learners.
-In the centralised setting, the Q-network updates of agent $i=1$ are automatically pushed to all other agents, and the true global mean-field distribution is always used in place of the local estimate i.e.
-$\tilde{\hat{\mu}}^{i}_{t}$ = ${\hat{\mu}}_{t}$.
-In the independent case, there are no links in $\mathcal{G}^{comm}_{t}$ or $\mathcal{G}^{vis}_{t}$, i.e.
-$\mathcal{E}^{comm}_{t}=\mathcal{E}^{vis}_{t}=\emptyset$.
-Hyperparameters are discussed in Appx.
-A.3.
-5.1 Results
-We evaluate our experiments via two metrics.
-Exploitability is the most common metric in works on MFGs, and is a measure of proximity to the MFNE.
-It quantifies how much a best-responding agent can benefit by deviating from the set of policies that generate the current mean-field distribution, with a decreasing exploitability meaning the population is closer to the MFNE.
-However, there are several issues with this metric in our setting, such that it may give limited or noisy information (see Appx.
-A.2 for a full discussion).
-We thus also give a second metric, as in Benjamin and Abate (2024): the population’s average discounted return.
-This allows us to compare how much agents are learning policies that increase their returns, even when the exploitability metric gives us limited ability to distinguish between the desirability of the MFNEs to which populations are converging.
-Population-independent policies in large state-spaces
-Figs.
-1 and 2 illustrate that introducing function approximation to algorithms in this setting allows them to converge within a practical number of iterations ($k\ll 100$), even for large state spaces (100x100 grids).
-By contrast, the tabular algorithms in Benjamin and Abate (2024) appear only just to converge by $k=200$ for the same tasks for the larger of their two grids, which is only 16x16.
-In Fig.
-1 the networked agents generally have lower exploitability than both centralised and independent agents, while in Fig.
-2 the networked agents’ exploitability is similar to that of the other cases.
-However, the networked agents significantly outperform the other architectures in terms of average return on both tasks, indicating that the communication scheme helps agents to find substantially ‘preferable’ equilibria.
-Moreover, the margin by which the networked agents can outperform the centralised agents is much greater than in Benjamin and Abate (2024), showing that the benefits of the communication scheme are even greater in non-tabular settings.
-See Appx.
-A for further experiments with these large state spaces.
-Population-dependent policies in complex environments
-Figs.
-3 and 4, where agents estimate the mean-field distribution via Alg.
-3, differ minimally from Figs.
-5 and 6 in Appx.
-A, where agents directly receive the global mean-field distribution.
-This shows that our estimation algorithm allows agents to appropriately estimate the distribution, even with only one round of communication for agents to help each other improve their local counts.
-Only in the ‘push object’ task in Fig.
-3, and there only with the smaller broadcast radii, do agents slightly underperform the returns of agents in the global observability case in Fig.
-5, as is reasonable.
-For the reasons given in Appx.
-A.2, the exploitability metric gives limited information in the ‘push object’ task in Fig.
-3.
-In the ‘evade’ task in Fig.
-4, exploitability suggests that centralised learners outperform the other cases.
-However, all of the networked cases significantly outperform the independent learners in terms of the average return to which they converge in both tasks.
-In the ‘push object’ task networked learners also significantly outperform centralised learners in all but the case with the smallest broad communication radius, while in the ‘evade’ task all networked cases perform similarly to the centralised case.
-Recall though that in the real world a centralised architecture is a strong assumption, a computational bottleneck and single point of failure.
-Appendix A Experiments
-For reproducibility, the code to run our experiments is provided with our Supplementary Materials, and will be made publicly available upon publication.
+13:  return  mean-field estimates $\{\tilde{\hat{\mu}}^{i}_{t}\}_{i=1}^{N}$
+Appendix B Additional remarks on mean-field estimation algorithms
+In our Algs.
+2 and 3, agents share their local counts with neighbours on the communication network $\mathcal{G}^{comm}_{t}$, and only after the $C_{e}$ communication rounds do they complete their estimated distribution by distributing the uncounted agents along their vectors.
+An alternative would be for each agent to immediately form a local estimate from their local count obtained via $\mathcal{G}^{obs}_{t}$ or $\mathcal{G}^{vis}_{t}$, which is only then communicated and updated via the communication network.
+However, we take the former approach to avoid poor local estimations spreading through the network and leading to widespread inaccuracies.
+Information that is certain (the count) is spread as widely as possible, before being locally converted into an estimate of the total mean field.
+The same would be the case in our extension proposed in Sec.
+E for averaging noisy counts, i.e.
+only the counts would be averaged, with the estimates completed by distributing the remaining agents after the $C_{e}$ communication rounds.
+Appendix C Experiments
 Experiments were conducted on a Linux-based machine with 2 x Intel Xeon Gold 6248 CPUs (40 physical cores, 80 threads total, 55 MiB L3 cache).
 We use the JAX framework to accelerate and vectorise our code.
-Random seeds are set in our code in a fixed way dependent on the trial number to allow easy replication of experiments - please see the code files in the Supplementary Materials for details.
-A.1 Games
-We conduct numerical tests with four games (defined by the agents’ reward/transition functions), chosen for being particularly amenable to intuitive and visualisable understanding of whether the agents are learning behaviours that are appropriate and explainable for the respective objective functions.
+Random seeds are set in our code in a fixed way dependent on the trial number to allow easy replication of experiments.
+C.1 Games
+We conduct numerical tests with five games.
+All are defined by the agents’ reward/transition functions, and chosen for being particularly amenable to intuitive and visualisable understanding of whether the agents are learning behaviours that are appropriate and explainable for the respective objective functions.
 In all cases, rewards are normalised in [0,1] after they are computed.
 Cluster.
-This is the inverse of the ‘exploration’ game in (Lauriere et al.
-2022), where in our case agents are encouraged to gather together by the reward function $R(s^{i}_{t},a^{i}_{t},\hat{\mu}_{t})=$ log$(\hat{\mu}_{t}(s^{i}_{t}))$.
+This is the inverse of the ‘exploration’ game in Laurière et al.
+[2022b], where in our case agents are encouraged to gather together by the reward function $R(s^{i}_{t},a^{i}_{t},\hat{\mu}_{t})=$ log$(\hat{\mu}_{t}(s^{i}_{t}))$.
 That is, agent $i$ receives a reward that is logarithmically proportional to the fraction of the population that is co-located with it at time $t$.
@@
 Define the magnitude of the horizontal and vertical distances between $x,y$ at $t$ as $dist^{h}_{t}(x,y)$ and $dist^{v}_{t}(x,y)$ respectively.
-The state $s^{i}_{t}$ now consists of agent $i$’s position $x^{i}_{t}$ and the ‘shark’s’ position $\phi_{t}$.
-At each time step, the shark steps towards the most populated grid point $x^{*}_{t}=\arg\max_{x\in\mathcal{S}}\hat{\mu}_{t}(x)$.
+The state $s^{i}_{t}$ now consists of agent $i$’s position $x^{i}_{t}$ and a ‘shark’s’ position $\phi_{t}$.
+At each time step, the shark steps towards the most populated grid point according to the empirical mean-field distribution i.e.
+$x^{*}_{t}=\arg\max_{x\in\mathcal{S}}\hat{\mu}_{t}(x)$.
 A horizontal step is taken if $dist^{h}_{t}(\phi_{t},x^{*}_{t})$geq $dist^{v}_{t}(\phi_{t},x^{*}_{t})$, otherwise a vertical step is taken.
-Additionally, the shark moves in a random direction with probability 0.01.
+As well as featuring a non-stationary distribution, we add ‘common noise’ to the environment, with the shark in a random direction with probability 0.01.
+Such noise that affects the local states of all agents in the same way, making the evolution of the distribution stochastic, makes population-independent policies sub-optimal Laurière et al.
+[2022a].
 Agents are rewarded more for being further from the shark, and also for clustering with other agents.
@@
 $dist^{h}_{t}(\phi_{t},\text{edge}^{*}_{t})+dist^{v}_{t}(\phi_{t},\text{edge}^{*}_{t}).$
-Diffuse.
-We conduct an additional experiment not mentioned in the main text, which is similar to the ‘exploration’ tasks in Lauriere et al.
-(2022), Wu et al.
-(2024) and other MFG works.
+Disperse.
+This is similar to the ‘exploration’ tasks in Laurière et al.
+[2022b], Wu et al.
+[2024] and other MFG works.
 In our version agents are rewarded for being located in more sparsely populated areas but only if they are stationary.
@@
 $r_{stationary}(x)=\begin{cases}x\quad&\text{if}\,a^{i}_{t}\text{ is `remain stationary'}\\ -1\quad&\text{otherwise.}\\ \end{cases}$
-A.2 Experimental metrics
+C.2 Experimental metrics
 To give as informative results as possible about both performance and proximity to the MFNE, we provide two metrics for each experiment.
 Both metrics are plotted with mean and standard deviation, computed over the ten trials (each with a random seed) of the system evolution in each setting.
-Exploitability
-Works on MFGs most commonly use the exploitability metric to evaluate how close a given policy $\pi$ is to a NE policy $\pi^{*}$ (Lauriere et al.
-2022; Perrin et al.
-2020; Laurière et al.
-2022; Algumaei et al.
-2023; Pérolat et al.
-2022b; Benjamin and Abate 2024; Wu et al.
-2024).
-The metric usually assumes that all agents are following the same policy $\pi$, and quantifies how much an agent can benefit by deviating from $\pi$ by measuring the difference between the return given by $\pi$ and that of a $BR$ policy with respect to the distribution indicated by $\pi$:
-Definition 9 (Exploitability of $\pi$)
+C.2.1 Exploitability
+Works on MFGs most commonly use the exploitability metric to evaluate how close a given policy $\pi$ is to a NE policy $\pi^{*}$ Laurière et al.
+[2022b]; Perrin et al.
+[2020]; Laurière et al.
+[2022a]; Algumaei et al.
+[2023]; Pérolat et al.
+[2022]; Benjamin and Abate [2023]; Wu et al.
+[2024].
+The metric usually assumes that all agents are following the same policy $\pi$, and quantifies how much an agent can benefit by deviating from $\pi$ by measuring the difference between the return given by $\pi$ and that of a $BR$ policy with respect to the distribution generated by $\pi$:
+Definition 9 (Exploitability of $\pi$).
 The exploitability ${E}x$ of policy $\pi$ is given by:
@@
 Prior works conducting empirical testing have generally focused on the centralised setting, so this classical definition, as well as most evaluations, only consider exploitability when all agents are following a single policy $\pi_{k}$.
-However, Benjamin and Abate (2024) notes that purely independent agents, as well as networked agents, may have divergent policies $\pi^{i}_{k}\neq\pi^{j}_{k}$ $\forall i,k\in 1,\dots,N$, as in our own setting.
+However, Benjamin and Abate [2023] notes that purely independent agents, as well as networked agents, may have divergent policies $\pi^{i}_{k}\neq\pi^{j}_{k}$ $\forall i,k\in 1,\dots,N$, as in our own setting.
 We therefore are interested in the ‘exploitability’ of the population’s joint policy $\boldsymbol{\pi}$ := ($\pi^{1},\dots,\pi^{N}$) $\in\Pi^{N}$.
-Since we do not have access to the exact $BR$ policy as in some related works (Lauriere et al.
-2022; Wu et al.
-2024), we must instead approximate the exploitability, similarly to (Perrin et al.
-2021; Benjamin and Abate 2024).
+Since we do not have access to the exact $BR$ policy as in some related works Laurière et al.
+[2022b]; Wu et al.
+[2024], we must instead approximate the exploitability, similarly to Perrin et al.
+[2021]; Benjamin and Abate [2023].
 We freeze the policy of all agents apart from a deviating agent, for which we store its current policy and then conduct 50 $k$ loops of policy improvement.
@@
 Due to the expensive computations required for this metric, we evaluate it every second $k$ iteration of the main algorithm for Figs.
-1, 2, 7, 8 and 9, and every fourth iteration for the population-dependent experiments.
+1, 2, 3, 4 and 5, and every fourth iteration for the population-dependent experiments.
 The exploitability metric has a number of limitations in our setting.
-Our approximation takes place via MOMD policy improvement steps (as in the main algorithm) for a independent, deviating agent while the policies of the rest of the population are frozen.
+Our approximation takes place via MOMD policy improvement steps (as in the main algorithm) for an independent, deviating agent while the policies of the rest of the population are frozen.
 As such, the quality of our approximation is limited by the number of policy improvement/expectation rounds, which must be restricted for the sake of running speed of the experiments.
@@
 This includes the exploitability going below zero, which should not be possible if the policies and distributions are computed exactly.
-Moreover, in coordination games (the setting on which we focus), agents benefit by following the same behaviour as others, and so a deviating agent generally stands to gain less from a $BR$ policy than it might in the non-coordination games on which many other works focus.
+Moreover, in coordination games (the setting for all tasks apart from the ‘disperse’ game), agents benefit by following the same behaviour as others, and so a deviating agent generally stands to gain less from a $BR$ policy than it might in the non-coordination games on which many other works focus.
 For example, the return of a best-responding agent in the ‘push object’ game still depends on the extent to which other agents coordinate on which direction in which to push the box, meaning it cannot significantly increase its return by deviating.
 This means that the downward trajectory of the exploitability metric is less clear in our plots than in other works.
-This is likely why the approximated exploitability gets lower in the ‘disperse’ task in Fig.
+This is likely why the approximated exploitability gets lower in the non-coordination ‘disperse’ task in Fig.
 5 than in the other tasks.
 Given the limitations presented by approximating exploitability, we also provide the second metric to indicate the progress of learning.
-Average discounted return
+C.2.2 Average discounted return
 We record the average discounted return of the agents’ policies $\pi_{k}^{i}$ during the $M$ iterations - this allows us to observe that settings that converge to similar exploitability values may not have similar average agent returns, suggesting that some algorithms are better than others at finding not just NE, but preferable NE.
 See for example Figs.
-1 and 7, where the networked agents converge to similar exploitability as the independent and centralised agents, but receive higher average returns.
-A.3 Hyperparameters
+3 and 1, where the networked agents converge to similar exploitability as the independent and centralised agents, but receive higher average returns.
+C.3 Hyperparameters
+Figure 3: ‘Target agreement’ task, population-independent policies, 50x50 grid.
+Figure 4: ‘Cluster’ task, population-independent policies, 50x50 grid.
+Figure 5: ‘Disperse’ task, population-independent policies, 100x100 grid.
 See Table 1 for our hyperparameter choices.
@@
 Gridsize 10x10 / 50x50 / 100x100 Experiments with population-dependent policies are run on the 10x10 grid (Figs.
-3, 4, 5 and 6), while experiments on large state spaces are run on 50x50 and 100x100 grids (Figs.
-1, 2, 7, 8 and 9).
+6, 7, 8 and 9), while experiments on large state spaces are run on 50x50 and 100x100 grids (Figs.
+1, 2, 3, 4 and 5).
 Population 500 We chose 500 for our demonstrations to show that our algorithm can handle large populations, indeed often larger than those demonstrated in other mean-field works, especially for grid-world environments, while also being feasible to simulate wrt.
-time and computation constraints (Yongacoglu, Arslan, and Yüksel 2022; Cui and Koeppl 2021; Cui, Fabian, and Koeppl 2023; Guo et al.
-2020; Subramanian and Mahajan 2019; Yang et al.
-2018; Subramanian et al.
-2020, 2022, 2021; Wu et al.
-2024; Benjamin and Abate 2024).
+time and computation constraints Yongacoglu et al.
+[2022]; Cui and Koeppl [2021]; Cui et al.
+[2023]; Guo et al.
+[2023]; Subramanian and Mahajan [2019]; Yang et al.
+[2018]; Ganapathi Subramanian et al.
+[2021, 2020]; Subramanian et al.
+[2022]; Wu et al.
+[2024]; Benjamin and Abate [2023].
 Number of neurons in input layer cf.
@@
 It may be possible to reduce this value further and still achieve similar results.
-$C_{p}$ 1 As in Benjamin and Abate (2024), we choose this value to show the convergence benefits brought by even a single communication round, even in networks that may have limited connectivity; higher choices are likely to have even better performance.
+$C_{p}$ 1 As in Benjamin and Abate [2023], we choose this value to show the convergence benefits brought by even a single communication round, even in networks that may have limited connectivity; higher choices are likely to have even better performance.
 $C_{e}$ 1 Similar to $C_{p}$, we choose this value to show the ability of our algorithm to appropriately estimate the mean field even with only a single communication round, even in networks that may have limited connectivity.
@@
 However, only 0.03 gave stable increase in return.
-Note that this is the value also chosen in Vieillard, Pietquin, and Geist (2020).
+Note that this is the value also chosen in Vieillard et al.
+[2020].
 $|B|$ 32 This is a common choice of batch size that trades off noisy updates and computational efficiency.
-$cl$1 We use the same value as in Vieillard, Pietquin, and Geist (2020).
+$cl$1 We use the same value as in Vieillard et al. [2020].
 $\nu$ $L-1$ We tested $\nu$ in $\{1,4,20,L-1\}$.
@@
 $\theta^{\prime}_{k+1,l}=\theta_{k,l}$ $\forall l$ gave sufficient learning that was similar to the other potential choices of $\nu$, so we do this for simplicity, rather than arbitrarily choosing a frequency to update $\theta^{\prime}$ during each $k$ loop.
-Setting the target to be the policy from the previous iteration is similar to the method in Lauriere et al.
-(2022).
+Setting the target to be the policy from the previous iteration is similar to the method in Laurière et al.
+[2022b].
 Whilst Wu et al.
-(2024) updates the target within the $L$ loops for stability, we do not find this to be a problem in our experiments.
-Optimiser Adam As in Vieillard, Pietquin, and Geist (2020), we use the Adam optimiser with initial learning rate 0.01.
-$\tau_{k}$ cf.
-comment $\tau_{k}$ increases linearly from 0.001 to 1 across the $K$ iterations.
-This is a simplification of the annealing scheme used in Benjamin and Abate (2024).
+[2024] updates the target within the $L$ loops for stability, we do not find this to be a problem in our experiments.
+Optimiser Adam As in Vieillard et al. [2020], we use the Adam optimiser with initial learning rate 0.01.
+$\tau^{comm}_{k}$ cf.
+comment $\tau^{comm}_{k}$ increases linearly from 0.001 to 1 across the $K$ iterations.
+This is a simplification of the annealing scheme used in Benjamin and Abate [2023].
 Further optimising the annealing process may lead to better results.
 Table 1: Hyperparameters
-A.4 Additional experiments
-Figure 5: ‘Push object’ task, population-dependent policies with global observability of mean-field distribution, 10x10 grid.
-Figure 6: ‘Evade’ task, population-dependent policies with global observability of mean-field distribution, 10x10 grid.
-Figure 7: ‘Target agreement’ task, population-independent policies, 100x100 grid.
-Figure 8: ‘Disperse’ task, population-independent policies, 100x100 grid.
-Figure 9: ‘Cluster’ task, population-independent policies, 50x50 grid.
-We provide additional experiments on large grids in Figs.
-7, 8 and 9.
+C.4 Additional experiments
+C.4.1 Additional experiments on large grids
+We provide additional experiments on large grids in Figs. 3, 4 and 5.
 In the ‘target agreement’ task in Fig.
-7, the networked agents all significantly outperform both centralised and independent agents in term of average return, despite the independent agents appearing to have slightly lower exploitability.
-This is because independent agents are hardly improving their policies at all, such that there is little a deviating agent can do to increase its return in this coordination game, such that exploitability appears low, despite this being an undesirable equilibrium (see Appx.
-A.2 for further discussion on the limited information provided by the exploitation metric).
+3, the networked agents generally have lower exploitability than both centralised and independent agents, and significantly outperform the other architectures in terms of average return.
+As before, the margin by which the networked agents can outperform the centralised agents is much greater than in Benjamin and Abate [2023], showing that the benefits of the communication scheme are even greater in non-tabular settings.
 In the ‘cluster’ task in Fig.
-9, the networked agents obtain significantly higher return than the independent agents.
+4, the networked agents obtain significantly higher return than the independent agents.
 While centralised agents have the lowest exploitability, networked agents of almost all communication radii outperform them in terms of average return, indicating that the communication scheme helps populations reach better performing equilibria.
-In the additional ‘disperse’ task in Fig.
-8, networked agents significantly outperform independent and centralised agents in terms of average return.
+In the ‘disperse’ task in Fig.
+5, networked agents significantly outperform independent and centralised agents in terms of average return.
 They also outperform centralised agents in terms of exploitability, and significantly outperform independent agents in terms of exploitability.
-Explanation of better performance by networked agents than centralised agents
-By adopting policy parameters $\pi^{j}_{k+1}$ from neighbours with probability related to their estimated value $\sigma^{j}_{k+1}$, networked agents can adopt parameters that are estimated to be better performing than their own, which intuitively means networked agents improve their return quicker than agents learning entirely independently.
-Moreover, since the centralised case involves a single central agent training its Q-network and then pushing its update automatically to all the others, this single set of updated parameters will be taken on by all agents whether or not they are actually particularly well performing.
-Since the networked agents are more likely to select better performing parameters out of a variety of options according to our algorithm, we find networked agents usually in fact learn quicker than centralised agents, and indeed the benefit is even greater in the function approximation setting than in the tabular case given in Benjamin and Abate (2024).
-It is significant that our communication scheme not only allows us to avoid the undesirable assumption of a centralised learner, but even to outperform it.
-Appendix B Related work
-MFGs are a quickly growing research area, so we only discuss the works most closely related to this present work, and instead refer the reader to Benjamin and Abate (2024) for detailed discussion around the setting of networked communication for MFGs, and to Laurière et al.
-(2022) for a broader survey of MFGs.
-Our work is most closely related to Benjamin and Abate (2024), which introduced networked communication to the infinite-horizon MFG setting.
+The fact that this happens in this non-coordination, competitive game shows that agents do have an incentive to communicate with each other even if they are self-interested.
+C.4.2 Population-dependent policies in complex environments
+Figure 6: ‘Push object’ task, population-dependent policies with estimated mean-field distribution, 10x10 grid.
+Figure 7: ‘Evade’ task, population-dependent policies with estimated mean-field distribution, 10x10 grid.
+Figure 8: ‘Push object’ task, population-dependent policies with global observability of mean field, 10x10 grid.
+Figure 9: ‘Evade’ task, population-dependent policies with global observability of mean field, 10x10 grid.
+We also showcase the ability of our algorithm to handle two more complex tasks, using population-dependent policies and estimated mean-field observations.
+Figs.
+6 and 7, where agents estimate the mean-field distribution via Alg.
+3, differ minimally from Figs.
+8 and 9, where agents directly receive the global mean-field distribution.
+This shows that our estimation algorithm allows agents to appropriately estimate the distribution, even with only one round of communication for agents to help each other improve their local counts.
+Only in the ‘push object’ task in Fig.
+6, and there only with the smaller broadcast radii, do agents slightly underperform the returns of agents in the global observability case in Fig.
+8, as is reasonable.
+For the reasons given in Sup.
+Mat.
+C.2.1, the exploitability metric gives limited information in the ‘push object’ task in Fig.
+6.
+In the ‘evade’ task in Fig.
+7, exploitability suggests that centralised learners outperform the other cases.
+However, all of the networked cases significantly outperform the independent learners in terms of the average return to which they converge in both tasks.
+In the ‘push object’ task networked learners also significantly outperform centralised learners in all but the case with the smallest broad communication radius, while in the ‘evade’ task all networked cases perform similarly to the centralised case.
+Recall though that in the real world a centralised architecture is a strong assumption, a computational bottleneck and single point of failure.
+Appendix D Related work
+MFGs are a quickly growing research area, so we only discuss the works most closely related to this present work, and instead refer the reader to Benjamin and Abate [2023] for detailed discussion around the setting of networked communication for MFGs, and to Laurière et al.
+[2022a] for a broader survey of MFGs.
+Our work is most closely related to Benjamin and Abate [2023], which introduced networked communication to the infinite-horizon MFG setting.
 However, this work focuses only on tabular settings rather than using function approximation as in ours, and only addresses population-independent policies.
-Lauriere et al.
-(2022) uses Munchausen Online Mirror Descent (MOMD), similar to our method for learning with neural networks, but there are numerous differences to our setting: most relevantly, they study a finite-horizon episodic setting, where the mean-field distribution is updated in an exact way and an oracle supplies a centralised learner with rewards and transitions for it to learn a population-independent policy.
+Laurière et al.
+[2022b] uses Munchausen Online Mirror Descent (MOMD), similar to our method for learning with neural networks, but there are numerous differences to our setting: most relevantly, they study a finite-horizon episodic setting, where the mean-field distribution is updated in an exact way and an oracle supplies a centralised learner with rewards and transitions for it to learn a population-independent policy.
 Wu et al.
-(2024) uses MOMD to learn population-dependent policies, albeit also with a centralised method that exactly updates the mean-field distribution in a finite-horizon episodic setting.
+[2024] uses MOMD to learn population-dependent policies, albeit also with a centralised method that exactly updates the mean-field distribution in a finite-horizon episodic setting.
 Perrin et al.
-(2022) learns population-dependent policies with function approximation in infinite-horizon settings like our own, but does so in a centralised, two-timescale manner without using the empirical mean-field distribution.
-Yongacoglu, Arslan, and Yüksel (2022) addresses decentralised learning from a continuous, non-episodic run of the empirical system using either full or compressed information about the mean-field distribution, but agents are assumed to receive this information directly, rather than estimating it locally as in the algorithm we now present.
+[2022] learns population-dependent policies with function approximation in infinite-horizon settings like our own, but does so in a centralised, two-timescale manner without using the empirical mean-field distribution.
+Zhang et al.
+[2024] also uses function approximation along a non-episodic path, but involves a generic central agent learning an estimate of the mean field rather than using an empirical population.
+Approaches that directly update an estimate of the mean field must be able to generate rewards from this arbitrary mean field, even if they otherwise claim to be oracle-free.
+They are thus inherently centralised algorithms and rely on strong assumptions that may not apply in real-world problems.
+Conversely, we are interested in practical convergence in online, deployed settings, where the reward is computed from the empirical finite population.
+Yongacoglu et al.
+[2022] addresses decentralised learning from a continuous, non-episodic run of the empirical system using either full or compressed information about the mean-field distribution, but agents are assumed to receive this information directly, rather than estimating it locally as in the algorithm we now present.
 They also do not consider function approximation or inter-agent communication in their algorithms.
-In the closely related but distinct area of Mean-Field RL, Subramanian et al.
-(2020) does estimate the empirical mean-field distribution from the local neighbourhood, however agents are seeking to estimate the mean action rather than the mean-field distribution over states as in our MFG setting.
+In the closely related but distinct area of Mean-Field RL, Ganapathi Subramanian et al.
+[2021] does estimate the empirical mean-field distribution from the local neighbourhood, however agents are seeking to estimate the mean action rather than the mean-field distribution over states as in our MFG setting.
 Their agents also do not have access to a communication network by which they can improve their estimates.
-Appendix C Extensions and future work
+Appendix E Limitations and future work
+Our work follows the gold standard in MFGs by presenting experiments on grid world toy environments, albeit we show our algorithms are able to handle much larger and more complex games than prior work.
+Nevertheless future work lies in moving from these environments to real-world settings.
+In Sec.
+5 we give theoretical results showing that our networked algorithm can outperform a centralised alternative.
+We leave more general analysis, such proof of convergence and sample guarantees in the function approximation setting, for future work.
 Alg.
-3, described in Sec.
 3 assumes that if a state $s^{\prime}$ is connected to $s$ on the visibility graph $\mathcal{G}^{vis}_{t}$, an agent in $s$ is able to accurately count all the agents in $s^{\prime}$, i.e.
 it either counts the exact total or cannot observe the state at all.
-While we assume this for simplicity, this is not inherently the case, since a real-world agent may have only noisy observations even of others located nearby, due to imperfect sensors.
-We suggest two ways to deal with this case.
+We assume this for simplicity but it is not inherently the case, since a real-world agent may have only noisy observations even of others located nearby, due to imperfect sensors.
+We suggest two ways to deal with this.
 Firstly, if agents share unique IDs as in Alg.
 2, then when communicating their vectors of collected IDs with each other via $\mathcal{G}^{comm}_{t}$, agents would gain the most accurate picture possible of all the agents that have been observed in a given state.
-However, as we note in the main body of the paper, there are various reasons why sharing IDs might be undesirable, including privacy and scalability.
+However, as we note above, there are various reasons why sharing IDs might be undesirable, including privacy and scalability.
 If instead only counts are taken, and if the noise on each agents’ count is assumed to be independent and, for example, subject to a Gaussian distribution, the algorithm can easily be updated such that communicating agents compute averages of their local and received counts to improve their accuracy, rather than simply using communication to fill in counts for previously unobserved states.
 (Note that we can also consider the original case without noise to involve averaging, since averaging identical values equates to using the original value).
-Since the algorithm is intended to aid in local estimation of the mean-field distribution, which is inherently approximate due to the uniform method for distributing the uncounted agents, we are not concerned with reaching exact consensus between agents on the communicated counts, such that we do not require repeated averaging to ensure asymptotic convergence.
-We may also wish to consider more sophisticated methods for distributing the uncounted agents across states, in place of the uniform distribution we use now.
+Since the algorithm is intended to aid in local estimation of the mean-field distribution, which is inherently approximate due to the uniform method for distributing the uncounted agents, we are not concerned with reaching exact consensus between agents on the communicated counts, so we do not require repeated averaging to ensure asymptotic convergence.
+We may wish to consider more sophisticated methods for distributing the uncounted agents across states, in place of the current uniform distribution.
 Such choices may be domain-specific based on knowledge of a particular environment.
-For example, one might in fact use the counts to perform Bayesian updates on a specific prior, where this prior may relate to the estimated mean-field distribution at the previous time step $t-1$.
+For example, one might use the counts to perform Bayesian updates on a specific prior, where this prior may relate to the estimated mean-field distribution at the previous time step $t-1$.
 If agents seek to learn to predict the evolution of the mean field based on their own policy or by learning a model, the Bayesian prior may also be based on forward prediction from the estimated mean-field distribution at $t-1$.
@@
 Perrin et al.
-(2022) notes that in grid-world settings such as those in our experiments, passing the (estimated or true global) mean-field distribution as a flat vector to the Q-network ignores the geometric structure of the problem.
+[2022] notes that in grid-world settings such as those in our experiments, passing the (estimated or true global) mean-field distribution as a flat vector to the Q-network ignores the geometric structure of the problem.
 They therefore propose to create an embedding of the distribution by first passing the vector to a convolutional neural network, essentially treating the categorical distribution as an image.
-This technique is also followed in (Wu et al.
-2024) (for their additional experiments, but not in the main body of their paper).
+This technique is also followed in Wu et al.
+[2024] (for their additional experiments, but not in the main body of their paper).
 As future work, we can test whether such a method improves the performance of our algorithms.
-As in the closely-related work by Wu et al.
-(2024), we conduct extensive numerical experiments to demonstrate the benefits of our algorithm over baselines, and leave the theoretical analysis, such proof of convergence in the function approximation setting, for future work.
-Appendix D Additional remarks
-In our Algs.
-2 and 3, agents share their local counts with neighbours on the communication network $\mathcal{G}^{comm}_{t}$, and only after the $C_{e}$ communication rounds do they complete their estimated distribution by distributing the uncounted agents along their vectors.
-An alternative would be for each agent to immediately form a local estimate from their local count obtained via $\mathcal{G}^{obs}_{t}$ or $\mathcal{G}^{vis}_{t}$, which is only then communicated and updated via the communication network.
-However, we take the former approach to avoid poor local estimations spreading through the network and leading to widespread inaccuracies.
-Information that is certain (the count) is spread as widely as possible, before being locally converted into an estimate of the total mean field.
-The same would be the case in our proposed extension for averaging noisy counts i.e.
-only the counts would be averaged, with the estimates completed by distributing the remaining agents after the $C_{e}$ communication rounds.

### v3

@@
 Networked Communication for Mean-Field Games with Function Approximation and Empirical Mean-Field Estimation
-Abstract
+Abstract.
 Recent algorithms allow decentralised agents, possibly connected via a communication network, to learn equilibria in mean-field games from a non-episodic run of the empirical system.
@@
 While this permits us to include the mean field in the observation for players’ policies, it is unrealistic to assume decentralised agents have access to this global information: we therefore also provide new algorithms allowing agents to locally estimate the global empirical distribution, and to improve this estimate via inter-agent communication.
-We show theoretically that exchanging policy information helps networked agents outperform both independent and even centralised agents in function-approximation settings.
-Our experiments demonstrate this happening empirically, by an even greater margin than in tabular settings, and show that the communication network allows decentralised agents to estimate the mean field for population-dependent policies.
-1 Introduction
-The mean-field game (MFG) framework Lasry and Lions (2007); Huang et al.
-(2006) can be used to address the difficulty faced by multi-agent reinforcement learning (MARL) regarding computational scalability as the number of agents grows Yardim and He (2024); Zeng et al.
-(2024).
+We prove theoretically that exchanging policy information helps networked agents outperform both independent and even centralised agents in function-approximation settings.
+Our experiments demonstrate this happening empirically, and show that the communication network allows decentralised agents to estimate the mean field for population-dependent policies.
+Key words and phrases:
+Mean-Field Games, Communication Networks, Reinforcement Learning
+1. Introduction
+The mean-field game (MFG) framework (lasry2007mean; huangMFG) can be used to circumvent the difficulty faced by multi-agent reinforcement learning regarding computational scalability as the number of agents grows (yardim2024exploiting; zeng2024single).
+[^1] Preliminary versions of this paper have been accepted at: MARW at AAAI’25; GAAMAL at IJCAI’25; MALTA at AAAI’25; ALA at AAMAS’25; EMAS at AAMAS’25; GAIW at AAMAS’25; OptLearnMAS-25 at AAMAS’25; RLDM’25.
 It models a representative agent as interacting not with other individual agents in the population on a per-agent basis, but instead with a distribution over the other agents, called the mean field.
 The MFG framework analyses the limiting case when the population consists of an infinite number of symmetric and anonymous agents, that is, they have identical reward and transition functions which depend on the mean-field distribution rather than on the actions of specific other players.
-The solution to this game is the mean-field Nash equilibrium (MFNE), which can be used as an approximation for the Nash equilibrium (NE) in a finite-agent game (which is harder to solve in itself), with the error in the solution reducing as the number of agents N tends to infinity Saldi et al.
-(2018); Anahtarci et al.
-(2023); Yardim et al.
-(2024); Toumi et al.
-(2024); Hu and Zhang (2024); Chen et al.
-(2024).
-MFGs have thus been applied to a wide variety of real-world problems: see Laurière et al.
-(2022a) for examples.
+The solution to this game is the mean-field Nash equilibrium (MFNE), which can be used as an approximation for the Nash equilibrium (NE) in a finite-agent game (which is harder to solve in itself), with the error in the solution reducing as the number of agents N tends to infinity (finite_proofs; Anahtarci2020QLearningIR; yardim2024mean; TOUMI2024111420; hu2024mfoml; chen2024periodic).
+MFGs have thus been applied to a wide range of real-world problems: see survey_learningMFGs for examples.
 Recent works argue that classical algorithms for solving MFGs rely on assumptions and methods that are likely to be undesirable in real-world applications (e.g.
-swarm robotics or autonomous vehicles), emphasising that desirable qualities for practical MFG algorithms include: learning from the empirical distribution of $N$ agents (i.e.
+swarm robotics, autonomous vehicles), emphasising that desirable qualities for practical MFG algorithms include: learning from the empirical distribution of $N$ agents (i.e.
 this distribution is generated only by the policies of the agents, rather than being updated by the algorithm itself or an external oracle/simulator); learning online from a single, non-episodic system run (i.e.
-similar to above, the population cannot be arbitrarily reset by an external controller); model-free learning; decentralisation; and fast practical convergence Yardim et al.
-(2023); Benjamin and Abate (2023).
+similar to above, the population cannot be arbitrarily reset by an external controller); model-free learning; decentralised learning; and fast practical convergence (policy_mirror_independent; benjamin2024networked).
 While these works address these desiderata, they do so only in settings in which the state and action spaces are small enough that the Q-function can be represented by a table, limiting their approaches’ scalability.
 Moreover, in those works, as in many others on MFGs, agents only observe their local state as input to their Q-function (which defines their policy).
-This is sufficient when the solved MFG is expected to have a stationary distribution (‘stationary MFGs’) Laurière et al.
-(2022a); Xie et al.
-(2021); Anahtarci et al.
-(2023); Zaman et al.
-(2023); Yardim et al.
-(2023); Benjamin and Abate (2023).
-However, in reality there are numerous reasons for which agents may benefit from being able to respond to the current distribution.
+This is sufficient when the solved MFG is expected to have a stationary distribution (‘stationary MFGs’) (survey_learningMFGs; one_that_may_sample; Anahtarci2020QLearningIR; zaman2023oraclefree; policy_mirror_independent; benjamin2024networked).
+However, in reality there are numerous reasons why agents may benefit from being able to respond to the current distribution (discussed further in Appx.
+B).
 Recent work has thus increasingly focused on these more general settings where it is necessary for agents to have so-called ‘master policies’ (a.k.a.
-population-dependent policies) which depend on both the mean-field distribution and their local state Cardaliaguet et al.
-(2015); Carmona et al.
-(2016); Perrin et al.
-(2022); Wu et al.
-(2024); Laurière et al.
-(2022a, b, a).
+population-dependent policies) which depend on both the mean-field distribution and their local state (cardaliaguet2015masterequationconvergenceproblem; carmona_common; perrin2022generalization; wu2024populationaware; survey_learningMFGs; scalable_deep).
 The distribution is a large, high-dimensional observation object, taking a continuum of values.
-Therefore a population-dependent Q-function cannot be represented exactly in a table and must be approximated.
-To address these limitations while maintaining the desiderata for real-world applications given in recent works, we introduce function approximation to the MFG setting of decentralised agents learning online from a single, non-episodic run of the empirical system, allowing this setting to handle larger state spaces and to accept the mean-field distribution as an observation input.To overcome the difficulties of training non-linear approximators in this context, we use the so-called ‘Munchausen’ trick, introduced by Vieillard et al.
-(2020) for single-agent RL, and extended to MFGs by Laurière et al.
-(2022b), and to MFGs with population-dependent policies by Wu et al.
-(2024).
-We particularly explore this in the context of networked communication between decentralised agents Benjamin and Abate (2023).
-We demonstrate that communication brings two specific benefits over the purely independent setting (while also removing the undesirable assumption of a centralised learner, which in the real world may be unrealistic, a computational bottleneck and a vulnerable single point of failure).
+Therefore a population-dependent Q-function cannot be represented in a table and must be approximated.
+To address these limitations while maintaining the desiderata for real-world applications given in recent works, we introduce function approximation to the MFG setting of decentralised agents learning online from a single, non-episodic run of the empirical system, allowing this setting to handle larger state spaces and to accept the mean-field distribution as an observation input.
+To overcome the difficulties of training non-linear approximators in this context, we use the so-called ‘Munchausen’ trick, introduced first for single-agent RL (NEURIPS2020_2c6a0bae), extended to MFGs (scalable_deep), and then to MFGs with population-dependent policies (wu2024populationaware).
+We particularly explore this in the context of networked communication between decentralised agents.
+Almost all prior work relies on a centralised node to learn on behalf of all the agents.
+In this context ‘centralised’ does not imply global observability of the whole population’s actions - which would generally make computation infeasible given the complexity of the problem - but rather that learning is only conducted from the samples of a single representative agent, whose policy updates are assumed to be automatically pushed to the rest of the population by the central node.
+Therefore to reduce confusion, we sometimes refer to ‘central-agent learning’ instead of ‘centralised learning’ in contrast to prior works.
+More recent works have recognised that the assumption of a central learner might be be unrealistic in the real world, as well as a computational bottleneck and a vulnerable single point of failure (benjamin2024networked; policy_mirror_independent).
+We demonstrate that networked communication brings two specific benefits over the purely independent setting, while also removing the undesirable assumption of a central learner.
 Firstly, when the Q-function is approximated rather than exact, some updates lead to better performing policies than others.
-Allowing networked agents to propagate better performing policies through the population (reminiscent of the use of fitness functions in distributed evolutionary algorithms Eiben and Smith (2015); Hart et al.
-(2015)) leads to faster learning than in the purely independent case and often even than in the centralised case, as we show both theoretically and empirically.
+Allowing networked agents to propagate better performing policies through the population leads to faster learning than in the purely independent case and very often even than in the central-agent case, as we show both theoretically and empirically (this method is reminiscent of the use of fitness functions in distributed evolutionary algorithms (Eiben2015; survivability), and similarly of ‘population-based training’ (jaderberg2017population)).
 Secondly, we argue that in the real world it is unrealistic to assume that decentralised agents, endowed with local state observations and limited (if any) communication radius, would be able to observe the global mean-field distribution and use it as input to their Q-functions / policy.
@@
 Thus our work can be applied to real-world problems in e.g.
-traffic signal control, formation control in swarm robotics, and consensus and synchronisation e.g.
-for sensor networks Soleimani et al.
-(2024).
-[^1] We further preempt concerns about the appropriateness of communication in competitive settings by wondering whether self-interested agents would be any more likely to want to obey a central learner as has usually been assumed.
+traffic signal control, robotic swarm formation control, vehicle platooning, and consensus and synchronisation e.g.
+for sensor networks (10757965).
+[^2] We further preempt concerns about communication in competitive settings by wondering whether self-interested agents would be any more likely to want to obey a central learner as has usually been assumed.
 Moreover we show that self-interested communicating agents can obtain higher returns than independent agents even in non-coordination games (Fig.
-5), indicating that they do have incentive to communicate.
+6), indicating that they do have incentive to communicate.
 In summary, our contributions are:
   • We introduce, for the first time, function approximation to MFG settings with decentralised agents. To do this:
-    • We use Munchausen RL for the first time in an infinite-horizon MFG context (for finite-horizon see Laurière et al. (2022b); Wu et al. (2024)).
-    • This constitutes the first use of function approximation for solving MFGs from a single, non-episodic run of the empirical system (for tabular settings see Yardim et al.
-(2023); Benjamin and Abate (2023)).
+    • We use Munchausen RL for the first time in an infinite-horizon MFG context (cf. finite-horizon (scalable_deep; wu2024populationaware)).
+    • This constitutes the first use of function approximation for solving MFGs from a single, non-episodic run of the empirical system (cf.
+tabular settings (policy_mirror_independent; benjamin2024networked)).
   • Function approximation allows us to explore larger state spaces, and also settings where agents’ policies depend on the mean-field distribution as well as their local state.
-  • Rather than assuming that agents have access to this global knowledge as in prior works, we present two additional novel algorithms allowing decentralised agents to locally estimate the empirical distribution and to improve these estimates by inter-agent communication.
-  • We show theoretically that networked agents may learn faster than both centralised and independent agents in the function-approximation setting.
-  • We support this with extensive experiments, where our results showcase the two benefits of the decentralised communication scheme, which significantly outperforms both the independent and centralised settings.
+  • Instead of assuming that agents have access to this global knowledge as in prior works, we present two additional novel algorithms allowing decentralised agents to locally estimate the empirical distribution and to improve these estimates by inter-agent communication.
+  • We prove theoretically that networked agents can learn faster than even central-agent populations in the function-approximation setting.
+  • We support this with extensive experiments, where our results demonstrate the two benefits of the decentralised communication scheme, which significantly outperforms both the independent and central-agent architectures.
 The paper is structured as follows.
+Related work is given in Sec.
+2.
 We give preliminaries in Sec.
-2 and our core learning and policy improvement algorithm in Sec.
-3.
+3 and our core learning and policy-improvement algorithm in Sec.
+4.
 We present our mean-field estimation and communication algorithms in Sec.
-4, theoretical results in Sec.
+5, theoretical results in Sec.
 6 and experiments in Sec.
-6.
-Sup.
-Mat.
-D contains an extended ‘Related work’ section.
-2 Preliminaries
-2.1 Mean-field games
+7.
+2. Related work
+We refer the reader to benjamin2024networked for detailed discussion around the setting of networked communication in MFGs, and to survey_learningMFGs for a broader survey of MFGs.
+Our work is most closely related to benjamin2024networked, which introduced networked communication to the infinite-horizon MFG setting.
+However, this work focuses only on tabular settings rather than using function approximation as in ours, and only addresses population-independent policies.
+scalable_deep uses Munchausen Online Mirror Descent (MOMD), similar to our method for learning with neural networks, but in a different setting: their mean-field distribution is updated in an exact way and an oracle supplies a central learner with rewards and transitions for it to learn a population-independent policy, in a finite-horizon, episodic setting.
+wu2024populationaware uses MOMD to learn population-dependent policies, albeit also with a central-agent method that exactly updates the mean-field distribution in a finite-horizon episodic setting.
+perrin2022generalization learns population-dependent policies with function approximation in infinite-horizon settings like our own, but does so in a central-agent, two-timescale manner without using the empirical mean-field distribution.
+zhang2024stochastic also uses function approximation along a non-episodic path, but involves a generic central agent learning an estimate of the mean field rather than using an empirical population.
+Approaches that directly update an estimate of the mean field must be able to generate rewards from this arbitrary mean field, even if they otherwise claim to be oracle-free.
+They are thus inherently centralised algorithms and rely on strong assumptions that may not apply in real-world problems.
+Conversely, we are interested in practical convergence in online, deployed settings, where the reward is computed from the empirical finite population.
+subjective_equilibria addresses decentralised learning from a continuous, non-episodic run of the empirical system using either full or compressed information about the mean field, but agents are assumed to receive this information directly, rather than estimating it locally as in the algorithm we now present.
+They also do not consider function approximation or inter-agent communication.
+In the closely related but distinct area of mean-field RL, pomfrl does estimate the empirical mean-field distribution from the local neighbourhood, however agents are seeking to estimate the mean action rather than the mean-field distribution over states as in our MFG setting.
+Their agents also do not have access to a communication network by which they can improve their estimates.
+3. Preliminaries
+3.1. Mean-field games
 We use the following notation.
@@
 The set of probability measures on a finite set $\mathcal{X}$ is denoted $\Delta_{\mathcal{X}}$, and $\mathbf{e}_{x}\in\Delta_{\mathcal{X}}$ for $x\in\mathcal{X}$ is a one-hot vector with only the entry corresponding to $x$ set to 1, and all others set to 0.
-For time $t\geq 0$, $\hat{\mu}_{t}$ = $\frac{1}{N}\sum^{N}_{i=1}\sum_{s\in\mathcal{S}}$ $\mathbb{1}_{s^{i}_{t}=s}\mathbf{e}_{s}$in $\Delta_{\mathcal{S}}$ is a vector of length $|\mathcal{S}|$ denoting the empirical categorical state distribution of the $N$ agents at time $t$.
-For agent $i\in 1\dots N$, $i$’s policy at time $t$ depends on its observation $o^{i}_{t}$.
+For time $t\geq 0$, $\hat{\mu}_{t}$ = $\frac{1}{N}\sum^{N}_{i=1}\sum_{s\in\mathcal{S}}$ $\mathds{1}_{s^{i}_{t}=s}\mathbf{e}_{s}$ $\in$ $\Delta_{\mathcal{S}}$ is a vector of length $|\mathcal{S}|$ denoting the empirical categorical state distribution of the $N$ agents at time $t$.
+For agent $\in 1\dots N, ii$’s policy at time $t$ depends on its observation $o^{i}_{t}$.
 We explore three different forms that this observation object can take:
@@
 Thus the set of policies is $\Pi$ = $\pi$ : $\mathcal{S}\times\Delta_{\mathcal{S}}\rightarrow\Delta_{\mathcal{A}}$}, and the set of Q-functions is denoted $\mathcal{Q}=\{q:\mathcal{S}\times\Delta_{\mathcal{S}}\times\mathcal{A}\rightarrow\mathbb{R}\}$.
-Definition 1 (N-player symmetric anonymous games).
+Definition 3.1 (N-player symmetric anonymous games).
 […]
-$s^{i}_{t+1}\sim P(\cdot|s^{i}_{t},a^{i}_{t},\hat{\mu}_{t}),\;\;\;r^{i}_{t}=R(s^{i}_{t},a^{i}_{t},\hat{\mu}_{t}).$
+$s^{i}_{t+1}\sim P(\cdot|s^{i}_{t},a^{i}_{t},\hat{\mu}_{t}),\;\;\;r^{i}_{t}=R(s^{i}_{t},a^{i}_{t},\hat{\mu}_{t})\;\;\;\;\;\;\;\;\;\forall i=1,\dots,N.$
 At the limit as $N\rightarrow\infty$, the infinite population of agents can be characterised as a limit distribution $\mu\in\Delta_{\mathcal{S}}$; the infinite-agent game is termed an MFG.
@@
 $\boldsymbol{\mu}=$$(\mu_{t})_{t\geq 0}$.
-Definition 2 (Induced mean-field flow).
+Definition 3.2 (Induced mean-field flow).
 We denote by $I(\pi)$ the mean-field flow $\boldsymbol{\mu}$ induced when all the agents follow $\pi$, where this is generated from $\pi$ as follows:
@@
 We also give the following definitions.
-Definition 3 (Mean-field discounted return).
+Definition 3.3 (Mean-field discounted return).
 In a MFG where all agents follow policy $\pi$ giving a mean-field flow $\boldsymbol{\mu}=$$(\mu_{t})_{t\geq 0}$, the expected discounted return of the representative agent is given by
 $V(\pi,\boldsymbol{\mu})=\mathbb{E}\left[\sum^{\infty}_{t=0}\gamma^{t}(R(s_{t},a_{t},\mu_{t}))\bigg|\begin{subarray}{c}s_{0}\sim\mu_{0}\\ a_{t}\sim\pi(\cdot|s_{t},\mu_{t})\\ s_{t+1}\sim P(\cdot|s_{t},a_{t},{\mu_{t}})\end{subarray}\right].$
-Definition 4 (Best-response (BR) policy).
-A policy $\pi^{*}$ is a best response (BR) against the mean-field flow $\boldsymbol{\mu}$ if it maximises the discounted return $V(\cdot,\boldsymbol{\mu})$; the set of these policies is denoted $BR(\boldsymbol{\mu})$, i.e.
+Definition 3.4 (Best-response (BR) policy).
+A policy $\pi^{*}$ is a best response (BR) against the mean-field flow $\boldsymbol{\mu}$ if it maximises the discounted return $V(\cdot,\boldsymbol{\mu})$; the set of these policies is denoted $BR(\boldsymbol{\mu})$:
 $\pi^{*}\in BR(\boldsymbol{\mu}):=\mathop{\arg\max}\limits_{\pi}V(\pi,\boldsymbol{\mu}).$
-Definition 5 (Mean-field Nash equilibrium (MFNE)).
+Definition 3.5 (Mean-field Nash equilibrium (MFNE)).
 A pair ($\pi^{*},\boldsymbol{\mu}^{*})$ is a mean-field Nash equilibrium (MFNE) if the following two conditions hold:
@@
 If a population-dependent policy is a MFNE policy for any initial distribution $\mu_{0}$, it is a ‘master policy’.
-Previous works have shown that, in tabular settings, it is possible for a finite population of decentralised agents (each of which is permitted to have a distinct population-independent policy $\pi^{i}$) to learn the MFNE using only the empirical distribution $\hat{\mu}_{t}$, rather than the exactly calculated infinite flow $\boldsymbol{\mu}$ Yardim et al.
-(2023); Benjamin and Abate (2023).
+Previous works have shown that, in tabular settings, it is possible for a finite population of decentralised agents (each of which is permitted to have a distinct population-independent policy $\pi^{i}$) to learn the MFNE using only the empirical distribution $\hat{\mu}_{t}$, rather than the exactly calculated infinite flow $\boldsymbol{\mu}$ (policy_mirror_independent; benjamin2024networked).
 This MFNE may be the goal in itself, or it can in turn serve as an approximate NE for the harder-to-solve game involving the finite population.
 In this work we provide algorithms to perform this process in non-tabular and population-dependent settings, and demonstrate them empirically.
-2.2 (Munchausen) Online Mirror Descent
-Instead of performing the computationally expensive process of finding a $BR$ at each iteration, we can use a form of policy iteration for MFGs called Online Mirror Descent (OMD).
-This involves beginning with an initial policy $\pi_{0}$, and then at each iteration $k$, evaluating the current policy $\pi_{k}$ with respect to its induced mean-field flow $\boldsymbol{\mu}=I(\pi_{k})$ to compute its Q-function $Q_{k+1}$.
+3.2. (Munchausen) Online Mirror Descent
+Instead of finding a $BR$ at each iteration, which is computationally expensive, we can use a form of policy iteration for MFGs called Online Mirror Descent (OMD).
+This begins with an initial policy $\pi_{0}$, and then at each iteration $k$, evaluating the current policy $\pi_{k}$ with respect to its induced mean-field flow $\boldsymbol{\mu}=I(\pi_{k})$ to compute its Q-function $Q_{k+1}$.
 To stabilise the learning process, we then use a weighted sum over this and past Q-functions, and set $\pi_{k+1}$ to be the softmax over this weighted sum, i.e.
 $\pi_{k+1}(\cdot|s,\mu)=softmax\left(\frac{1}{\tau_{q}}\sum^{k}_{\kappa=0}{Q_{\kappa}}(s,\mu,\cdot)\right)$.
-$\tau_{q}$ is a temperature parameter that scales the entropy in Munchausen RL Vieillard et al.
-(2020); note that this is a different temperature to the one agents use when selecting which communicated parameters to adopt, denoted $\tau^{comm}_{k}$ (Sec.
-3.2).
+$\tau_{q}$ is a temperature parameter that scales the entropy in Munchausen RL (NEURIPS2020_2c6a0bae); note that this is a different temperature to the one agents use when selecting which communicated parameters to adopt, denoted $\tau^{comm}_{k}$ (Sec.
+4.2).
 If the Q-function is approximated non-linearly using neural networks, it is difficult to compute this weighted sum.
-The so-called ‘Munchausen trick’ addresses this by computing a single Q-function that mimics the weighted sum using implicit regularisation based on the Kullback-Leibler (KL) divergence between $\pi_{k}$ and $\pi_{k+1}$ Vieillard et al.
-(2020).
+The so-called ‘Munchausen trick’ addresses this by computing a single Q-function that mimics the weighted sum using implicit regularisation based on the Kullback-Leibler (KL) divergence between $\pi_{k}$ and $\pi_{k+1}$ (NEURIPS2020_2c6a0bae).
 Using this reparametrisation gives Munchausen OMD (MOMD), detailed further in Sec.
-3.1 Laurière et al.
-(2022b); Wu et al.
-(2024).
-MOMD does not bias the MFNE, and has the same convergence guarantees as OMD Hadikhanloo (2017); Perolat et al.
-(2021); Wu et al.
-(2024).
-2.3 Networks
+4.1 (scalable_deep; wu2024populationaware).
+MOMD does not bias the MFNE, and has the same convergence guarantees as OMD (hadikhanloo2017learninganonymousnonatomicgames; perolat2021scaling; wu2024populationaware).
+3.3. Networks
 We conceive of the finite population as exhibiting two time-varying networks. The basic definition of such a network is:
-Definition 6 (Time-varying network).
-[…]
+Definition 3.6 (Time-varying network).
+The network ($\mathcal{G}_{t}$)t≥0 is given by $\mathcal{G}_{t}$ = ($\mathcal{N},\mathcal{E}_{t}$), where $\mathcal{N}$ is the set of vertices representing agents $i\in\{1,\dots,N\}$, and the edge set $\mathcal{E}_{t}$ $\subseteq$ {(i,j) : i,j $\in$ $\mathcal{N}$, i $\neq$ j} is the set of undirected links present at time t.
+A network’s diameter $d_{\mathcal{G}_{t}}$ is the maximum of the shortest path length between any pair of nodes.
 One of these networks $\mathcal{G}^{comm}_{t}$ defines which agents can communicate information to each other at time $t$.
@@
 When this is the case, we usually think of agents’ ability to observe each other as depending more abstractly on whether states are visible to each other.
-We define this visibility graph as follows:
-Definition 7 (Time-varying state-visibility graph).
-[…]
-We view an agent in $s$ as able to obtain a count of the number of agents in $s^{\prime}$ if $s^{\prime}$ is visible to $s$.
+This visibility graph is:
+Definition 3.7 (Time-varying state-visibility graph).
+The state visibility graph ($\mathcal{G}^{vis}_{t}$)t≥0 is given by $\mathcal{G}^{vis}_{t}$ = ($\mathcal{S^{\prime}},\mathcal{E}^{vis}_{t}$), where $\mathcal{S^{\prime}}$ is the set of vertices representing the environment states $\mathcal{S}$, and the edge set $\mathcal{E}^{vis}_{t}$ $\subseteq$ {(m,n) : m,n $\in$ $\mathcal{S^{\prime}}$} is the set of undirected links present at time t, indicating which states are visible to each other.
+We say an agent in $s$ can obtain a count of the number of agents in $s^{\prime}$ if $s^{\prime}$ is visible to $s$.
 The benefit of this graph $\mathcal{G}^{vis}_{t}$ over $\mathcal{G}^{obs}_{t}$ is that there is mutual exclusivity: either an agent in state $s$ is able to obtain a total count of all of the agents in state $s^{\prime}$ (if $s^{\prime}$ is visible to $s$), or it cannot obtain information about any agent in state $s^{\prime}$ (if those states are not visible to each other).
@@
 These benefits are not available if the observability graph is defined strictly between agents as in $\mathcal{G}^{obs}_{t}$, such that using $\mathcal{G}^{vis}_{t}$ facilitates more efficient estimation of the global mean-field distribution from local information in settings where $\mathcal{G}^{vis}_{t}$ is applicable (see Sec.
-4).
-Algorithm 1 Networked learning with non-linear function approximation
-0:  loop parameters $K,M,L,E,C_{p}$, learning parameters $\gamma,\tau_{q},|B|,cl,\nu$, $\{\tau^{comm}_{k}\}_{k\in\{0,\dots,K-1\}}$
-0:  initial states {$s^{i}_{0}$}${}_{i=1}^{N}$; $t\leftarrow 0$
-1:  $\forall i$ : Randomly initialise parameters $\theta^{i}_{0}$ of Q-networks $\check{Q}_{\theta^{i}_{0}}(o,\cdot)$, and set $\pi^{i}_{0}(a|o)=$ softmax$\left(\frac{1}{\tau_{q}}\check{Q}_{\theta^{i}_{0}}(o,\cdot)\right)(a)$
+5).
+Algorithm 1 Networked learning with function approximation
+0: loop parameters $K,M,L,E,C_{p}$; learning parameters $\gamma,\tau_{q},|B|,cl,\nu$, $\{\tau^{comm}_{k}\}_{k\in\{0,\dots,K-1\}}$; initial states $s^{i}_{0}$$_{i=1}^{N}$; $t\leftarrow 0$
+[…]
 2:for $k=0,\dots,K-1$ do
@@
 9:   $\forall i:$ Sample batch $B^{i}_{k,l}$ from $i$’s buffer
-10:        Update $\theta$ to minimise ^$\hat{\mathcal{L}}(\theta,\theta^{\prime})$ as in Def. 8
+10:   Update $\theta$ to minimise $\hat{\mathcal{L}}(\theta,\theta^{\prime})$ as in Def. 4.1
 11:   If $l\mod$ $\nu=0$, set $\theta^{\prime}\leftarrow\theta$
@@
 13:  $\check{Q}_{\theta^{i}_{k+1}}(o,\cdot)\leftarrow\check{Q}_{\theta^{i}_{k,L}}(o,\cdot)$
-14:     $\forall i$ : $\pi^{i}_{k+1}(a|o)\leftarrow$ softmax$\left(\frac{1}{\tau_{q}}\check{Q}_{\theta^{i}_{k+1}}(o,\cdot)\right)(a)$
+[…]
 15:  $\forall i:$ $\sigma^{i}_{k+1}\leftarrow 0$
@@
 22:   $\forall i:$ Broadcast $\sigma^{i}_{k+1},\pi^{i}_{k+1}$
-23:        $\forall i:J^{i}_{t}\leftarrow\{j\in\mathcal{N}:(i,j)\in\mathcal{E}^{comm}_{t}$}
-24:        $\forall i:$ Select $\mathrm{adopted}^{i}\sim$ Pr$\left(\mathrm{adopted}^{i}=j\right)$ = $\frac{\exp{(\sigma^{j}_{k+1}}/\tau^{comm}_{k})}{\sum_{x\in J^{i}_{t}}\exp{(\sigma^{x}_{k+1}}/\tau^{comm}_{k})}$ $\forall j\in J^{i}_{t}$
+23:   $\forall i:J^{i}_{t}\leftarrow i\cup\{j\in\mathcal{N}:(i,j)\in\mathcal{E}^{comm}_{t}$}
+[…]
 25:   $\forall i:\sigma^{i}_{k+1}\leftarrow\sigma^{\mathrm{adopted}^{i}}_{k+1},\pi^{i}_{k+1}\leftarrow\pi^{\mathrm{adopted}^{i}}_{k+1}$
@@
 29:return policies $\pi^{i}_{K}$$_{i=1}^{N}$
-3 Learning and policy improvement
-3.1 Q-network and update
-Lines 1-14 of our novel Alg.
-1 contain the core Q-function/policy update method.
-Agent $i$ has a neural network parametrised by $\theta^{i}_{k}$ to approximate its Q-function: $\check{Q}_{\theta^{i}_{k}}(o,\cdot)$.
-The agent’s policy is given by $\pi_{\theta^{i}_{k}}(a|o)=$ softmax$\left(\frac{1}{\tau_{q}}\check{Q}_{\theta^{i}_{k}}(o,\cdot)\right)(a)$.
-We denote the policy $\pi^{i}_{k}(a|o)$ for simplicity when appropriate.
-Each agent maintains a buffer (of size $M$) of collected transitions of the form $\left(o^{i}_{t},a^{i}_{t},r^{i}_{t},o^{i}_{t+1}\right)$.
-At each iteration $k$, they empty their buffer (Line 3) before collecting $M$ new transitions (Lines 4-7); each decentralised agent $i$ then trains its Q-network $\check{Q}_{\theta^{i}_{k}}$ via $L$ training updates as follows (Lines 8-12).
-For training purposes, $i$ also maintains a target network $\check{Q}_{\theta^{i,^{\prime}}_{k,l}}$ with the same architecture but parameters $\theta^{i,^{\prime}}_{k,l}$ copied from $\theta^{i}_{k,l}$ less regularly than $\theta^{i}_{k,l}$ themselves are updated, i.e.
-only every $\nu$ learning iterations (Line 11).
-At each iteration $l$, the agent samples a random batch $B^{i}_{k,l}$ of $|B|$ transitions from its buffer (Line 9), and trains its neural network via stochastic gradient descent to minimise the empirical loss (Def.
-8, Line 10).
-For $cl<0$, $[\cdot]^{0}_{cl}$ is a clipping function used in Munchausen RL to prevent numerical issues if the policy is too close to deterministic, as the log-policy term is otherwise unbounded Vieillard et al.
-(2020); Wu et al.
-(2024):
-Definition 8 (Empirical loss for Q-network).
-This is given by
-^$\hat{\mathcal{L}}(\theta,\theta^{\prime})=\frac{1}{|B|}\sum_{transition\in B^{i}_{k,l}}\left|\check{Q}_{\theta^{i}_{k,l}}(o_{t},a_{t})-T\right|^{2},$
-where the target $T$ is
-$T=r_{t}\;+\;[\tau_{q}\ln\pi_{\theta^{i,^{\prime}}_{k,l}}(a_{t}|o_{t})]^{0}_{cl}\;+$
-$\gamma\sum_{a\in\mathcal{A}}\pi_{\theta^{i,^{\prime}}_{k,l}}(a|o_{t+1})\left(\check{Q}_{\theta^{i,^{\prime}}_{k,l}}(o_{t+1},a)-\tau_{q}\ln\pi_{\theta^{i,^{\prime}}_{k,l}}(a|o_{t+1})\right).$
-3.2 Communication and adoption of parameters
+4. Learning and policy improvement
+4.1. Q-network and update
+[…]
+Definition 4.1 (Empirical loss for Q-network).
+For the target $T$:
+$T=r_{t}$ $+[\tau_{q}\ln\pi_{\theta^{i,^{\prime}}_{k,l}}(a_{t}|o_{t})]^{0}_{cl}$
+$+\gamma\sum_{a\in\mathcal{A}}\pi_{\theta^{i,^{\prime}}_{k,l}}(a|o_{t+1})\left(\check{Q}_{\theta^{i,^{\prime}}_{k,l}}(o_{t+1},a)-\tau_{q}\ln\pi_{\theta^{i,^{\prime}}_{k,l}}(a|o_{t+1})\right),$
+$\text{the loss is:}\;\;\;\;\;\;\;\;\;\;\hat{\mathcal{L}}(\theta,\theta^{\prime})=\frac{1}{|B|}\sum_{transition\in B^{i}_{k,l}}\left|\check{Q}_{\theta^{i}_{k,l}}(o_{t},a_{t})-T\right|^{2}.$
+4.2. Communication and adoption of parameters
 We use the communication network $\mathcal{G}^{comm}_{t}$ to share two types of information at different points in Alg 1.
-One is used to improve local estimates of the mean-field distribution (see Sec.
-4).
-The other, described here, is used to privilege the spread of better performing policy updates through the population, allowing faster convergence in this networked case than in the independent and even centralised cases.
-We adapt the work in Benjamin and Abate (2023) for the function-approximation case, where in our work agents broadcast the parameters of the Q-network that defines their policy, rather than the Q-function table.
-At each iteration $k$, after independently updating their Q-network and policy (Lines 3-14), agents approximate the infinite discounted return (Def.
-3) of their new policies by collecting rewards for $E$ steps, and assign the finite-step discounted sum to $\sigma^{i}_{k+1}$ (Lines 15-20).
+One is used to improve local estimates of the mean field (Sec.
+5).
+The other, described here, is used to privilege the spread of better performing policy updates through the population, allowing faster convergence in this networked case than in the independent and even central-agent cases.
+We adapt benjamin2024networked for the function-approximation case, where in our work agents broadcast the parameters of the Q-network that defines their policy, rather than the Q-function table.
+At each iteration $k$, after independently updating their Q-network and policy (Lines 3-14), agents estimate the infinite discounted return (Def.
+3.3) of their new policies by collecting rewards for $E$ steps, and assign the finite-step discounted sum to $\sigma^{i}_{k+1}$ (Lines 15-20).
 They then broadcast their Q-network parameters along with $\sigma^{i}_{k+1}$ (Line 22).
@@
 This allows decentralised agents to adopt policy parameters estimated to perform better than their own, accelerating learning as shown in Sec.
-5.
-4 Mean-field estimation and communication
-We first describe the most general version of our algorithm for decentralised estimation of the empirical categorical mean-field distribution, assuming the more general setting where $\mathcal{G}^{obs}_{t}$ applies (see discussion in Sec.
-2.3).
+6.
+5. Mean-field estimation and communication
+We now give our algorithms for decentralised estimation of the empirical categorical mean-field distribution.
+We first describe the general version, assuming the more general setting where $\mathcal{G}^{obs}_{t}$ applies (see discussion in Sec.
+3.3).
 We subsequently detail how the algorithm can be made more efficient in environments where the more abstract visibility graph $\mathcal{G}^{vis}_{t}$ applies, as in our experimental settings.
@@
 to produce $o^{i}_{t}=(s^{i}_{t},\tilde{\hat{\mu}}^{i}_{t})$ for the steps $a^{i}_{t}\sim\pi^{i}_{k}(\cdot|o^{i}_{t})$ in Lines 5, 17 and 26.
-Both versions of the algorithm are subject to implicit assumptions, which we highlight and discuss methods for addressing in Sup.
-Mat.
-E.
-4.0.1 Algorithm for the general setting
-Algorithm 2 Mean-field estimation and communication in general settings
+Note that if $\mathcal{G}^{obs}_{t}$$\mathcal{G}^{vis}_{t}$ are fully connected, all agents’ estimated mean-field observations will be equivalent to the true categorical distribution.
+Both versions of the algorithm are subject to implicit assumptions, which we discuss methods for addressing as future work in Appx.
+C.
+5.1. Algorithm for the general setting
+Algorithm 2 Mean-field estimation in general settings
 0: Time-dependent observation graph $\mathcal{G}^{obs}_{t}$, time-dependent communication graph $\mathcal{G}^{comm}_{t}$, states $\{s^{i}_{t}\}_{i=1}^{N}$, number of communication rounds $C_{e}$
-1:  $\forall i,s:$ Initialise count vector $\hat{\upsilon}^{i}_{t}[s]$ with $\emptyset$
-2:  $\forall i:$ $\hat{\upsilon}^{i}_{t}[s^{j}_{t}]$leftarrow $\{ID^{j}\}_{j\in\mathcal{N}:(i,j)\in\mathcal{E}^{obs}_{t}}$
+[…]
+[…]
 3:for $c_{e}$ in $1,\dots,C_{e}$ do
@@
 7:end for
-8:  $\forall i:$ $counted\_agents^{i}_{t}\leftarrow\sum_{s\in\mathcal{S}:\hat{\upsilon}^{i}_{t}[s]\neq\emptyset}|\hat{\upsilon}^{i}_{t}[s]|$
-9:  $\forall i:$ $uncounted\_agents^{i}_{t}\leftarrow N-counted\_agents^{i}_{t}$
-10:  $\forall i,s$ : $\tilde{\hat{\mu}}^{i}_{t}[s]\leftarrow\frac{uncounted\_agents^{i}_{t}}{N\times|\mathcal{S}|}$
-11:  $\forall i,s$ where $\hat{\upsilon}^{i}_{t}[s]$ is not $\emptyset$ : $\tilde{\hat{\mu}}^{i}_{t}[s]\leftarrow\tilde{\hat{\mu}}^{i}_{t}[s]+\frac{|\hat{\upsilon}^{i}_{t}[s]|}{N}$
+[…]
+[…]
+[…]
+[…]
 12:return mean-field estimates $\{\tilde{\hat{\mu}}^{i}_{t}\}_{i=1}^{N}$
@@
 Each state position in the vector can hold a list of IDs.
-Before any actions are taken at each time step $t$, each agent’s count vector ${\hat{\upsilon}}^{i}_{t}$ is initialised as full of $\emptyset$ (‘no count’) markers for each state.
-Then, for each agent $j$ with which agent $i$ is connected via the observation graph, $i$ places $j$’s unique ID in its count vector in the correct state position.
-Next, for $C_{e}\geq 0$ communication rounds, agents exchange their local counts with neighbours on the communication network, and merge these counts with their own count vector, filtering out the unique IDs of those that have already been counted.
+Before any actions are taken at each time step $t$, each agent’s count vector ${\hat{\upsilon}}^{i}_{t}$ is initialised as full of $\emptyset$ (‘no count’) markers for each state (Line 1).
+Then, for each agent $j$ with which agent $i$ is connected via the observation graph, $i$ places $j$’s unique ID in its count vector in the correct state position (Line 2).
+Next, for $C_{e}\geq 0$ communication rounds, agents exchange their local counts with neighbours on the communication network (Line 4), and merge these counts with their own count vector, filtering out the unique IDs of those that have already been counted (Line 6).
 If $C_{e}=0$ then the local count will remain purely independent.
@@
 The local mean-field estimate $\tilde{\hat{\mu}}^{i}_{t}$ is then obtained from ${\hat{\upsilon}}^{i}_{t}$ as follows.
-All states that have a count $x_{s}$ have this count converted into the categorical probability $x_{s}/N$ (we assume that agents know the total number of agents in the finite population, even if they cannot observe them all at each $t$).
-The total number of agents counted in ${\hat{\upsilon}}^{i}_{t}$ is given by $counted\_agents$ = $\sum_{s\in\mathcal{S}}x_{s}$, and the agents that have not been observed are $uncounted\_agents$ = $N$ $counted\_agents$.
-In this general setting, the unobserved agents are assumed to be uniformly distributed across all the states, so $uncounted\_agents/(N\times|\mathcal{S}|)$ is added to all the values in $\tilde{\hat{\mu}}^{i}_{t}$, replacing the $\emptyset$ marker for states for which no agents have been observed.
-4.0.2 Algorithm for visibility-based environments
-We give now the differences in our estimation algorithm (Alg.
-3, Sup.
-Mat.
-A) for the subclass of environments where $\mathcal{G}^{vis}_{t}$ applies in place of $\mathcal{G}^{obs}_{t}$, i.e.
+All states that have a count $x_{s}$ have this count converted into the categorical probability $x_{s}/N$ (we assume that agents know the total number of agents in the finite population, even if they cannot observe them all at each $t$) (Line 11).
+The total number of agents counted in ${\hat{\upsilon}}^{i}_{t}$ is given by counted_agentscounted\_agents = $\sum_{s\in\mathcal{S}}x_{s}$, and the agents that have not been observed are uncounted_agentsuncounted\_agents = $N$ - counted_agentscounted\_agents.
+In this general setting, the unobserved agents are assumed to be uniformly distributed across all the states, so $uncounted\_agents/(N\times|\mathcal{S}|)$ is added to all the values in $\tilde{\hat{\mu}}^{i}_{t}$, replacing the $\emptyset$ marker for states for which no agents have been observed (Line 10).
+Algorithm 3 Mean-field estimation for environments with $\mathcal{G}^{vis}_{t}$
+0: Time-dependent visibility graph $\mathcal{G}^{vis}_{t}$, time-dependent communication graph $\mathcal{G}^{comm}_{t}$, states $\{s^{i}_{t}\}_{i=1}^{N}$, number of communication rounds $C_{e}$
+[…]
+[…]
+3:for $c_{e}$ in $1,\dots,C_{e}$ do
+4:  $\forall i:$ Broadcast $\hat{\upsilon}^{i}_{t,c_{e}}$
+5:  $\forall i:J^{i}_{t}=i\cup\{j\in\mathcal{N}:(i,j)\in\mathcal{E}^{comm}_{t}\}$
+6:  $\forall i,s:$ Initialise new count vector $\hat{\upsilon}^{i}_{t,(c_{e}+1)}[s]$ with $\emptyset$
+7:  $\forall i,s$ and $\forall j\in J^{i}_{t}:\hat{\upsilon}^{i}_{t,(c_{e}+1)}[s]\leftarrow\hat{\upsilon}^{j}_{t,c_{e}}[s]$ if $\hat{\upsilon}^{j}_{t,c_{e}}[s]\neq\emptyset$
+8:end for
+[…]
+[…]
+[…]
+[…]
+[…]
+14:return mean-field estimates $\{\tilde{\hat{\mu}}^{i}_{t}\}_{i=1}^{N}$
+5.2. Algorithm for visibility-based environments
+We explain now the differences in our estimation algorithm (Alg.
+3) for the subclass of environments where $\mathcal{G}^{vis}_{t}$ applies in place of $\mathcal{G}^{obs}_{t}$, i.e.
 the mutual observability of agents depends in turn on the mutual visibility of states.
@@
 Unique IDs are thus not required as there is no risk of counting the same agent twice when receiving communicated counts: either all agents in $s^{\prime}$ have been counted, or no count has yet been obtained for $s^{\prime}$.
-This simplifies the algorithm and helps preserve agent privacy.
+This simplifies the algorithm and helps preserve agent anonymity and privacy.
 Secondly, uncounted agents cannot be in states for which a count has already been obtained, since the count is complete and correct, even if the count is $x_{s^{\prime}}=0$.
-Therefore after the $C_{e}$ communication rounds, the $uncounted\_agents$ proportion needs to be uniformly distributed only across the positions in the vector that still have the $\emptyset$ marker, and not across all states as in the general setting.
-This makes the estimation more accurate in this special setting.
-5 Theoretical results
-To demonstrate the benefits of the networked architecture by comparison, we also consider (theoretically here and experimentally in Sec.
-6) the results of modified versions of our algorithm for centralised and independent learners.
-In the centralised setting, the Q-network updates of arbitrary agent $i=1$ are automatically pushed to all other agents, and the true global mean-field distribution is always used in place of the local estimate i.e.
+Therefore after the $C_{e}$ communication rounds, the uncounted_agentsuncounted\_agents proportion needs to be uniformly distributed only across the positions in the vector that still have the $\emptyset$ marker (Line 13), and not across all states as in the general setting.
+This makes the estimation more accurate in this setting.
+[^3] In our Algs.
+2 and 3, agents share their local counts with neighbours on the communication network $\mathcal{G}^{comm}_{t}$, and only after the $C_{e}$ communication rounds do they complete their estimated distribution by distributing the uncounted agents along their vectors.
+An alternative would be for each agent to immediately form a local estimate from their local count obtained via $\mathcal{G}^{obs}_{t}$ or $\mathcal{G}^{vis}_{t}$, which is only then communicated and updated via the communication network.
+However, we take the former approach to avoid poor local estimations spreading through the network and leading to widespread inaccuracies.
+Information that is certain (the count) is spread as widely as possible, before being locally converted into an estimate of the total mean field.
+The same would be the case in our extension proposed in Appx.
+C for averaging noisy counts, i.e.
+only the counts would be averaged, with the estimates completed by distributing the remaining agents after the $C_{e}$ rounds.
+6. Theoretical results
+To demonstrate the benefits of the networked architecture by comparison, we also consider the results of baseline central-agent and independent architectures given by alternative versions of our algorithm.
+As in previous MFG works (policy_mirror_independent; benjamin2024networked), in the central-agent setting, the Q-network updates of arbitrary agent $i=1$ are automatically pushed to all other agents, and the true global mean-field distribution is always used in place of the local estimate i.e.
 $\tilde{\hat{\mu}}^{i}_{t}$ = ${\hat{\mu}}_{t}$.
-In the independent case, there are no links in $\mathcal{G}^{comm}_{t}$ or $\mathcal{G}^{vis}_{t}$, i.e.
-$\mathcal{E}^{comm}_{t}=\mathcal{E}^{vis}_{t}=\emptyset$.
-Networked agents often learn faster than centralised ones in our experiments; we justify theoretically this possibly counterintuitive result here.
-We first make two strong assumptions that give conditions under which networked agents do outperform centralised ones.
-The fact that these strong assumptions do not always hold in reality explains why networked agents may not always outperform centralised ones.
+In the independent case, there are no links in $\mathcal{G}^{comm}_{t}$ or $\mathcal{G}^{obs}_{t}$$\mathcal{G}^{vis}_{t}$, i.e.
+$\mathcal{E}^{comm}_{t}=\mathcal{E}^{obs}_{t}=\mathcal{E}^{vis}_{t}=\emptyset$.
+Networked populations often learn faster than central-agent ones in our experiments.
+To indicate how this is possible while allowing simplicity of the theory, we give a proof for a special case with relatively strong assumptions that give conditions under which networked populations do outperform central-agent ones.
+Nevertheless the intuition provided by Thm.
+6.3’s proof suggests why networked agents can learn faster even without enforcing the assumptions, and we discuss loosening them subsequently.
 Recall that at each iteration $k$ of Alg.
@@
 There is randomness in these independent policy updates, stemming from the random sampling of each agent’s independently collected buffer.
-In Lines 15-20, agents approximate the infinite discounted returns {$V(\pi^{i}_{k+1},I(\pi^{i}_{k+1}))\}_{i=1}^{N}$ (Def.
-3) of their updated policies by computing $\sigma^{i}_{k+1}$}${}_{i=1}^{N}$: the $E$-step discounted return with respect to the empirical mean field generated when agents follow policies $\pi^{i}_{k+1}$}${}_{i=1}^{N}$ (i.e.
+In Lines 15-20, agents estimate the infinite discounted returns $V(\pi^{i}_{k+1},I(\pi^{i}_{k+1}))\}_{i=1}^{N}$ (Def.
+3.3) of their updated policies by computing $\sigma^{i}_{k+1}$$_{i=1}^{N}$: the $E$-step discounted return with respect to the empirical mean field generated when agents follow policies $\pi^{i}_{k+1}$$_{i=1}^{N}$ (i.e.
 they do not at this stage all follow a single identical policy).
 We can characterise the approximation as $\sigma^{i}_{k+1}$$_{i=1}^{N}$ = $\widehat{V}(\pi^{i}_{k+1},I(\pi^{i}_{k+1}))\}_{i=1}^{N}$.
-Assumption 1.
-Assume that $\sigma^{i}_{k+1}$}${}_{i=1}^{N}$ are sufficiently good approximations so as to respect the ordering of the true values {$V(\pi^{i}_{k+1},I(\pi^{i}_{k+1}))\}_{i=1}^{N}$, i.e.
-$\forall i,j\in\{1,\dots,N\}$: $\sigma^{i}_{k+1}>\sigma^{j}_{k+1}\iff V(\pi^{i}_{k+1},I(\pi^{i}_{k+1}))>V(\pi^{j}_{k+1},I(\pi^{j}_{k+1})).$
-Assumption 2.
-Assume that after the $C_{p}$ rounds in Lines 21-27 in which agents exchange and adopt policies from neighbours, the population is left with a single policy such that $\forall i,j\in\{1,\dots,N\}$ $\pi^{i}_{k+1}=\pi^{j}_{k+1}$.
-[^2] Most simply we can think of Ass.
-2 holding if 1) $\tau^{comm}_{k}\rightarrow 0\;\forall k$ such that the softmax essentially becomes a max function, and 2) the communication network $\mathcal{G}^{comm}_{t}$ is static and connected during the $C_{p}$ communication rounds, where $C_{p}$ is larger than the network diameter $d_{\mathcal{G}^{comm}_{t}}$ Under these conditions, previous results on max-consensus algorithms show that all agents in the network will converge on the highest $\sigma^{max}_{k+1}$ value (and hence the unique associated $\pi^{max}_{k+1}$ within a number of rounds equal to the diameter $d_{\mathcal{G}^{comm}_{t}}$ Nejad et al.
-(2009); Benjamin and Abate (2023).
-However, policy consensus as in Ass.
-2 might be achieved even outside of these conditions, including if the network is dynamic and not connected at every step, given appropriate values for $C_{p}$ and $\tau^{comm}_{k+1}\in\mathbb{R}_{>0}$.
-Call the network consensus policy $\pi^{\mathrm{net}}_{k+1}$, and its associated finitely approximated return $\sigma^{\mathrm{net}}_{k+1}$.
-Recall that the centralised case is where the Q-network update of arbitrary agent $i=1$ is automatically pushed to all the others instead of the policy evaluation and exchange in Lines 15-27; this is equivalent to a networked case where policy consensus is reached on a random one of the policies $\pi^{i}_{k+1}$}${}_{i=1}^{N}$.
-Call this policy arbitrarily given to the whole population $\pi^{\mathrm{cent}}_{k+1}$, and its associated finitely approximated return $\sigma^{\mathrm{cent}}_{k+1}$.
-Theorem 1.
-Given Ass.
-1 and 2, $\mathbb{E}[V(\pi^{\mathrm{net}}_{k+1},I(\pi^{\mathrm{net}}_{k+1}))]>\mathbb{E}[V(\pi^{\mathrm{cent}}_{k+1},I(\pi^{\mathrm{cent}}_{k+1}))]$.
-Thus in expectation networked agents will increase their returns faster than centralised ones.
+We now assume the following:
+Assumption 6.1.
+After the $C_{p}$ rounds in Lines 21-27 in which agents exchange and adopt policies from neighbours, the population is left with a single policy such that $\forall i,j\in\{1,\dots,N\}$ $\pi^{i}_{k+1}=\pi^{j}_{k+1}$.
+[^4] Most simply Assumption 6.1 holds if 1) $\tau^{comm}_{k}$ is a positive constant sufficiently close to zero that the softmax essentially becomes a max function, and 2) the communication network $\mathcal{G}^{comm}_{t}$ is static and connected during the $C_{p}$ communication rounds, where $C_{p}$ is larger than the network diameter $d_{\mathcal{G}^{comm}_{t}}$.
+Under these conditions, previous results on max-consensus algorithms show that all agents in the network will converge on the highest $\sigma^{max}_{k+1}$ value (and hence the unique associated $\pi^{max}_{k+1}$) within a number of rounds equal to the diameter $d_{\mathcal{G}^{comm}_{t}}$ (5348437; benjamin2024networked).
+However, policy consensus can be achieved even outside of these conditions, including if the network is dynamic and not connected at every step, given appropriate values for $C_{p}$ and $\tau^{comm}_{k+1}\in\mathbb{R}_{>0}$.
+Assumption 6.2.
+Assume that $\sigma^{i}_{k+1}$$_{i=1}^{N}$ are sufficiently good approximations so as to respect the ordering of the true values
+$V(\pi^{i}_{k+1},I(\pi^{i}_{k+1}))\}_{i=1}^{N}$, i.e. $\forall i,j\in\{1,\dots,N\}$:
+$\sigma^{i}_{k+1}>\sigma^{j}_{k+1}\iff V(\pi^{i}_{k+1},I(\pi^{i}_{k+1}))>V(\pi^{j}_{k+1},I(\pi^{j}_{k+1})).$
+Call the network consensus policy $\pi^{\mathrm{net}}_{k+1}$, and its associated finitely estimated return $\sigma^{\mathrm{net}}_{k+1}$.
+Recall that the central-agent case is where the Q-network update of arbitrary agent $i=1$ is automatically pushed to all the others instead of the policy exchange in Lines 15-27; this is equivalent to a networked case where policy consensus is reached on a random one of the policies $\pi^{i}_{k+1}$$_{i=1}^{N}$.
+Call this policy arbitrarily given to the whole population $\pi^{\mathrm{cent}}_{k+1}$, and its associated finitely estimated return $\sigma^{\mathrm{cent}}_{k+1}$.
+Now we can say:
+Theorem 6.3.
+Given Assumptions 6.1 and 6.2,
+$\mathbb{E}[V(\pi^{\mathrm{net}}_{k+1},I(\pi^{\mathrm{net}}_{k+1}))]>\mathbb{E}[V(\pi^{\mathrm{cent}}_{k+1},I(\pi^{\mathrm{cent}}_{k+1}))].$
+Thus in expectation networked populations will increase their returns faster than central-agent ones.
 Proof.
 Recall that before the communication rounds in Line 21 (Alg.
-1), the randomly updated policies $\pi^{i}_{k+1}$}${}_{i=1}^{N}$ have associated approximated returns $\sigma^{i}_{k+1}$}${}_{i=1}^{N}$.
-Denote the mean and maximum of this set $\sigma^{\mathrm{mean}}_{k+1}$ and $\sigma^{\mathrm{max}}_{k+1}$ respectively.
+1), the randomly updated policies $\pi^{i}_{k+1}$$_{i=1}^{N}$ have associated estimated returns $\sigma^{i}_{k+1}$$_{i=1}^{N}$.
+Call the mean and maximum of this set $\sigma^{\mathrm{mean}}_{k+1}$ and $\sigma^{\mathrm{max}}_{k+1}$ respectively.
 Since $\pi^{\mathrm{cent}}_{k+1}$ is chosen arbitrarily from $\pi^{i}_{k+1}$$_{i=1}^{N}$, it will obey $\mathbb{E}[\sigma^{\mathrm{cent}}_{k+1}]=\sigma^{\mathrm{mean}}_{k+1}$ $\forall k$, though there will be high variance.
@@
 1) for the networked case means by definition that policies with higher $\sigma^{i}_{k+1}$ are more likely to be adopted at each communication round.
-Thus the $\pi^{\mathrm{net}}_{k+1}$ that gets adopted by the whole networked population will obey $\mathbb{E}[\sigma^{\mathrm{net}}_{k+1}]>\sigma^{\mathrm{mean}}_{k+1}$ (if $\tau^{comm}_{k+1}\rightarrow 0$, it will obey $\mathbb{E}[\sigma^{\mathrm{net}}_{k+1}]=\sigma^{\mathrm{max}}_{k+1}$ $\forall k$).
-As such, $\mathbb{E}[\sigma^{\mathrm{net}}_{k+1}]>\mathbb{E}[\sigma^{\mathrm{cent}}_{k+1}]$, which by Ass.
-1 implies the result.
+Thus the $\pi^{\mathrm{net}}_{k+1}$ that is adopted by the whole networked population will obey $\mathbb{E}[\sigma^{\mathrm{net}}_{k+1}]>\sigma^{\mathrm{mean}}_{k+1}$ (if $\tau^{comm}_{k+1}$ is a positive value near zero, it will obey $\mathbb{E}[\sigma^{\mathrm{net}}_{k+1}]=\sigma^{\mathrm{max}}_{k+1}$ $\forall k$).
+So $\mathbb{E}[\sigma^{\mathrm{net}}_{k+1}]>\mathbb{E}[\sigma^{\mathrm{cent}}_{k+1}]$, which by Assumption 6.2 implies the result.
 ∎
-The adoption scheme in Line 24 biases the spread of policies towards those estimated to be better, which, given sufficiently good approximations (Ass.
-1), results in higher discounted returns in practice.
-By choosing updates in a more principled way, networked agents learn faster than the centralised case that adopts updates regardless of quality.
-This intuition applies even if we loosen Ass.
-2 that the networked population converges on a single consensus policy within the $C_{p}$ communication rounds.
-Similar logic can also be applied to understand why networked agents outperform entirely independent ones, combined with the fact that divergence between policies in the independent case worsens sample complexity over the networked and centralised cases by biasing approximations of the Q-function Yardim et al.
-(2023); Benjamin and Abate (2023).
-Significantly, the communication scheme not only allows us to avoid the undesirable assumption of a central learner, but even to outperform it.
-Moreover, the benefit of networked communication over centralised learning is even greater in our function approximation setting than in the tabular case of Benjamin and Abate (2023), perhaps due to greater variance in the quality of Q-function estimates in our case.
-6 Experiments
+The adoption scheme in Line 24 biases the spread of policies towards those estimated to be better, which, given sufficiently good approximations (Assumption 6.2), results in higher discounted returns in practice.
+By choosing updates in a more principled way, networked agents learn faster than the central-agent case that pushes updates regardless of quality.
+[^5] Assumption 6.1 may not hold if $C_{p}$ is not large enough or if parts of the population remain isolated.
+Thus in our experiments, where we use $C_{p}=1$ to show the benefit of even one communication round, networked populations with smaller broadcast radii outperform central-agent populations by a smaller margin.
+Nevertheless the intuition provided by Thm.
+6.3’s proof indicates how the former are still able to perform better even if Assumption 6.1 is loosened.
+[^6] Even if Assumption 6.2 does not strictly hold, the softmax parameter $\tau^{comm}_{k}$ allows a smooth degradation as the ordering of the approximations worsens with respect to the ordering of the true values.
+I.e.
+if instead of the exact correct policy ordering we have that better policies are simply more likely to be given higher estimated evaluations, then the softmax means that these policies remain more likely to spread, and the best policy may still be adopted even if it is not evaluated as the best.
+Similar logic can be applied to understand why networked agents outperform independent ones, coupled with the fact that greater policy diversity in the independent case worsens sample complexity over the networked and central-agent cases by biasing approximations of the Q-function (policy_mirror_independent; benjamin2024networked).
+Significantly, our communication scheme not only avoids the undesirable assumption of a central node, but even outperforms it.
+Moreover, the benefit of the scheme over central-agent learning is greater with our function approximation than in the tabular case (cf.
+(benjamin2024networked)), perhaps due to greater variance in the quality of Q-function estimates in our case.
+This shows that networked communication facilitates greater scalability than the central-agent paradigm.
+7. Experiments
 We provide two sets of experiments.
-The first set showcases that our function-approximation algorithm (Alg.
-1) can scale to large state spaces for population-independent policies, and that in such settings networked, communicating agents can outperform purely-independent and even centralised agents, and do so by an even greater margin than in the tabular settings from Benjamin and Abate (2023).
-The second set (given in Sup.
-Mat.
-C.4) demonstrates that Alg.
+The first shows that our function-approximation algorithm (Alg.
+1) can scale to large state spaces for population-independent policies, and that in such settings networked, communicating populations can outperform purely-independent and even central-agent populations.
+The second set demonstrates that Alg.
 1 can handle population-dependent policies, as well as the ability of Alg.
-3 to practically estimate the mean-field distribution locally.
-For the types of game used in our demonstrations we follow the gold standard in prior MFG works, i.e.
-grid-world environments where agents can move in the four cardinal directions or remain in place Laurière (2021); Laurière et al.
-(2022b); Zaman et al.
-(2023); Algumaei et al.
-(2023); Cui et al.
-(2023); Benjamin and Abate (2023); Wu et al.
-(2024).
-We present results from five tasks defined by the agents’ reward/transition functions, all but one of which are coordination tasks.
-The first two tasks are those used with population-independent policies in Benjamin and Abate (2023), but while they show results for an 8x8 and a ‘larger’ 16x16 grid, our results are for 100x100 and 50x50 grids:
+3 to estimate the mean field locally.
+7.1. Experimental set-up
+For the types of game in our experiments we follow the gold standard in prior MFG works, i.e.
+grid-world environments where agents can move in the four cardinal directions or remain in place (lauriere2021numerical; scalable_deep; zaman2023oraclefree; algumaei2023regularization; cui2023multiagent; benjamin2024networked; wu2024populationaware).
+We present results from four tasks defined by the agents’ reward/transition functions, all of which are coordination tasks - see Appx.
+A.1 for full technical descriptions, and also a fifth, non-coordination task.
+In all cases, rewards are normalised in [0,1] after they are computed.
+The first two tasks are those used with population-independent policies in benjamin2024networked, but while they show results for an 8x8 and a ‘larger’ 16x16 grid, our results are for $100\times 100$ and $50\times 50$ grids:
   • Cluster. Agents are rewarded for gathering but given no indication where to do so, agreeing it over time.
@@
 Agents must coordinate on which single target they will all meet at to maximise their individual rewards.
-See Sup.
-Mat.
-C.1 for a full technical description of these tasks, as well as descriptions of additional more complex tasks.
-In these spatial environments, both the communication network $\mathcal{G}^{comm}_{t}$ and the visibility graph $\mathcal{G}^{vis}_{t}$ are determined by the physical distance from agent $i$; we show plots for various radii, expressed as fractions of the maximum possible distance (the grid’s diagonal length).
-We discuss hyperparameters in Sup.
-Mat.
-C.3.
-Figure 1: ‘Target agreement’, pop.-independent, 100x100 grid.
-Figure 2: ‘Cluster’, pop.-independent, 100x100 grid.
+We also show our algorithms handling more complex tasks:
+  • Evade shark in shoal.
+At each $t$, a ‘shark’ takes a step towards the grid point containing the most agents according to the empirical mean field.
+The shark’s position is part of agents’ local states in addition to their own position.
+Agents are rewarded more for being further from the shark and for clustering with other agents.
+As well as having a non-stationary distribution, we add ‘common noise’, with the shark taking a random step with probability 0.01.
+Such noise that affects the local states of all agents in the same way, making the evolution of the distribution stochastic, makes population-independent policies sub-optimal (survey_learningMFGs).
+  • Push object to edge.
+Agents are rewarded for how close an ‘object’ is to the grid’s edge.
+The object’s position forms part of agents’ local states in addition to their own position.
+The object moves in a direction with a probability proportional to the number of agents on its opposite side, i.e.
+agents must coordinate on which side of the object from which to ‘push’ it, to ensure it moves toward the edge of the grid.
 We evaluate our experiments via two metrics.
-Exploitability is the most common metric in works on MFGs, and is a measure of proximity to the MFNE.
-It quantifies how much a best-responding agent can benefit by deviating from the set of policies that generate the current mean-field distribution, with a decreasing exploitability meaning the population is closer to the MFNE.
-However, there are several issues with this metric in our setting, particularly for our coordination tasks where competitive agents benefit from aligning behaviours, such that it may give limited or noisy information (discussed further in Sup.
-Mat.
-C.2.1).
-We thus also give a second metric, as in Benjamin and Abate (2023): the population’s average discounted return.
-This allows us to compare how quickly agents are learning to increase their returns, even when ‘exploitability’ gives us limited ability to distinguish between the desirability of the MFNEs to which populations converge.
-6.1 Results and discussion
+Exploitability is the most common metric in MFG works, and is a measure of proximity to the MFNE.
+It quantifies how much an agent can benefit by deviating from the set of policies that generate the current mean field, with a lower exploitability meaning the population is closer to the MFNE.
+However, there are several issues with this metric in our setting, particularly for our coordination tasks where competitive agents benefit from aligning behaviours, such that it may give limited or noisy information (discussed further in Appx.
+A.2.1).
+We thus also give a second metric, as in benjamin2024networked: the population’s average discounted return.
+This allows us to compare how quickly agents are learning to increase their returns, even when exploitability gives us limited ability to distinguish between the desirability of the MFNEs to which populations converge.
+7.2. Results and discussion
+In our spatial environments, the physical distance from $i$ determines the communication graph $\mathcal{G}^{comm}_{t}$ and the visibility graph $\mathcal{G}^{vis}_{t}$.
+Our plots show various radii, given as fractions of the maximum possible distance (the grid’s diagonal length).
+We set $C_{p}=C_{e}=1$ to show the benefit to learning speed brought by even a single communication round.
+Note that the networked population with the largest radius is always fully connected, and therefore these agents are always able to accurately estimate ${\hat{\mu}}_{t}$ even for $C_{e}=0$.
+That is, their observations are equivalent to those that the central-agent population would receive, albeit that policies are updated and spread differently.
+Other hyperparameters are detailed in Appx.
+A.3.
+Figure 1.
+‘Target agreement’, pop.-independent, $100\times 100$ grid.
+Reproduced larger in Fig.
+3.
+The networked populations of all broadcast radii significantly outperform the central-agent and independent populations in terms of average return, where the latter two cases hardly appear to learn at all.
+[…]
+‘Target agreement’, population-independent, $100\times 100$ grid.
+The networked populations of all broadcast radii significantly outperform the central-agent and independent populations in terms of average return, where the latter two cases hardly appear to learn at all.
+Figure 2.
+‘Cluster’, pop.-independent, $100\times 100$ grid.
+Larger version in Fig.
+4.
+Networked populations of all broadcast radii outperform the central-agent and independent populations wrt.
+average return; independent agents hardly learn at all.
+[…]
+‘Cluster’, population-independent, $100\times 100$ grid.
+The networked populations of all broadcast radii outperform the central-agent and independent populations in terms of average return; independent agents hardly appear to learn at all.
+7.2.1. Population-independent policies in large state-spaces
 Figs.
-1 and 2 illustrate that introducing function approximation to algorithms in this setting allows them to converge within a practical number of iterations ($k\ll 100$), even for large state spaces (100x100 grids).
-By contrast, the tabular algorithms in Benjamin and Abate (2023) appear only just to converge by $k=200$ for the same tasks for the larger of their two grids, which is only 16x16.
+1 and 2 (for $100\times 100$ and grids; reproduced larger in Figs.
+3 and 4, Appx.
+A), and Figs.
+5 and 7 (for $50\times 50$ grids; Appx.
+A) illustrate that introducing function approximation to algorithms in this setting allows them to converge within a practical number of iterations ($k\ll 100$), even for large state spaces.
+By contrast, the tabular algorithms in benjamin2024networked appear only just to converge by $k=200$ for the same tasks for the larger of their two grids, which is only $16\times 16$.
 In Figs.
-1 and 2, networked agents all significantly outperform both centralised and independent agents in term of average return, despite the centralised agents appearing to have similar exploitability, and the independent agents having similar or slightly lower exploitability.
-This is because independent agents (and also the centralised ones in Fig.
-1) hardly improve their policies at all, such that there is little a deviating agent can do to increase its return in these coordination games, meaning exploitability appears low, despite this being an undesirable equilibrium (see Sup.
-Mat.
-C.2 for further discussion on the limited information provided by the exploitation metric).
-The fact that the networked agents nevertheless significantly outperform the other architectures in terms of average return indicates that communication helps agents to find substantially ‘preferable’ equilibria.
-Moreover, the margin by which the networked agents can outperform the centralised agents is much greater than in Benjamin and Abate (2023), showing that the benefits of the communication scheme are even greater in non-tabular settings.
-See Sup.
-Mat.
-C.4 for further experiments with large state spaces, as well as others demonstrating the ability of our algorithm to handle two more complex tasks, using population-dependent policies and estimated mean-field observations.
-7 Conclusion
-We novelly contributed function approximation to the online-learning setting for empirical MFGs, and also contributed two novel algorithms for locally estimating the empirical mean field for population-dependent policies.
-We have justified theoretically why our networked communication algorithm is able to learn faster than both centralised and independent agents in this function approximation setting, and demonstrated empirically the ability of our algorithms to handle large state spaces and estimate the mean field.
-Limitations and ways to extend our algorithms are in Sup.
-Mat.
-E.
-Supplementary material to ‘Networked Communication for Mean-Field Games with Function Approximation and Empirical Mean-Field Estimation’
-Appendix A Mean-field estimation algorithm for visibility-based environments
-Algorithm 3 Mean-field estimation and communication for environments with $\mathcal{G}^{vis}_{t}$
-0:  Time-dependent visibility graph $\mathcal{G}^{vis}_{t}$, time-dependent communication graph $\mathcal{G}^{comm}_{t}$, states $\{s^{i}_{t}\}_{i=1}^{N}$, number of communication rounds $C_{e}$
-1:  $\forall i,s:$ Initialise count vector $\hat{\upsilon}^{i}_{t}[s]$ with $\emptyset$
-2:  $\forall i$ and $\forall s^{\prime}\in\mathcal{S^{\prime}}:(s^{i}_{t},s^{\prime})\in\mathcal{E}^{vis}_{t}$ : $\hat{\upsilon}^{i}_{t}[s^{\prime}]$leftarrow $\sum_{j\in 1,\dots,N:s^{j}_{t}=s^{\prime}}1$
-3:  for $c_{e}$ in $1,\dots,C_{e}$ do
-4:     $\forall i:$ Broadcast $\hat{\upsilon}^{i}_{t,c_{e}}$
-5:     $\forall i:J^{i}_{t}=i\cup\{j\in\mathcal{N}:(i,j)\in\mathcal{E}^{comm}_{t}\}$
-6:     $\forall i,s$ and $\forall j\in J^{i}_{t}:\hat{\upsilon}^{i}_{t,(c_{e}+1)}[s]\leftarrow\hat{\upsilon}^{j}_{t,c_{e}}[s]$ if $\hat{\upsilon}^{j}_{t,c_{e}}[s]\neq\emptyset$
-7:  end for
-8:  $\forall i:$ $counted\_agents^{i}_{t}\leftarrow\sum_{s\in\mathcal{S}:\hat{\upsilon}^{i}_{t}[s]\neq\emptyset}\hat{\upsilon}^{i}_{t}[s]$
-9:  $\forall i:$ $uncounted\_agents^{i}_{t}\leftarrow N-counted\_agents^{i}_{t}$
-10:  $\forall i:$ $unseen\_states^{i}_{t}\leftarrow\sum_{s\in\mathcal{S}:\hat{\upsilon}^{i}_{t}[s]=\emptyset}1$
-11:  $\forall i,s$ where $\hat{\upsilon}^{i}_{t}[s]$ is not $\emptyset$ : $\tilde{\hat{\mu}}^{i}_{t}[s]\leftarrow\frac{\hat{\upsilon}^{i}_{t}[s]}{N}$
-12:  $\forall i,s$ where $\hat{\upsilon}^{i}_{t}[s]$ is $\emptyset$ : $\tilde{\hat{\mu}}^{i}_{t}[s]\leftarrow\frac{uncounted\_agents^{i}_{t}}{N\times unobserved\_states^{i}_{t}}$
-13:  return  mean-field estimates $\{\tilde{\hat{\mu}}^{i}_{t}\}_{i=1}^{N}$
-Appendix B Additional remarks on mean-field estimation algorithms
-In our Algs.
-2 and 3, agents share their local counts with neighbours on the communication network $\mathcal{G}^{comm}_{t}$, and only after the $C_{e}$ communication rounds do they complete their estimated distribution by distributing the uncounted agents along their vectors.
-An alternative would be for each agent to immediately form a local estimate from their local count obtained via $\mathcal{G}^{obs}_{t}$ or $\mathcal{G}^{vis}_{t}$, which is only then communicated and updated via the communication network.
-However, we take the former approach to avoid poor local estimations spreading through the network and leading to widespread inaccuracies.
-Information that is certain (the count) is spread as widely as possible, before being locally converted into an estimate of the total mean field.
-The same would be the case in our extension proposed in Sec.
-E for averaging noisy counts, i.e.
-only the counts would be averaged, with the estimates completed by distributing the remaining agents after the $C_{e}$ communication rounds.
-Appendix C Experiments
+1, 2 and 5, no populations appear to have significantly different exploitability to each other, while in Fig.
+7 the central-agent population may have lower exploitability, but not significantly so.
+As discussed in Appx.
+A.2.1, the exploitability metric is noisy and provides limited information in these coordination games.
+Nevertheless we can see that in all four plots the independent agents hardly improve their returns at all, while the central-agent populations hardly improve their returns in the ‘target agreement’ games in Figs.
+1 and 5.
+There is therefore little a deviating agent can do to increase its return in these coordination games, meaning exploitability appears low, despite these being undesirable equilibria.
+Meanwhile, the networked agents do improve their returns and thus significantly outperform the stagnant independent agents in Figs.
+1, 2, 5 and 7 and the stagnant central-agent populations in Figs.
+1 and 5.
+This indicates that our communication scheme helps agents to reach ‘preferable’ equilibria, even when exploitability is similar (for reasons related to how networked populations can increase their returns faster as per Sec.
+6).
+While central-agent populations do appear to increase their returns in the ‘cluster’ task in Figs.
+2 and 7, they do so more slowly and reaching a lower final value than all networked agents in the $100\times 100$ grid case, and than all networked agents apart from those with the smallest broadcast radii in the $50\times 50$ grid case.
+Indeed in the $100\times 100$ grids in Figs.
+1 and 2, the central-agent populations appear to perform less well than they do in the $50\times 50$ grids in Figs.
+5 and 7, whereas the networked populations do not suffer a performance decrease, indicating that our networked communication scheme scales better and is robust to larger state spaces than the central-agent paradigm.
+Similarly, in the ‘target agreement’ and ‘cluster’ tasks in the tabular setting in benjamin2024networked, the central-agent populations generally perform similarly to the networked populations, indicating that our networked architecture is more robust than the central-agent alternative when moving to non-tabular settings.
+7.2.2. Population-dependent policies in complex environments
+We also show the ability of our algorithms to handle more complex tasks, using population-dependent policies and estimated mean-field observations.
+Figs.
+8(a) and 9(a) (Appx.
+A), where agents estimate the mean field via Alg.
+3, differ minimally from Figs.
+8(b) and 9(b), where agents directly receive the global mean field.
+This indicates that our estimation algorithm allows agents to appropriately estimate the distribution, even with only one round of communication for agents to help each other improve their local counts.
+Only in the ‘push object’ task in Fig.
+8(a), and there only with the smaller broadcast radii, do agents slightly underperform the returns of agents in the global observability case in Fig.
+8(b), as might be expected.
+For the reasons given in Appx.
+A.2.1 regarding coordination games, the exploitability metric gives limited information in the ‘push object’ and ‘evade’ tasks in Figs.
+8 and 9: for example, the return of a best-responding agent in the ‘push object’ task still depends on the extent to which other agents coordinate on which direction in which to push the box, meaning it cannot significantly increase its return by deviating.
+However, all of the networked cases significantly outperform the independent learners in terms of the average return to which they converge in both tasks.
+In the ‘push object’ task networked learners also appear to outperform central-agent populations, while in the ‘evade’ task all networked cases perform similarly to the central-agent case.
+Recall though that in the real world a central-agent architecture is a strong assumption, a computational bottleneck and single point of failure.
+8. Conclusion
+We introduced function approximation to the online-learning setting for empirical MFGs, and also contributed two novel algorithms for locally estimating the empirical mean field for population-dependent policies.
+We proved theoretically that our networked communication algorithms can learn faster than even central-agent architectures in this setting, and showed empirically that they can handle large state spaces and estimate the mean field.
+For future work, see Appx.
+C.
+Appendices
+Appendix A Experimental set-up
 Experiments were conducted on a Linux-based machine with 2 x Intel Xeon Gold 6248 CPUs (40 physical cores, 80 threads total, 55 MiB L3 cache).
 We use the JAX framework to accelerate and vectorise our code.
+For reproducibility, our code is included in our Supplementary Material.
 Random seeds are set in our code in a fixed way dependent on the trial number to allow easy replication of experiments.
-C.1 Games
-We conduct numerical tests with five games.
-All are defined by the agents’ reward/transition functions, and chosen for being particularly amenable to intuitive and visualisable understanding of whether the agents are learning behaviours that are appropriate and explainable for the respective objective functions.
-In all cases, rewards are normalised in [0,1] after they are computed.
+Figure 3.
+Larger version of Fig.
+1.
+‘Target agreement’, population-independent, $100\times 100$ grid.
+The networked populations of all broadcast radii significantly outperform the central-agent and independent populations in terms of average return, where the latter two cases hardly appear to learn at all.
+[…]
+‘Target agreement’, population-independent, $100\times 100$ grid.
+The networked populations of all broadcast radii significantly outperform the central-agent and independent populations in terms of average return, where the latter two cases hardly appear to learn at all.
+Figure 4.
+Larger version of Fig.
+2.
+‘Cluster’, population-independent, $100\times 100$ grid.
+Networked populations of all broadcast radii outperform the central-agent and independent populations wrt.
+average return; independent agents hardly appear to learn at all.
+[…]
+‘Cluster’, population-independent, $100\times 100$ grid.
+The networked populations of all broadcast radii outperform the central-agent and independent populations in terms of average return; independent agents hardly appear to learn at all.
+Figure 5.
+‘Target agreement’ task, population-independent policies, $50\times 50$ grid.
+The networked populations of all broadcast radii significantly outperform the central-agent and independent populations in terms of average return, where the latter two cases hardly appear to learn at all.
+[…]
+‘Target agreement’ task, population-independent policies, $50\times 50$ grid.
+The networked populations of all broadcast radii significantly outperform the central-agent and independent populations in terms of average return, where the latter two cases hardly appear to learn at all.
+Figure 6.
+‘Disperse’ task, population-independent policies, $100\times 100$ grid.
+The networked populations of all broadcast radii significantly outperform the central-agent and independent populations in terms of average return (and exploitability in the case of independent agents), with independent agents not learning at all.
+[…]
+‘Disperse’ task, population-independent policies, $100\times 100$ grid.
+The networked populations of all broadcast radii significantly outperform the central-agent and independent populations in terms of average return (and exploitability in the case of independent agents), with independent agents not learning at all.
+Figure 7.
+‘Cluster’ task, population-independent policies, $50\times 50$ grid.
+The networked agents of all broadcast radii significantly outperform the independent agents in terms of average return, where the latter case hardly appears to learn at all.
+The higher broadcast radii also appear to outperform the central-agent case in terms of average return, with the latter outperforming all others in terms of exploitability.
+[…]
+‘Cluster’ task, population-independent policies, $50\times 50$ grid.
+The networked agents of all broadcast radii significantly outperform the independent agents in terms of average return, where the latter case hardly appears to learn at all.
+The higher broadcast radii also appear to outperform the central-agent case in terms of average return, with the latter outperforming all others in terms of exploitability.
+(a) Estimated mean-field distribution.
+(b) Global observability of mean field.
+Figure 8.
+‘Push object’ task, population-dependent policies on a $10\times 10$ grid.
+The networked populations of all broadcast radii appear to outperform the central-agent and independent populations in terms of average return; the latter two cases hardly appear to learn at all.
+[…]
+‘Push object’ task, population-dependent policies on a $10\times 10$ grid.
+The networked populations of all broadcast radii appear to outperform the central-agent and independent populations in terms of average return; the latter two cases hardly appear to learn at all.
+(a) Estimated mean-field distribution.
+(b) Global observability of mean field.
+Figure 9.
+‘Evade’ task, population-dependent policies on a $10\times 10$ grid.
+The networked agents of all broadcast radii significantly outperform the independent agents in terms of average return, and perform similarly to the central-agent populations.
+[…]
+‘Evade’ task, population-dependent policies on a $10\times 10$ grid.
+The networked agents of all broadcast radii significantly outperform the independent agents in terms of average return, and perform similarly to the central-agent populations.
+A.1. Technical description of tasks
 Cluster.
-This is the inverse of the ‘exploration’ game in Laurière et al.
-[2022b], where in our case agents are encouraged to gather together by the reward function $R(s^{i}_{t},a^{i}_{t},\hat{\mu}_{t})=$ log$(\hat{\mu}_{t}(s^{i}_{t}))$.
-That is, agent $i$ receives a reward that is logarithmically proportional to the fraction of the population that is co-located with it at time $t$.
-We give the population no indication where they should cluster, agreeing this themselves over time.
-Agree on a single target.
+[…]
+Target agreement.
 Unlike in the above ‘cluster’ game, the agents are given options of locations at which to gather, and they must reach consensus among themselves.
 If the agents are co-located with one of a number of specified targets $\phi\in\Phi$ (in our experiments we place one target in each of the four corners of the grid), and other agents are also at that target, they get a reward proportional to the fraction of the population found there; otherwise they receive a penalty of -1.
-In other words, the agents must coordinate on which of a number of mutually beneficial points will be their single gathering place.
+In other words, the agents must coordinate on which of a number of mutually beneficial points will be their single gathering place to maximise their individual rewards.
 Define the magnitude of the distances between $x,y$ at $t$ as $dist_{t}(x,y)$.
 The reward function is given by $R(s^{i}_{t},a^{i}_{t},\hat{\mu}_{t})=r_{targ}(r_{collab}(\hat{\mu}_{t}(s^{i}_{t})))$, where
-$r_{targ}(x)=\begin{cases}x\quad&\text{if}\,\exists\phi\in\Phi\text{ s.t. dist}_{t}(s^{i}_{t},\phi)=0\\ -1\quad&\text{otherwise,}\\ \end{cases}$
-$r_{collab}(x)=\begin{cases}x\quad&\text{if}\,\hat{\mu}_{t}(s^{i}_{t})>1/N\\ -1\quad&\text{otherwise.}\\ \end{cases}$
+[…]
+[…]
 Evade shark in shoal.
+This is a similar idea to the task found in liu2025achievement.
 Define the magnitude of the horizontal and vertical distances between $x,y$ at $t$ as $dist^{h}_{t}(x,y)$ and $dist^{v}_{t}(x,y)$ respectively.
@@
 $x^{*}_{t}=\arg\max_{x\in\mathcal{S}}\hat{\mu}_{t}(x)$.
-A horizontal step is taken if $dist^{h}_{t}(\phi_{t},x^{*}_{t})$geq $dist^{v}_{t}(\phi_{t},x^{*}_{t})$, otherwise a vertical step is taken.
-As well as featuring a non-stationary distribution, we add ‘common noise’ to the environment, with the shark in a random direction with probability 0.01.
-Such noise that affects the local states of all agents in the same way, making the evolution of the distribution stochastic, makes population-independent policies sub-optimal Laurière et al.
-[2022a].
+A horizontal step is taken if $dist^{h}_{t}(\phi_{t},x^{*}_{t})$ $\geq$ $dist^{v}_{t}(\phi_{t},x^{*}_{t})$, otherwise a vertical step is taken.
+As well as featuring a non-stationary distribution, we add ‘common noise’ to the environment, with the shark moving in a random direction with probability 0.01.
+Such noise that affects the local states of all agents in the same way, making the evolution of the distribution stochastic, makes population-independent policies sub-optimal (survey_learningMFGs).
 Agents are rewarded more for being further from the shark, and also for clustering with other agents.
@@
 $dist^{h}_{t}(\phi_{t},\text{edge}^{*}_{t})+dist^{v}_{t}(\phi_{t},\text{edge}^{*}_{t}).$
+The above tasks are all coordination tasks, in that agents receive higher returns by aligning their policies and hence have an incentive to communicate their policy parameters even though the MFG framework is non-cooperative.
+Our fifth game is a non-coordination task: the reward function is not designed to give higher returns for more aligned policies.
+We include this game to demonstrate that even in such non-cooperative scenarios our networked architecture receives higher returns than both independent and central-agent alternatives, such that decentralised selfish agents may still have incentive to communicate.
+The fifth game is:
 Disperse.
-This is similar to the ‘exploration’ tasks in Laurière et al.
-[2022b], Wu et al.
-[2024] and other MFG works.
-In our version agents are rewarded for being located in more sparsely populated areas but only if they are stationary.
+This is similar to the ‘exploration’ tasks in scalable_deep, wu2024populationaware and other MFG works.
+In our version agents are rewarded for being located in more sparsely populated areas, but only if they are stationary.
 The reward function is given by $R(s^{i}_{t},a^{i}_{t},\hat{\mu}_{t})=r_{stationary}(-\hat{\mu}_{t}(s^{i}_{t}))$, where
-$r_{stationary}(x)=\begin{cases}x\quad&\text{if}\,a^{i}_{t}\text{ is `remain stationary'}\\ -1\quad&\text{otherwise.}\\ \end{cases}$
-C.2 Experimental metrics
+[…]
+A.2. Experimental metrics
 To give as informative results as possible about both performance and proximity to the MFNE, we provide two metrics for each experiment.
-Both metrics are plotted with mean and standard deviation, computed over the ten trials (each with a random seed) of the system evolution in each setting.
-C.2.1 Exploitability
-Works on MFGs most commonly use the exploitability metric to evaluate how close a given policy $\pi$ is to a NE policy $\pi^{*}$ Laurière et al.
-[2022b]; Perrin et al.
-[2020]; Laurière et al.
-[2022a]; Algumaei et al.
-[2023]; Pérolat et al.
-[2022]; Benjamin and Abate [2023]; Wu et al.
-[2024].
+Both metrics are plotted with mean and standard deviation, computed over the ten trials (each with a random seed) of the system run in each setting.
+A.2.1. Exploitability
+Works on MFGs most commonly use the exploitability metric to evaluate how close a given policy $\pi$ is to a NE policy $\pi^{*}$ (scalable_deep; continuous_fictitious_play; survey_learningMFGs; algumaei2023regularization; scaling_MFG; benjamin2024networked; wu2024populationaware).
 The metric usually assumes that all agents are following the same policy $\pi$, and quantifies how much an agent can benefit by deviating from $\pi$ by measuring the difference between the return given by $\pi$ and that of a $BR$ policy with respect to the distribution generated by $\pi$:
-Definition 9 (Exploitability of $\pi$).
-The exploitability ${E}x$ of policy $\pi$ is given by:
+Definition A.1 (Exploitability of $\pi$).
+The exploitability Ex{E}x of policy $\pi$ is given by:
 ${E}x(\pi)=V(BR(I(\pi)),I(\pi))-V(\pi,I(\pi)).$
 If $\pi$ has a large exploitability then an agent can significantly improve its return by deviating from $\pi$, meaning that $\pi$ is far from $\pi^{*}$, whereas an exploitability of 0 implies that $\pi=\pi^{*}$.
-Prior works conducting empirical testing have generally focused on the centralised setting, so this classical definition, as well as most evaluations, only consider exploitability when all agents are following a single policy $\pi_{k}$.
-However, Benjamin and Abate [2023] notes that purely independent agents, as well as networked agents, may have divergent policies $\pi^{i}_{k}\neq\pi^{j}_{k}$ $\forall i,k\in 1,\dots,N$, as in our own setting.
+Prior works conducting empirical testing have generally focused on the central-agent setting, so this classical definition, as well as most evaluations, only consider exploitability when all agents are following a single policy $\pi_{k}$.
+However, benjamin2024networked notes that purely independent agents, as well as networked agents, may have divergent policies $\pi^{i}_{k}\neq\pi^{j}_{k}$ $\forall i,k\in 1,\dots,N$, as in our own setting.
 We therefore are interested in the ‘exploitability’ of the population’s joint policy $\boldsymbol{\pi}$ := ($\pi^{1},\dots,\pi^{N}$) $\in\Pi^{N}$.
-Since we do not have access to the exact $BR$ policy as in some related works Laurière et al.
-[2022b]; Wu et al.
-[2024], we must instead approximate the exploitability, similarly to Perrin et al.
-[2021]; Benjamin and Abate [2023].
+Since we do not have access to the exact $BR$ policy as in some related works (scalable_deep; wu2024populationaware), we must instead approximate the exploitability, similarly to flock; benjamin2024networked.
 We freeze the policy of all agents apart from a deviating agent, for which we store its current policy and then conduct 50 $k$ loops of policy improvement.
 To approximate the expectations in Def.
-9, we take the best return of the deviating agent across 10 additional $k$ loops, as well as the mean of all the other agents’ returns across these same 10 loops.
-(While the policies of all non-deviating agents is $\pi_{k}$ in the centralised case, if the non-deviating agents do not share a single policy, then this method is in fact approximating the exploitability of their joint policy $\boldsymbol{\pi}^{-d}_{k}$, where $d$ is the deviating agent.) We then revert the agent back to its stored policy, before learning continues for all agents as per the main algorithm.
+A.1, we take the best return of the deviating agent across 10 additional $k$ loops, as well as the mean of all the other agents’ returns across these same 10 loops.
+(While the policies of all non-deviating agents is $\pi_{k}$ in the central-agent case, if the non-deviating agents do not share a single policy, then this method is in fact approximating the exploitability of their joint policy $\boldsymbol{\pi}^{-d}_{k}$, where $d$ is the deviating agent.) We then revert the agent back to its stored policy, before learning continues for all agents as per the main algorithm.
 Due to the expensive computations required for this metric, we evaluate it every second $k$ iteration of the main algorithm for Figs.
-1, 2, 3, 4 and 5, and every fourth iteration for the population-dependent experiments.
+1, 2, 5, 7 and 6, and every fourth iteration for the population-dependent experiments.
 The exploitability metric has a number of limitations in our setting.
-Our approximation takes place via MOMD policy improvement steps (as in the main algorithm) for an independent, deviating agent while the policies of the rest of the population are frozen.
-As such, the quality of our approximation is limited by the number of policy improvement/expectation rounds, which must be restricted for the sake of running speed of the experiments.
-Moreover, since one of the findings of our paper is that networked agents can improve their policies faster than independent or centralised agents, especially when non-linear function approximation is used, it is arguably unsurprising that approximating the $BR$ by an independently deviating agent sometimes gives an unclear and noisy metric.
+In coordination games (the setting for all tasks apart from the ‘disperse’ game), agents benefit by following the same behaviour as others, and so a deviating agent generally stands to gain less from switching to a $BR$ policy than it might in the non-coordination games on which many other works focus.
+For example, the return of a best-responding agent in the ‘cluster’ task still depends on the extent to which other agents coordinate on where to cluster, meaning it cannot significantly increase its return by deviating from a badly clustering policy.
+This means that the downward trajectory of the exploitability metric is less clear in our plots than in other works that do not focus on coordination games.
+This is likely why the difference between the approximated exploitability of the independent agents and the other populations is clearer in the non-coordination ‘disperse’ task in Fig.
+6 than in the other tasks.
+Moreover, our approximation takes place via MOMD policy improvement steps (as in the main algorithm) for an independent, deviating agent while the policies of the rest of the population are frozen.
+As such, the quality of our approximation is limited by the number of policy-improvement/expectation-estimation rounds, which must be restricted for the sake of the running speed of the experiments.
+Moreover, since one of the findings of our paper is that independent-learning agents increase their returns significantly slower (if at all) than networked or central-agent populations, it is arguably unsurprising that approximating the $BR$ by an independently deviating agent sometimes gives an unclear and noisy metric.
 This includes the exploitability going below zero, which should not be possible if the policies and distributions are computed exactly.
-Moreover, in coordination games (the setting for all tasks apart from the ‘disperse’ game), agents benefit by following the same behaviour as others, and so a deviating agent generally stands to gain less from a $BR$ policy than it might in the non-coordination games on which many other works focus.
-For example, the return of a best-responding agent in the ‘push object’ game still depends on the extent to which other agents coordinate on which direction in which to push the box, meaning it cannot significantly increase its return by deviating.
-This means that the downward trajectory of the exploitability metric is less clear in our plots than in other works.
-This is likely why the approximated exploitability gets lower in the non-coordination ‘disperse’ task in Fig.
-5 than in the other tasks.
-Given the limitations presented by approximating exploitability, we also provide the second metric to indicate the progress of learning.
-C.2.2 Average discounted return
-We record the average discounted return of the agents’ policies $\pi_{k}^{i}$ during the $M$ iterations - this allows us to observe that settings that converge to similar exploitability values may not have similar average agent returns, suggesting that some algorithms are better than others at finding not just NE, but preferable NE.
+Given the limitations presented by approximating exploitability, we also provide a second metric to indicate the progress of learning, as in benjamin2024networked.
+A.2.2. Average discounted return
+We record the average discounted return of the agents’ policies $\pi_{k}^{i}$ during the $M$ iterations.
+This allows us to compare how quickly agents are learning to increase their returns, even when exploitability gives us limited ability to distinguish between the desirability of the MFNEs to which populations converge.
+We can observe that settings that converge to similar exploitability values may not have similar average agent returns, suggesting that some algorithms are better than others at finding not just NE, but preferable NE (for reasons related to why networked populations can increase their returns faster as per Sec.
+6).
 See for example Figs.
-3 and 1, where the networked agents converge to similar exploitability as the independent and centralised agents, but receive higher average returns.
-C.3 Hyperparameters
-Figure 3: ‘Target agreement’ task, population-independent policies, 50x50 grid.
-Figure 4: ‘Cluster’ task, population-independent policies, 50x50 grid.
-Figure 5: ‘Disperse’ task, population-independent policies, 100x100 grid.
+1 and 5, where the networked populations converge to similar exploitability as the independent and central-agent populations, but receive higher average returns.
+A.3. Hyperparameters
 See Table 1 for our hyperparameter choices.
 We can group our hyperparameters into those controlling the size of the experiment, those controlling the size of the Q-network, those controlling the number of iterations of each loop in the algorithms and those affecting the learning/policy updates or policy adoption.
-In our experiments we generally want to demonstrate that our communication-based algorithms outperform the centralised and independent architectures by allowing policies that are estimated to be better performing to proliferate through the population, such that convergence occurs within fewer iterations and computationally faster, even when the Q-function is poorly approximated and/or the mean-field is poorly estimated, as is likely to be the case in real-world scenarios.
+In our experiments we generally want to demonstrate that our communication-based algorithms outperform the central-agent and independent architectures by allowing policies that are estimated to be better performing to proliferate through the population, such that convergence occurs within fewer iterations and computationally faster, even when the Q-function is poorly approximated and/or the mean-field is poorly estimated, as is likely to be the case in real-world scenarios.
 Moreover we want to show that there is a benefit even to a small amount of communication, so that communication rounds themselves do not excessively add to time complexity.
 As such, we generally select hyperparameters at the lowest end of those we tested during development, to show that our algorithms are particularly successful given what might otherwise be considered ‘undesirable’ hyperparameter choices.
+Table 1. Hyperparameters
 Hyperparam. Value Comment
 Trials 10 We run 10 trials with different random seeds for each experiment. We plot the mean and standard deviation for each metric across the trials.
-Gridsize 10x10 / 50x50 / 100x100 Experiments with population-dependent policies are run on the 10x10 grid (Figs.
-6, 7, 8 and 9), while experiments on large state spaces are run on 50x50 and 100x100 grids (Figs.
-1, 2, 3, 4 and 5).
+Gridsize $\times1010$/ $\times5050$/ $\times100100$ Experiments on large state spaces are run on $50\times 50$ and $100\times 100$ grids (Figs.
+1-6), while experiments with population-dependent policies are run on the $10\times 10$ grid (Figs.
+8(a), 9(a), 8(b) and 9(b)).
 Population 500 We chose 500 for our demonstrations to show that our algorithm can handle large populations, indeed often larger than those demonstrated in other mean-field works, especially for grid-world environments, while also being feasible to simulate wrt.
-time and computation constraints Yongacoglu et al.
-[2022]; Cui and Koeppl [2021]; Cui et al.
-[2023]; Guo et al.
-[2023]; Subramanian and Mahajan [2019]; Yang et al.
-[2018]; Ganapathi Subramanian et al.
-[2021, 2020]; Subramanian et al.
-[2022]; Wu et al.
-[2024]; Benjamin and Abate [2023].
+time and computation constraints (subjective_equilibria; approximately_entropy; cui2023multiagent; general_framework; rl_stationary; mfrl_yang; pomfrl; subramanian2022multi; DMFG; wu2024populationaware; benjamin2024networked).
 Number of neurons in input layer cf.
@@
 For population-dependent policies, the mean-field distribution is a flattened vector of the same size as the grid.
-As such, the input size in the ‘evade’ and ’push object’ tasks is $[(4\times\mathrm{dimension})+(\mathrm{dimension}^{2})]$; in the other settings it is $[2\times\mathrm{dimension}]$.
+As such, the input size in the ‘evade’ and ’push object’ tasks is $(4\times\mathrm{dimension})+(\mathrm{dimension}^{2})]$; in the other settings it is $2\times\mathrm{dimension}]$.
 Neurons per hidden layer cf.
@@
 It may be possible to reduce this value further and still achieve similar results.
-$C_{p}$ 1 As in Benjamin and Abate [2023], we choose this value to show the convergence benefits brought by even a single communication round, even in networks that may have limited connectivity; higher choices are likely to have even better performance.
+$C_{p}$ 1 As in benjamin2024networked, we choose this value to show the convergence benefits brought by even a single communication round, even in networks that may have limited connectivity; higher choices are likely to have even better performance.
 $C_{e}$ 1 Similar to $C_{p}$, we choose this value to show the ability of our algorithm to appropriately estimate the mean field even with only a single communication round, even in networks that may have limited connectivity.
@@
 However, only 0.03 gave stable increase in return.
-Note that this is the value also chosen in Vieillard et al.
-[2020].
+Note that this is the value also chosen in NEURIPS2020_2c6a0bae.
 $|B|$ 32 This is a common choice of batch size that trades off noisy updates and computational efficiency.
-$cl$1 We use the same value as in Vieillard et al. [2020].
-$\nu$ $L-1$ We tested $\nu$ in $\{1,4,20,L-1\}$.
+clcl -1 We use the same value as in NEURIPS2020_2c6a0bae.
+$\nu$ L1L-1 We tested $\nu$ in $\{1,4,20,L-1\}$.
 We found that in our setting, updating $\theta^{\prime}\leftarrow\theta$ once per $k$ iteration s.t.
 $\theta^{\prime}_{k+1,l}=\theta_{k,l}$ $\forall l$ gave sufficient learning that was similar to the other potential choices of $\nu$, so we do this for simplicity, rather than arbitrarily choosing a frequency to update $\theta^{\prime}$ during each $k$ loop.
-Setting the target to be the policy from the previous iteration is similar to the method in Laurière et al.
-[2022b].
-Whilst Wu et al.
-[2024] updates the target within the $L$ loops for stability, we do not find this to be a problem in our experiments.
-Optimiser Adam As in Vieillard et al. [2020], we use the Adam optimiser with initial learning rate 0.01.
+Setting the target to be the policy from the previous iteration is similar to the method in scalable_deep.
+Whilst wu2024populationaware updates the target within the $L$ loops for stability, we do not find this to be a problem in our experiments.
+Optimiser Adam As in NEURIPS2020_2c6a0bae, we use the Adam optimiser with initial learning rate 0.01.
 $\tau^{comm}_{k}$ cf.
 comment $\tau^{comm}_{k}$ increases linearly from 0.001 to 1 across the $K$ iterations.
-This is a simplification of the annealing scheme used in Benjamin and Abate [2023].
+This is a simplification of the annealing scheme used in benjamin2024networked.
 Further optimising the annealing process may lead to better results.
-Table 1: Hyperparameters
-C.4 Additional experiments
-C.4.1 Additional experiments on large grids
-We provide additional experiments on large grids in Figs. 3, 4 and 5.
-In the ‘target agreement’ task in Fig.
-3, the networked agents generally have lower exploitability than both centralised and independent agents, and significantly outperform the other architectures in terms of average return.
-As before, the margin by which the networked agents can outperform the centralised agents is much greater than in Benjamin and Abate [2023], showing that the benefits of the communication scheme are even greater in non-tabular settings.
-In the ‘cluster’ task in Fig.
-4, the networked agents obtain significantly higher return than the independent agents.
-While centralised agents have the lowest exploitability, networked agents of almost all communication radii outperform them in terms of average return, indicating that the communication scheme helps populations reach better performing equilibria.
-In the ‘disperse’ task in Fig.
-5, networked agents significantly outperform independent and centralised agents in terms of average return.
-They also outperform centralised agents in terms of exploitability, and significantly outperform independent agents in terms of exploitability.
-The fact that this happens in this non-coordination, competitive game shows that agents do have an incentive to communicate with each other even if they are self-interested.
-C.4.2 Population-dependent policies in complex environments
-Figure 6: ‘Push object’ task, population-dependent policies with estimated mean-field distribution, 10x10 grid.
-Figure 7: ‘Evade’ task, population-dependent policies with estimated mean-field distribution, 10x10 grid.
-Figure 8: ‘Push object’ task, population-dependent policies with global observability of mean field, 10x10 grid.
-Figure 9: ‘Evade’ task, population-dependent policies with global observability of mean field, 10x10 grid.
-We also showcase the ability of our algorithm to handle two more complex tasks, using population-dependent policies and estimated mean-field observations.
-Figs.
-6 and 7, where agents estimate the mean-field distribution via Alg.
-3, differ minimally from Figs.
-8 and 9, where agents directly receive the global mean-field distribution.
-This shows that our estimation algorithm allows agents to appropriately estimate the distribution, even with only one round of communication for agents to help each other improve their local counts.
-Only in the ‘push object’ task in Fig.
-6, and there only with the smaller broadcast radii, do agents slightly underperform the returns of agents in the global observability case in Fig.
-8, as is reasonable.
-For the reasons given in Sup.
-Mat.
-C.2.1, the exploitability metric gives limited information in the ‘push object’ task in Fig.
-6.
-In the ‘evade’ task in Fig.
-7, exploitability suggests that centralised learners outperform the other cases.
-However, all of the networked cases significantly outperform the independent learners in terms of the average return to which they converge in both tasks.
-In the ‘push object’ task networked learners also significantly outperform centralised learners in all but the case with the smallest broad communication radius, while in the ‘evade’ task all networked cases perform similarly to the centralised case.
-Recall though that in the real world a centralised architecture is a strong assumption, a computational bottleneck and single point of failure.
-Appendix D Related work
-MFGs are a quickly growing research area, so we only discuss the works most closely related to this present work, and instead refer the reader to Benjamin and Abate [2023] for detailed discussion around the setting of networked communication for MFGs, and to Laurière et al.
-[2022a] for a broader survey of MFGs.
-Our work is most closely related to Benjamin and Abate [2023], which introduced networked communication to the infinite-horizon MFG setting.
-However, this work focuses only on tabular settings rather than using function approximation as in ours, and only addresses population-independent policies.
-Laurière et al.
-[2022b] uses Munchausen Online Mirror Descent (MOMD), similar to our method for learning with neural networks, but there are numerous differences to our setting: most relevantly, they study a finite-horizon episodic setting, where the mean-field distribution is updated in an exact way and an oracle supplies a centralised learner with rewards and transitions for it to learn a population-independent policy.
-Wu et al.
-[2024] uses MOMD to learn population-dependent policies, albeit also with a centralised method that exactly updates the mean-field distribution in a finite-horizon episodic setting.
-Perrin et al.
-[2022] learns population-dependent policies with function approximation in infinite-horizon settings like our own, but does so in a centralised, two-timescale manner without using the empirical mean-field distribution.
-Zhang et al.
-[2024] also uses function approximation along a non-episodic path, but involves a generic central agent learning an estimate of the mean field rather than using an empirical population.
-Approaches that directly update an estimate of the mean field must be able to generate rewards from this arbitrary mean field, even if they otherwise claim to be oracle-free.
-They are thus inherently centralised algorithms and rely on strong assumptions that may not apply in real-world problems.
-Conversely, we are interested in practical convergence in online, deployed settings, where the reward is computed from the empirical finite population.
-Yongacoglu et al.
-[2022] addresses decentralised learning from a continuous, non-episodic run of the empirical system using either full or compressed information about the mean-field distribution, but agents are assumed to receive this information directly, rather than estimating it locally as in the algorithm we now present.
-They also do not consider function approximation or inter-agent communication in their algorithms.
-In the closely related but distinct area of Mean-Field RL, Ganapathi Subramanian et al.
-[2021] does estimate the empirical mean-field distribution from the local neighbourhood, however agents are seeking to estimate the mean action rather than the mean-field distribution over states as in our MFG setting.
-Their agents also do not have access to a communication network by which they can improve their estimates.
-Appendix E Limitations and future work
+A.4. Additional experiments and plots
+In the non-coordination ‘disperse’ task in Fig.
+6, networked agents significantly outperform independent and central-agent populations in terms of average return.
+They also significantly outperform independent agents in terms of exploitability, and perhaps also central-agent populations though not significantly so.
+The fact that this happens in this non-coordination, competitive game shows that agents may have an incentive to communicate with each other even if they are self-interested.
+Indeed the agents learning independently do not appear to improve their returns at all, despite this being the paradigm that might be expected to perform best in a non-cooperative setting.
+Appendix B Further discussion on the necessity of population-dependent policies
+There are numerous reasons why agents may benefit from being able to respond to the current distribution. These include when:
+  • There may in fact be any initial distribution (perrin2022generalization; wu2024populationaware; survey_learningMFGs).
+  • The solved MFG has a non-stationary equilibrium, e.g.
+if there is a finite time horizon meaning that actions become more urgent as time passes such that policies (and therefore also the distribution) are non-stationary (scalable_deep; wu2024populationaware).
+  • The distribution is subject to ‘common noise’.
+This is noise that affects the local states of all agents in the same way, making the evolution of the distribution stochastic.
+In such cases, even if the agent knows the policy used by all other agents, it cannot perfectly predict the evolution of the mean-field distribution, making population-independent policies suboptimal (carmona_common; cardaliaguet2015masterequationconvergenceproblem; survey_learningMFGs; li2024incomplete; becherer2024common).
+  • The distribution deviates from the equilibrium for some other reason, such that agents need to be able to generalise their response to other (possibly previously unseen) distributions (survey_learningMFGs).
+Appendix C Future work
 Our work follows the gold standard in MFGs by presenting experiments on grid world toy environments, albeit we show our algorithms are able to handle much larger and more complex games than prior work.
@@
 In Sec.
-5 we give theoretical results showing that our networked algorithm can outperform a centralised alternative.
-We leave more general analysis, such proof of convergence and sample guarantees in the function approximation setting, for future work.
+6 we give theoretical results showing that our networked algorithm can outperform a central-agent alternative.
+We leave more general analysis, such as proof of convergence and sample guarantees in the function approximation setting, for future work.
+Since the MFG setting is technically non-cooperative, we have preempted objections that agents would not have incentive to communicate their policies by focusing on coordination games, i.e.
+where agents seek to maximise only their individual returns, but receive higher rewards when they follow the same strategy as other agents.
+In this case they stand to benefit by exchanging their policies with others.
+Future work lies in extending our networked communication algorithms to mean-field control, the cooperative counterpart to MFGs, where agents would have incentive to communicate across different types of game.
+Nevertheless, in real-world settings, the communication network could still be vulnerable to malfunctioning agents or adversarial actors poisoning the equilibrium by broadcasting untrue policy information (agrawal2024impact), or equally to unreliable communication channels.
+It is outside the scope of this paper to analyse how much false information would have to be broadcast by how many agents to affect the equilibrium, but real-world applications may need to compute this and prevent it.
+Future research to mitigate this risk might build on work such as piazza2024power, where ‘power regularisation’ of information flow is proposed to limit the adverse effects of communication by misaligned agents.
+Similarly, for the non-tabular MFG setting in benjamin2024networked, the work considers the robustness of the population to scenarios when agents fail to update their policies by the time they are required to communicate.
+Similar studies in our setting are the matter of future work.
+Likewise, our experiments only use one round of communication within each iteration to show the benefit that even this can have on learning speed over the independent case, but it may be informative to study the effect of communication costs further.
 Alg.
@@
 However, as we note above, there are various reasons why sharing IDs might be undesirable, including privacy and scalability.
-If instead only counts are taken, and if the noise on each agents’ count is assumed to be independent and, for example, subject to a Gaussian distribution, the algorithm can easily be updated such that communicating agents compute averages of their local and received counts to improve their accuracy, rather than simply using communication to fill in counts for previously unobserved states.
+If instead only counts are taken, and if the noise on each agents’ count is assumed to be independent and, for example, subject to a Gaussian distribution, the algorithm can easily be updated such that communicating agents compute averages of their local and received counts to improve their accuracy, rather than simply using communication to fill in counts for previously unseen states.
 (Note that we can also consider the original case without noise to involve averaging, since averaging identical values equates to using the original value).
@@
 Such choices may be domain-specific based on knowledge of a particular environment.
-For example, one might use the counts to perform Bayesian updates on a specific prior, where this prior may relate to the estimated mean-field distribution at the previous time step $t-1$.
-If agents seek to learn to predict the evolution of the mean field based on their own policy or by learning a model, the Bayesian prior may also be based on forward prediction from the estimated mean-field distribution at $t-1$.
+For example, one might use the counts to perform Bayesian updates on a specific prior, where this prior may relate to the estimated mean-field distribution at the previous time step t1t-1.
+If agents seek to learn to predict the evolution of the mean field based on their own policy or by learning a model, the Bayesian prior may also be based on forward prediction from the estimated mean-field distribution at t1t-1.
 Future work lies in conducting experiments in all of these more general settings.
-Perrin et al.
-[2022] notes that in grid-world settings such as those in our experiments, passing the (estimated or true global) mean-field distribution as a flat vector to the Q-network ignores the geometric structure of the problem.
+perrin2022generalization notes that in grid-world settings such as those in our experiments, passing the (estimated or true global) mean-field distribution as a flat vector to the Q-network ignores the geometric structure of the problem.
 They therefore propose to create an embedding of the distribution by first passing the vector to a convolutional neural network, essentially treating the categorical distribution as an image.
-This technique is also followed in Wu et al.
-[2024] (for their additional experiments, but not in the main body of their paper).
+This technique is also followed in wu2024populationaware (for their additional experiments, but not in the main body of their paper).
 As future work, we can test whether such a method improves the performance of our algorithms.